### v1

# Semiparametric Skew-Elliptical Distributions For High-Dimensional Graphical Models

Abstract
We propose a semiparametric approach called elliptical skew-(S)KEPTIC for efficiently and robustly estimating non-Gaussian graphical models. Relaxing the assumption of semiparametric elliptical distributions to the family of meta skew-elliptical that accommodates a skewness component, we derive a new estimator which is an extension of the SKEPTIC estimator in Liu et al., 2012a , based on semiparametric Gaussian copula graphical models, to the case of skew-elliptical copula graphical models. Theoretically, we demonstrate that the elliptical skew-(S)KEPTIC estimator achieves robust parametric convergence rates in both graph recovery and parameters estimation. We conduct numerical simulations to prove the reliable graph recovery performance of the elliptical skew-(S)KEPTIC estimator. Finally, the new method is applied to the daily log-returns of the stocks of the S&P500 index and shows better interpretability compared to the Gaussian copula graphical models.
Keywords—High-dimensional statistics, undirected graphical models, elliptical distributions, skewness, robust statistics.
1 Introduction
We consider the problem of estimating high-dimensional undirected graphical models. Given a $p$-dimensional random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$, we aim to estimate the undirected graph $G:=(V,E)$, where the vertices set $V:=\{1,\ldots,p\}$ contains the number of nodes corresponding to the $p$ variables, and the edge set $E$ indicates the conditional independence relationships between $X_{1},\ldots,X_{p}$. Let $X_{i}$ and $X_{j}$ be random variables associated with the multivariate distribution $\boldsymbol{X}$ and define $X_{\setminus{\{i,j}\}}$ the set of variables excluding $X_{i}$ and $X_{j}$. If $X_{i}$ and $X_{j}$ are conditionally independent given $X_{\setminus{\{i,j}\}}$, then there exists no edge between the corresponding nodes $i$ and $j$ in the graph G, denoted as $(i,j)\notin E$.
One reliable approach to estimate the undirected graphical models is the semiparametric nonparanormal (NPN) method proposed in Liu et al., (2009), which relaxes the assumption of Gaussian data, considering a set of functions $\{f_{j}\}_{j=1}^{p}$ such that: $f(\boldsymbol{X})=(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim N(0,\Sigma^{0})$, where $\Omega=(\Sigma^{0})^{-1}$. When $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$, no edge connects $X_{i}$ and $X_{j}$ if $\Omega_{ij}=0$. Uncovering sparsity is subject to the estimation of the unknown correlation matrix. Liu et al., 2012a propose the SKEPTIC (Spearman/Kendall estimates preempt transformations to infer correlation) estimator for the nonparanormal distribution, a regularized rank-based correlation coefficient estimator based on Kendall’s tau and Spearman’s rho statistics. The estimated correlation matrix is then puggled into shrinkage procedures that reveal the sparsity of the graph.
The nonparanormal distribution exploits a Gaussian-based kernel and belongs to the broad class of meta-elliptical distributions described in Han and Liu, (2014), as well as Liu et al., 2012b , where a set of functions $\{f_{j}\}_{j=1}^{p}$ such that $f(\boldsymbol{X})\sim\mathcal{E}C(0,\Sigma^{0},\xi)$ is considered. In particular, Liu et al., 2012b show that the graph $G$ is encoded in the latent generalized partial correlation matrix $\Xi:=-\left[\text{diag}(\Omega)\right]^{-1/2}\Omega\left[\text{diag}(\Omega)\right]^{-1/2}$, thus no edge connects $X_{i}$ and $X_{j}$ if $\Xi_{ij}=0$, which implies $\Omega_{ij}=0$. The main drawback of the meta-elliptical class of distributions is that, even though it embeds skewed distributions, the skewness term does not contribute to reveal the conditional independence structure. For this reason, we introduce the meta skew-elliptical family of distributions defined as $\boldsymbol{X}\sim MS\mathcal{E}(0,\Sigma^{0},\Lambda;f)$ considering a set of functions $f$ such that $f(\boldsymbol{X})\sim SU\mathcal{E}(0,\Sigma^{0},\Lambda,h^{(p+q)},\tau,\Gamma)$ described in Arellano-Valle and Genton, 2010b .
In this paper, we propose a rank-based correlation matrix estimator, which we refer to as elliptical skew-(S)KEPTIC estimator, extending the SKEPTIC estimator to the $MS\mathcal{E}$ family of distributions, adding respect to nonparanormal models the possibility to rely on graphical models that take into account other distributional features such as directional skewness and heavy tails. The main idea is to consider the $\operatorname{SU\mathcal{E}}$ as a combination of the closed skew-Normal distribution introduced in Domìnguez-Molina et al., (2003), which belongs to the class of $\operatorname{SU\mathcal{E}}$ family of distributions, and possesses the stochastic representation defined through Normal distribution and a half-Normal distribution, for which the SKEPTIC estimator of the correlation matrix is defined. The estimated correlation matrix is plugged into existing shrinkage procedures. For instance, Meinshausen and Bühlmann, (2006) study a neighborhood selection approach with the Lasso operator, Banerjee et al., (2008) and Friedman et al., (2008) propose an approach based on penalized likelihood, and Yuan, (2010) and Cai et al., (2011) introduce, respectively, Graphical Dantzig selector and CLIME which are two methods based on linear programming that provide an improvement in terms of theoretical properties respect to the other shrinkage procedures.
We generalize the results in Liu et al., 2012a extending the SKEPTIC estimator to the elliptical skew-(S)KEPTIC estimator by introducing a directional skewness term, encoded in the matrix $\Lambda$, which assigns greater importance to the asymmetry present in data distributions. The proposed estimator is a version of the SKEPTIC estimator conforming to the ${MS}\mathcal{E}$ distribution and embeds two new subclasses of estimators based on Kendall’s tau and Sparman’s rho statistics. While the nonparanormal distribution is a flexible approach to model non-Normal data, the meta skew-elliptical distribution allows us to quantify the asymmetry (and the kurtosis) presented in the data. When the $\operatorname{SU\mathcal{E}}$ distribution reduces to the $\operatorname{CSN}$ distribution, we consider both the statistics for estimating the unknown correlation matrix and we define the elliptical skew-SKEPTIC estimator. Otherwise, we only consider Kendall’s tau statistic, and we define elliptical skew-KEPTIC estimator since Spearman’s rho statistic is not invariant for the class of elliptical distributions (Hult and Lindskog, , 2002).
The proposed framework extends to a greater variety of semiparametric graphical models, introducing the possibility of modeling the combination of directional skewness and heavy tails data and adding flexibility to the nonparanormal-type of data transformations.
We investigate the theoretical properties of the elliptical skew-(S)KEPTIC estimator, and find bounds equal to twice the bounds of the SKEPTIC estimator. The increase in model flexibility justifies the slight reduction in the robustness of the proposed estimators.
We provide numerical studies to support our theory. Through a backward approach, we define the true sparsity pattern given by $\Omega$ and provide the graph recovery using the elliptical skew-(S)KEPTIC estimator which is compared to the SKEPTIC estimator. The elliptical skew-(S)KEPTIC estimator proves an optimal graph recovery, giving statistical efficiency to the proposed method.
Prior studies by Zareifard et al., (2016), Nghiem et al., (2022), and Sheng et al., (2023) have explored graphical models while incorporating skewness into the sparsity pattern estimation. Zareifard et al., (2016) exploit a multivariate $\operatorname{CSN}$ distribution to define the skew Gaussian graphical models which are estimated through a Bayesian approach. Nghiem et al., (2022) propose a novel nodewise regression approach to estimate the graph on data generated from a generalized multivariate skew-Normal distribution. Instead, Sheng et al., (2023) include the shape parameter in the determination of the precision matrix using the skew-Normal distribution of Azzalini, (1985). They build an algorithm that penalizes the likelihood of this distribution in order to estimate $\Omega$ and thus the graph. However, these studies do not emphasize the direct estimation of the correlation matrix while accounting for skewness.
The paper is organized as follows: in Section 2 we recall the background, in Section 3 we define the ${MS}\mathcal{E}$ distribution, which we use to derive the elliptical skew-(S)KEPTIC estimator and in Section 4 we study the theoretical properties of the proposed estimator. In Section 5, we conduct experiments comparing elliptical skew-(S)KEPTIC estimator with the SKEPTIC estimator through numerical comparisons on synthetic data and with an empirical analysis on the S&P 500 stocks data. Section 6 concludes.
2 Background
In this section, we briefly recall the literature preceding our article: the SKEPTIC estimator of Liu et al., 2012a , the semiparametric ellitpical distributions in Liu et al., 2012b and Han and Liu, (2014), in this work, we rely of the meta-elliptical distribution.. The section concludes with the introduction of $\operatorname{SU\mathcal{E}}$ family of distributions in Arellano-Valle and Genton, 2010b , and its link with the undirected graphical models.
2.1 Notation
Here is the notation used throughout the paper. Let $A=\left[A_{ij}\right]\in\mathbb{R}^{p\times p}$ and $\boldsymbol{a}=(a_{1},\ldots,a_{p})\in\mathbb{R}^{p}$. For $1\leq q<\infty$, we define $\|\boldsymbol{a}\|_{q}=\left(\sum_{i=1}^{p}|\boldsymbol{a}|^{q}\right)^{1/q}$, when $q=\infty$: $\|\boldsymbol{a}\|_{\infty}=\max_{1\leq i\leq p}|\boldsymbol{a}_{i}|$. Regarding the $A$ matrix, when $q=1$: $\|A\|_{1}=\max_{1\leq j\leq q}\sum_{i=1}^{p}|A_{ij}|$, and when $q=\infty$, $|A|_{\infty}=\max_{1\leq i\leq q}\sum_{j=1}^{p}|A_{ij}|$. We denote $\|A\|_{\max}=\max_{i,j}|A_{ij}|$. A vector without its $j$-th element is defined: $\boldsymbol{a}_{\backslash j}=(\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{j-1},\boldsymbol{a}_{j+1},\ldots,\boldsymbol{a}_{p})^{\top}\in\mathbb{R}^{p-1}$, and we denote $A_{\backslash i,\backslash j}\in\mathbb{R}^{p-1\times p-1}$ the matrix without the $i$-$th$ row and the $j$-th column. We define: $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ respectively the minimum and the maximum eigenvalue of $A$. $\Omega$ denotes a sparse matrix, and we define the degree of $\Omega$: $\operatorname{deg}(\Omega)=\max_{1\leq i\leq p}\sum_{j=1}^{p}\operatorname{I}(|\Omega_{ij}|\neq 0)$.
2.2 The Nonparanormal SKEPTIC
A random vector $\boldsymbol{X}=(X_{1},...,X_{p})^{\top}$ has a nonparanormal distribution if there exists functions $\{f_{j}\}_{j=1}^{p}$ such that $\boldsymbol{Z}\equiv f(\boldsymbol{X})\sim\operatorname{N}(0,\Sigma^{0})$, where $f(\boldsymbol{X})=(f_{1}(X_{1}),...,f_{p}(X_{p}))^{\top}$, we then write $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$. Further, if the functions $\{f_{j}\}_{j=1}^{p}$ are monotone and differentiable, the joint distribution of $\boldsymbol{X}$ is a Gaussian copula. The focal point is to find a function, $f(\boldsymbol{X})$, that transforms our data into normally distributed data. Suppose we have $n$ observations for the $j$-th variable, $x_{1j},...,x_{nj}$, we compute the rank of each observation, and indicate the rank with $r_{1j},...,r_{nj}$. Let $\Phi(\cdot)$ denote the Gaussian cumulative distribution function, the $f(\boldsymbol{X})$ that transforms the data is:
$\hat{f}_{j}(x_{ij})=\Phi^{-1}(\hat{r}_{ij}),\quad\text{where:}\quad\hat{r}_{ij}=\frac{r_{ij}}{n+1}.$ (1)
We consider the following statistics:
[…]
Both $\hat{\rho}_{ij}$ and $\hat{\tau}_{ij}$ are non-parametric correlations between the empirical realizations of random variables $X_{j}$ and $X_{i}$. Let $\widetilde{X}_{i}$ and $\widetilde{X}_{j}$ be independent copies of the random variables $X_{i}$ and $X_{j}$, respectively. The cumulative distribution functions (CDFs) of $X_{i}$ and $X_{j}$ are denoted by $F_{i}$ and $F_{j}$. The population versions of Spearman’s rho and Kendall’s tau are defined as $\rho_{ij}:=\operatorname{Corr}(F_{i}(X_{i}),F_{j}(X_{i}))$ and $\tau_{ij}:=\operatorname{Corr}(\operatorname{sign}(X_{i}-\widetilde{X}_{i}),\operatorname{sign}(X_{j}-\widetilde{X}_{j}))$, respectively. For nonparanormal distributions, there is a connection between Spearman’s rho and Kendall’s tau to the underlying Pearson’s correlation coefficient $\Sigma^{0}_{ij}$. We recall the SKEPTIC estimator.
Lemma 2.1.
(Kruskal, (1958), Liu et al., 2012a ): Assuming $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$, we have $\Sigma_{ij}^{0}=2\operatorname{sin}\left(\frac{\pi}{6}\rho_{ij}\right)=\operatorname{sin}\left(\frac{\pi}{2}\tau_{ij}\right)$. From these results, we define two estimators for the unknown correlation matrix $(\Sigma^{0}_{ij})$ that are $\hat{S}^{\rho}=\left[\hat{S}^{\rho}_{ij}\right]$ and $\hat{S}^{\tau}=\left[\hat{S}^{\tau}_{ij}\right]$, where: $\hat{S}^{\rho}_{ij}=2\operatorname{sin}\left(\frac{\pi}{6}\hat{\rho}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ and $\hat{S}^{\tau}_{ij}=\operatorname{sin}\left(\frac{\pi}{2}\hat{\tau}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$,
Spearman’s rho and Kendall’s tau statistics are invariant under strictly increasing transformations of the underlying random variable. This concept will be fundamental in the definition of the elliptical skew-(S)KEPTIC estimator.
2.3 The meta-elliptical distribution
The meta-elliptical distribution introduced in Han and Liu, (2014) relaxes the assumption of elliptically distributed data, in the same way as the nonparanormal of Liu et al., (2009) do for Gaussian distributions. The meta-elliptical family is a strict extension of the nonparanormal distribution.
Definition 2.2.
[…]
Let $\mathcal{R}_{p}:=\{\Sigma\in\mathbb{R}^{p\times p}:\Sigma^{\top}=\Sigma,\text{diag}\left(\Sigma\right)=1,\Sigma\succeq 0\}$, in analogy to the nonparanormal case, the meta-elliptical distribution is defined as follows.
Definition 2.3.
(Han and Liu, (2014)) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has a meta-elliptical distribution, denoted by $\boldsymbol{X}\sim M\mathcal{E}_{p}(\Sigma^{0},\xi;f)$, if there exist a set of univariate monotone functions $f_{1},\ldots,f_{p}$ such that $(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\mathcal{EC}_{p}(0,\Sigma^{0},\xi)$, where $\Sigma^{0}\in\mathcal{R}_{p}$ is defined as the latent generalized correlation matrix.
For the distributions that belong to the meta-elliptical family, Kendall’s tau statistic in Lemma 2.1 can be exploited to estimate the unknown correlation matrix. Nevertheless, the meaning of a missing edge is different with respect to the nonparanormal case. Indeed, when $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$ a missing edge implies conditional independence. On the other hand, when $\boldsymbol{X}\sim M\mathcal{E}(\Sigma^{0},\xi,f)$ a missing edge means conditional linear independence.
The meta-elliptical family contains many useful distributions among which asymmetric distributions. However, the estimate of the correlation matrix does not directly take into account the skewness parameters. We extend the meta-elliptical family giving a direct contribution of the skewness parameters for the conditional independence relations, and thus considering the $\mathcal{E}$ family of distributions.
2.4 $\mathcal{E}$ distributions
We recall the family of $\operatorname{SU\mathcal{E}}$ distributions described in Arellano-Valle and Genton, 2010b . If a random vector has a $\operatorname{SU\mathcal{E}}$ distribution, we write $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,q}(\mu,\Sigma,\Lambda,h^{(p+q)},\tau,\Gamma)$, where $h^{(p+q)}$ is the density generator. We focus on two specifications of the family which are: the unified skew-$t$ ($\operatorname{SUT}$) distribution (Arellano-Valle and Genton, 2010a ), and the unified skew-Normal ($\operatorname{SUN}$) distribution (Arellano-Valle and Azzalini, (2006)). Table 1 shows how the $\operatorname{SU\mathcal{E}}$ distribution reduces into $\operatorname{SUN}$ and $\operatorname{SUT}$ distributions. Formally, we define the $\operatorname{SU\mathcal{E}}$ distribtuion.
Definition 2.4.
(Arellano-Valle and Genton, 2010b ) A continuous p-dimensional random vector $X$ has a multivariate unified skew-elliptical ($\mathcal{E}$) distribution, denoted by $\boldsymbol{X}\sim SU\mathcal{E}_{p,q}(\mu,\Sigma,\Lambda,h^{(p+q)},\tau,\Gamma)$, if its density function at $x\in\mathbb{R}^{p}$ is
$f_{X}(x)=\frac{1}{F_{q}\left(\tau;\Gamma+\Lambda\bar{\Sigma}\Lambda^{T},h^{(q)}\right)}f_{p}\left(y;\mu,\Sigma,h^{(p)}\right)F_{q}\left(\Lambda z+\tau;\Gamma,h_{Q(z)}^{(q)}\right),$ (4)
where $w=\sigma^{-1}(y-\mu),Q(w)=w^{\top}(\Sigma^{0})^{-1}w,\Lambda$ is a $q\times p$ real matrix controlling shape, $\tau\in\mathbb{R}^{q}$ is the extension parameter, $\Gamma$ is a $q\times q$ positive definite correlation matrix, $f_{p}\left(y;\mu,\Sigma,h^{(p)}\right)=$ $|\Sigma|^{-1/2}h^{(p)}(Q(w))$ denotes the density function of an elliptically contoured distribution with location $\mu\in\mathbb{R}^{p}$, positive definite $p\times p$ dispersion matrix $\Sigma$, with $p\times p$ scale and correlation matrices $\sigma=\operatorname{diag}(\Sigma)^{1/2}$ and $\Sigma^{0}=\sigma^{-1}\Sigma\sigma^{-1}$, respectively, and density generator $h^{(p)},F_{r}\left(y;\Sigma,h^{(r)}\right)$ denotes the $r$-dimensional centered elliptical cumulative distribution function with $r\times r$ dispersion matrix $\Sigma$ and density generator $h^{(q)}$, and $h_{Q(w)}^{(q)}(u)=h^{(p+q)}(u+Q(w))/h^{(p)}(Q(w))$.
$\operatorname{SU\mathcal{E}}$ $h^{(p)}$
$\operatorname{SUN}$ $h^{(p)}=\phi^{(p)}$$(u)$where: $\phi^{(p)}(u)=\frac{1}{(2\pi)^{p/2}}\exp\left(-\frac{u^{2}}{2}\right)$
$\operatorname{SUT}$ $h^{(p)}=t^{(p)}(u,\nu)$ where: $t^{(p)}(u,\nu)=\frac{\Gamma\left(\frac{\nu+p}{2}\right)}{\Gamma\left(\frac{\nu}{2}\right)\left(\pi a\right)^{p/2}}\left(1+\frac{u}{\nu}\right)^{-\frac{\nu+p}{2}}$
Table 1: Subclasses of skew-elliptical distribution based on the density generator $h^{(p)}$.
The relation between the $\operatorname{SUN}$ and the $\operatorname{SU\mathcal{E}}$ is defined through the following stochastic representation.
Proposition 2.5.
(Arellano-Valle and Genton, 2010b ) Given $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,q}(\mu,\Sigma,\Lambda,h^{(p+q)},\tau,\Gamma)$, when the extension parameter $\tau=0$, we can write $\boldsymbol{X}$ as a mixture of a $\operatorname{SUN}$ distribution: $\boldsymbol{X}=\mu+\sigma V^{-1/2}\boldsymbol{Z}_{0}$, where $V_{0}$ is the cumulative distribution function of a uni-dimensional random variable, called mixture variable, such that: $V_{0}(0)=0$, and $V_{0}$ is independent of $\boldsymbol{Z}_{0}\sim SUN_{p,q}(0,\Sigma^{0},\Lambda,0,\Gamma)$.
The $\operatorname{SUN}$ distribution results in a different parameterization of the closed skew-Normal (CSN) distribution in Domìnguez-Molina et al., (2003). The CSN distribution corresponds to the case in which the cumulative distribution functions in Definition 2.4 are the Gaussian cumulative distribution functions. Therefore, the $\operatorname{SU\mathcal{E}}$ distribution can be expressed in terms of the $\operatorname{CSN}$ distribution. This is useful for deriving the moments and for sampling random numbers from the $\operatorname{SU\mathcal{E}}$ distribution. The $\operatorname{CSN}$ distribution has the following stochastic representation as a combination of a Normal distribution and a half-Normal distribution.
Definition 2.6.
(Domìnguez-Molina et al., (2003)) Let $\boldsymbol{Y}\sim\operatorname{CSN}_{p,q}(\mu,\Sigma,\Gamma,0,\Delta)$, the random vector $\boldsymbol{Y}$ admits the following stochastic representation:
[…]
where: $\boldsymbol{U}\sim\operatorname{N}_{p}(0,I_{p})$ and $\boldsymbol{V}\sim\operatorname{HN}_{q}(0,\Delta+\Gamma\Sigma\Gamma)$, the Normal and the half-Normal distribution, respectively.
The stochastic representation in (5) allows us to estimate an undirected decomposable graph.
Definition 2.7.
(Zareifard et al., (2016)) Consider an undirected decomposable graph $G=(V,E)$, a random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ is called a skew decomposable graphical model concerning the graph G with mean $\mu$, the precision matrix $\Omega$ and the skewness parameter $\alpha\in\mathbb{R}^{p}$ if its density is $\operatorname{CSN}_{p,p}(\mu,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ or equivalently
$f_{\boldsymbol{X}}(x)=2^{p}\phi_{p}\left(x;\mu,\Omega^{-1}\right)\Phi_{p}\left(D_{\alpha}L(x-\mu);0,D^{-1}_{\kappa}\right),$ (6)
where $L$ and $D_{\kappa}$ are the matrix resulting from the modified Cholesky of $\Omega=LD_{\kappa}L^{\top}$, and zero in $L$ matrix determines the conditional independence relations between the $X_{1},\ldots,X_{p}$ variables, further, when the matrix $D_{\alpha}=0$, the random vector $\boldsymbol{X}\sim\operatorname{CSN}_{p,p}(\mu,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ reduces to $\boldsymbol{X}\sim\operatorname{N}_{p}(\mu,\Omega^{-1})$.
We introduce an estimator for skew-Gaussian graphical models defined in Zareifard et al., (2016) that extends the SKEPTIC including a skewness component to accommodate non-symmetric data.
3 The elliptical skew-(S)KEPTIC estimator
Definition 3.1.
(Meta skew-elliptical distribution) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has meta skew-elliptical distribution, denoted by $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},0,\Lambda;f)$, if there exists a set of monotone univariate functions $f_{1},\ldots,f_{p}$ and a nonnegative random variable $\xi$, such that
$(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\operatorname{SU}\mathcal{E}_{p,p}(0,\Sigma^{0},\Lambda,0,\Gamma),$ (7)
where $\Lambda$ is a matrix containing the skewness terms.
Proposition 3.2.
When $\Lambda=0$, then the meta skew-elliptical distribution reduces to the meta-elliptical distribution.
Proposition 3.3.
When $\Lambda=0$ and $h^{(p)}=\phi^{(p)}(u)$, then the meta skew-elliptical distribution reduces to the nonparanornmal distribution.
We exploit the stochastic representation of $\operatorname{SU\mathcal{E}}$ distribution through a CSN distribution (Proposition 2.5) to define the skew-elliptical graphical model. That is, we propose a nonparametric rank-based estimator, which we call elliptical skew-(S)KEPTIC estimator, based on the stochastic representation (5), according to the parametrization of $\operatorname{CSN}$ distribution in (6):
$\boldsymbol{Z}\stackrel{{\scriptstyle p}}{{=}}\boldsymbol{\mu}+L^{-1}D_{\kappa}^{-\frac{1}{2}}D_{\alpha}\left(I+D_{\alpha}^{2}\right)^{-\frac{1}{2}}\boldsymbol{U}+L^{-1}D_{\kappa}^{-\frac{1}{2}}\left(I+D_{\alpha}^{2}\right)^{-\frac{1}{2}}\boldsymbol{V},$ (8)
where: $\boldsymbol{U}\sim N_{p}(0,I_{p})$ and $\boldsymbol{V}\sim HN_{p}(0,I_{p})$, where $N_{p}(0,I_{p})$ is a $p$-dimensional Normal distribution, and $HN_{p}(0,I_{p})$ is a $p$-dimensional half-Normal distribution. The distributions of $\boldsymbol{U}$ and $\boldsymbol{V}$ are independent. The first two moments of $\boldsymbol{Z}$ are:
$\mathbb{E}(\boldsymbol{Z})$ $=\mu+L^{\top}D_{\kappa}^{-\frac{1}{2}}D,$
[…]
where $D=\sqrt{\frac{2}{\pi}}D_{\alpha}(I_{p}+D^{2}_{\alpha})^{-\frac{1}{2}}$.
Lemma 3.4.
[…]
$\boldsymbol{X}\sim\operatorname{SU}\mathcal{E}_{p,p}(\xi,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$ reduces to a $\operatorname{SUT}$ distribution: $\boldsymbol{X}\sim\operatorname{SUT}_{p,p}(\xi,\Sigma,\Lambda,\nu,0,\Gamma)$.
Theorem 3.5 (elliptical skew-KEPTIC estimator).
Assuming $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},0,\Lambda;f)$, such that $\boldsymbol{Z}=f(\boldsymbol{X})$ denotes a generic $\operatorname{SU\mathcal{E}}$ distribution: $\boldsymbol{Z}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Sigma^{0},\Lambda,h^{(2p)},0,\Gamma)$. The elliptical skew-KEPTIC estimator is:
${S}^{\tau,\alpha}_{ij}=\begin{cases}2\operatorname{sin}\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)&i\neq j\\ 1&i=j\end{cases}.$
As discussed in Hult and Lindskog, (2002), Spearman’s rho is not invariant in the class of elliptical distributions, and its relation with linear correlation coefficient in Lemma 2.1 holds if and only if the distribution is nonparanonrmal.
Proof.
We consider $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},0,\Lambda;f)$, where $f(\boldsymbol{X})\sim V_{0}^{-1/2}Z_{0}$, and $Z_{0}\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ where $D_{\kappa}=I_{p}$, and $V_{0}=1$. We compute $\mathbb{C}orr(X)$ using Kendall’s tau statistic, $\Sigma^{0}_{\tau}=\left[\tau_{ij}\right]$, and we decompose $\Sigma^{0}_{\tau}=L^{\top}D_{\kappa}L$, aligning it $\Sigma^{0}_{\tau}$ to the assumption of $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},0,\Lambda;f)$ exploiting (9).
[…]
$=(L^{\top}ML)^{-1}\quad\text{where:}\quad M=I-D^{2},$
$=\frac{Adj(L^{\top}ML)}{det(L^{\top}ML)},$
$=\frac{Adj(L^{\top}ML)}{det(L)^{2}det(M)}.$
Where $Adj$ is the adjoint matrix, and $det$ is the determinant of the matrix. The matrices $I$ and $D^{2}$ are both diagonal matrices, and the matrix $M$ is. Thus, we have a new decomposition of the correlation matrix in which the skewness parameter $\alpha$, belonging to the matrix $D_{\alpha}$, contributes to the determination of the linear relations. The matrix $(L^{\top}ML)$ is:
$\left(\begin{array}[]{ccccc}m_{11}&\ell_{12}m_{11}&\ell_{13}m_{11}&\ldots&\ell_{1p}m_{11}\\ \ell_{21}m_{11}&\sum^{p-(p-2)}_{k=1}\ell_{k2}\ell_{2k}m_{kk}&\sum_{k=1}^{p-(p-2)}\ell_{k2}\ell_{2k}m_{kk}&\ldots&\sum_{k=1}^{p-(p-2)}\ell_{k2}\ell_{pk}m_{kk}\\ \ell_{31}m_{11}&\sum_{k=1}^{p-(p-2)}\ell_{k2}\ell_{3k}m_{kk}&\sum_{k=1}^{p-(p-3)}\ell_{k3}\ell_{3k}m_{kk}&\ldots&\sum_{k=1}^{p-(p-3)}\ell_{k3}\ell_{pk}m_{kk}\\ \vdots&\ldots&\ldots&\ddots&\vdots\\ \ell_{p1}m_{11}&\sum_{k=1}^{p-(p-2)}\ell_{k2}\ell_{pk}m_{kk}&\sum_{k=1}^{p-(p-3)}\ell_{k3}\ell_{pk}m_{kk}&\ldots&\sum_{k=1}^{p}\ell_{kp}\ell_{pk}m_{kk}\end{array}\right)$
where the generic $m_{k,k}$ term is:
$m_{kk}=I_{kk}-D^{2}_{kk}=1-\frac{2}{\pi}\frac{\alpha_{kk}^{2}}{1+\alpha_{kk}^{2}}.$
Let $A_{{\tau}}$ denote the final correlation matrix of $X$ with Kendall’s tau statistic. We recall that the stochastic representation (8) of the $\operatorname{CSN}$ is a sum of two comonotic copula functions that have the invariance property, so the SKEPTIC estimator is the same for both the distributions, thus the elliptical skew-KEPTIC estimator is the sum of two SKEPTIC estimators and an estimated matrix, $A_{\tau}$, from $f(\boldsymbol{X})\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$.
When $V_{0}\sim Gamma(\nu/2,\nu/2)$, we have that $f(\boldsymbol{X})\sim\operatorname{SUT}_{p,p}(\xi,\Sigma,\Lambda,\nu,0,\Gamma)$ or equivalently: $f(\boldsymbol{X})\sim\frac{Z_{0}}{\sqrt{V_{0}}}$. Since the positive support of Gamma random variable, and the fact that Kendall’s tau is invariant under monotone transformations, we state that the skew-KEPTIC estimator maintains the same ordinal measures of dependencies of CSN distribution even when $f(\boldsymbol{X})\sim\operatorname{SUT}_{p,p}(0,\Sigma^{0},\Lambda,\nu,0,\Gamma)$, where $\Gamma=I_{p}$. ∎
Lemma 3.6.
(skew-Normal SKEPTIC estimator) Assuming $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},0,\Lambda;f)$, such that $\boldsymbol{Z}=f(\boldsymbol{X})$ denotes a generic $\operatorname{SU\mathcal{E}}$ distribution, when $\boldsymbol{Z}\sim\operatorname{SU}\mathcal{E}_{p,p}(0,\Sigma^{0},\Lambda,h^{(2p)},0,\Gamma)$ reduces to a $\operatorname{CSN}$ distribution, the skew-Normal SKEPTIC estimator is:
${S}^{\rho,\alpha}_{ij}=\begin{cases}4\sin\left(\frac{\pi}{6}A_{{\rho}_{ij}}\right)&\text{if }i\neq j\\ 1&\text{if }i=j\end{cases},\quad\quad{S}^{\tau,\alpha}_{ij}=\begin{cases}2\operatorname{sin}\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)&i\neq j\\ 1&i=j\end{cases}.$
Proof.
We recall Lemma 2.1, where for the Normal distribution, and for $i\neq j$, we have: ${S}_{ij}=2\sin\left(\frac{\pi}{6}{\rho}_{ij}\right)=\sin\left(\frac{\pi}{2}{\tau}_{ij}\right)$. Let $A_{\rho_{ij}}$ denote the final correlation matrix of $\boldsymbol{X}$ with Spearman’s rho statistic, as in the Normal distribution, we have: ${S}^{\rho,\alpha}_{ij}\equiv 4\sin\left(\frac{\pi}{6}A_{{\rho}_{ij}}\right)=2\sin\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)$. ∎
We generally refer to the estimators in Theorem 3.5 and Lemma 3.6 as elliptical skew-(S)KEPTIC estimators. The specification of the inclusion of the Spearman’s rho will be clear in the application. To obtain the correlation matrix using the elliptical skew-(S)KEPTIC estimator, we need to estimate the parameters of the $\operatorname{CSN}$ distribution. The estimation procedure is based on the EM algorithm of Abe et al., (2021) [^1] The estimated parameters come from the skew-Normal distribution of Azzalini and Capitanio, (1999), with an adequate parametrization explained in Arellano-Valle and Azzalini, (2006) we conform the parameters to a CSN distribution..
3.1 Estimating Marginal Distributions
We find estimators $\{f_{j}\}_{j=1}^{p}$ such that the transformed data are distributed according to a $\operatorname{CSN}$. Let $\Psi$ denote the cumulative distribution function of the $\operatorname{CSN}$ distribution, then the estimators are:
$\hat{f}_{csn,j}(x_{ij})=\Psi^{-1}(\hat{r}_{ij}),\quad\text{where: }\hat{r}_{ij}=\frac{r_{ij}}{n+1}.$ (11)
The functions $\{f_{csn,j}(x_{ij})\}_{j=1}^{p}$ are monotone and differentiable, and Gupta et al., (2004) show that $\operatorname{CSN}$ distribution is closed under conditioning therefore, according to Sklar’s theorem, (6) is a $\operatorname{CSN}$ copula.
3.2 Methods for the Estimation of the Precision Matrix and the Undirected Graph
Upon completing the estimation of the correlation matrix, we follow Liu et al., 2012a and estimate the precision matrix capturing the interrelationships among variables within an undirected graph.
3.2.1 The Dantzig selector
The Dantzig selector estimator (Yuan, , 2010) exploits the relation between the coefficients $\theta$ stemming from a multivariate linear regression and the elements of the precision matrix. Sparse precision matrix implies sparse regression coefficients estimation that relies on an algorithm that depends on the tuning parameter $\delta$. Dantzig procedure involves two steps:
  • Estimation: for $j=1,\ldots,p$, calculate
    $\hat{\theta}^{j}$ $=\text{arg}\min_{\theta\in\mathbb{R}^{p-1}}\|\theta\|_{1}\quad\text{subject to }\quad\|\hat{S}_{\backslash j,j}-\hat{S}_{\backslash j,\backslash j}\theta\|_{\infty}\leq\delta,$ (12)
    $\hat{\Omega}_{jj}$ $=\left[1-2(\hat{\theta}^{j})\hat{S}_{\backslash j,j}+(\hat{\theta}^{j})^{\top}\hat{S}_{\backslash j,\backslash j}\hat{\theta}^{j}\right]\quad\text{and }\quad\hat{\Omega}_{\backslash j.j}=-\hat{\Omega}_{\backslash jj}\hat{\omega}^{j}.$
  • Symmetrization:
    $\hat{\Omega}^{gDS}=\text{arg}\min_{\Omega=\Omega^{T}}\|\Omega-\hat{\Omega}\|_{1}.$ (13)
3.2.2 The Glasso Algorithm
Glasso algorithm is developed by Friedman et al., (2008) and it consists in adding a $L_{1}$-norm penalty term in the maximum likelihood estimation of $\Omega$ as follows:
$\Omega^{glasso}=arg\max_{\Omega\succ 0}\left\{\text{log}\ \text{det}\Omega-\text{tr}(S\Omega)-\lambda_{G}\|\Omega\|_{1}\right\}.$ (14)
The penalty term, $\lambda_{G}$, is the parameter that controls the sparsity of the estimated matrix, i.e., the number of zeros in the precision matrix. Banerjee et al., (2008) shows that the shrinking procedure based on the tuning parameter $\lambda_{G}$ is directly proportional to the input covariance matrix. This means that greater values of $\lambda_{G}$ imply a greater presence of zeros in the estimated $\Omega$, therefore a minor number of interconnections are unveiled.
3.2.3 The CLIME Algorithm
CLIME stands for constrained $\ell_{1}$-minimization for inverse matrix estimation. This algorithm, proposed by Cai et al., (2011), is very attractive for high-dimensional data since the estimator is obtained by combining vector solutions solving a linear program. Let $\Omega$ denote a $p\times p$ precision matrix and $\{\hat{\Omega}_{1}\}$ be the solution set of the optimization problem:
$\text{min}\|\Omega\|_{1}\quad\text{subject to:}$: (15)
$\|\Sigma_{n}\Omega-I\|_{\infty}\leq\lambda_{C},$
where $\lambda_{C}$ is the tuning parameter. The final CLIME estimator of $\Omega_{0}$ is defined by the symmetrization of $\hat{\Omega}_{1}$:
$\hat{\Omega}=(\hat{\omega}_{ij}),\quad\text{where:}$: (16)
$\hat{\omega}_{ij}=\hat{\omega}_{ji}=\hat{\omega}^{1}_{ij}I\{|\hat{\omega}^{1}_{ij}\leq|\hat{\omega}^{1}_{ij}|\}+\hat{\omega}^{1}_{ji}I\{|\hat{\omega}^{1}_{ji}\leq|\hat{\omega}^{1}_{ji}|\}.$
The CLIME estimator has the property that the optimization problem defined in (15) can be decomposed into $p$ optimization problems defined as:
$\min\|\boldsymbol{\beta}\|_{1}\quad\text{subject to}\quad\|\Sigma\boldsymbol{\beta}-\boldsymbol{e}_{i}\|_{\infty}\leq\lambda_{n},$ (17)
where $\boldsymbol{e}_{i}$ is defined as a standard unit vector in $\mathbb{R}^{p}$ with 1 in the $i$-th coordinate and 0 in the all other coordinates and $\boldsymbol{\beta}$ is a vector in $\mathbb{R}^{p}$. Solving (17) $p$ times, we obtain the CLIME estimator: $\{\hat{\Omega}_{1}\}=\{\hat{B}\}:=\{(\hat{\boldsymbol{\beta}}_{1},...,\hat{\boldsymbol{\beta}}_{p})\}$.
4 Theoretical properties
We analyze the theoretical properties of the elliptical skew-(S)KEPTIC estimator. The results suggest that our estimator is reliable when dealing with highly skewed data. As Liu et al., 2012a , our results show an exponential concentration rate in $\|\cdot\|_{\max}$. Moreover, we prove optimal bounds for shrinkage procedures. We show the results for the Dantzig selector, however, they can be easily extended to the CLIME and the Glasso procedures.
Theorem 4.1.
Let $x_{1},...,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim\operatorname{MS\mathcal{E}}_{p}(\Sigma^{0},\xi,\Lambda;f)$, restricted to the case $h^{(p)}=\phi^{(p)}(u)$, and let $\hat{S}^{\rho,\alpha}_{ij}=4\sin(\frac{\pi}{6}A_{\hat{\rho,}_{ij}})\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ denote the elliptical skew-SKEPTIC, then with probability at least $1-1/p^{2}$ we have:
$\|\hat{S}^{\rho,\alpha}_{ij}-\Sigma^{0}_{ij}\|_{\max}\leq 16\pi\sqrt{\frac{logp}{n}}.$ (18)
Theorem 4.2.
Let $x_{1},...,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim\operatorname{MS\mathcal{E}}_{p}(\Sigma^{0},\xi,\Lambda;f)$ and let $\hat{S}^{\tau,\alpha}_{ij}=2\sin(\frac{\pi}{2}A_{\hat{\tau}_{ij}})\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ denote the elliptical skew-(S)KEPTIC estimator, then with probability at least $1-1/p^{2}$ we have:
$\|\hat{S}^{\tau,\alpha}_{ij}-\Sigma^{0}_{ij}\|_{\max}\leq 4\pi\sqrt{\frac{logp}{n}}.$ (19)
Proofs of Theorems 4.1 and 4.2 are in Appendix A. The elliptical skew-(S)KEPTIC estimator achieves convergence in high-dimension, thus it is possible to define the following result.
Theorem 4.3.
When we substitute the estimated matrices $\hat{S}^{\rho,\alpha}$ and $\hat{S}^{\tau,\alpha}$ into the parametric graphical lasso (or the CLIME or graphical Dantzig selector), the estimators provide a slower convergence than the Gaussian case ensuring consistency and graph recovery as a Gaussian model.
Proof.
The proof is based on the observation that the sample correlation matrix $\hat{S}$ is a sufficient statistic for all three shrinkage procedures: the graphical lasso, graphical Dantzig selector, and CLIME. As Yuan, (2010) points out, a sufficient condition to ensure that an estimator is consistent is if there exists some constant $c$, such that $\mathbb{P}\left(\|\hat{S}-\Sigma^{0}\|_{\max}>c\sqrt{\frac{\log p}{n}}\right)\leq 1-\frac{1}{p}$, which can be replaced by (18) and (19) in Theorems 4.1 and 4.2. ∎
The high-dimensional convergence of the elliptical skew-(S)KEPTIC estimator also means that the precision matrix is bounded. Similar results are found by Cai et al., (2011) and Yuan, (2010). In particular, let $\hat{\Omega}^{mse-s}$ denote the inverse correlation matrix using the elliptical skew-(S)KEPTIC estimator and consider the Dantzig selector as the shrinkage procedure to find $\hat{\Omega}$. We introduce a class of inverse matrix defined as $\mathcal{M}_{1}(\kappa,\tau,M):=\{\Omega:\Omega\succ 0,\text{diag}(\Omega^{-1})=1,\|\Omega\|_{1}\leq\kappa,\frac{1}{\tau}\leq\lambda_{\min}(\Omega)\leq\lambda_{\max}(\Omega)\leq\tau,\deg(\Omega)\leq M\}$, where $\kappa,\tau>1$. Given Theorem 4.3, we obtain the following result:
Theorem 4.4.
For $1\leq q<\infty$, there exists a constant $c$ that depends on $\kappa$, $\tau$, $\lambda_{\min}(\Omega)$ and $\lambda_{\max}(\Omega)$, such that
$\sup_{\Omega\in\mathcal{M}_{1}(\kappa,\tau,M)}\|\hat{\Omega}^{mse-s}-\Omega\|_{q}=O_{P}\left(M\sqrt{\frac{\log p}{n}}\right).$ (20)
provided that $\lim_{n\rightarrow\infty}\frac{n}{M^{2}\log d}=\infty$ and $\delta=c\sqrt{\frac{\log d}{n}}$, for sufficiently large $c$, where $\delta$ is the parameter in (12).
Proof.
This result follows Yuan, (2010), which proves that the difference between the estimated precision matrix and the theoretical precision matrix is bounded by $\|\hat{S}-\Sigma^{0}\|_{\max}$, where $\hat{S}$ is the estimated correlation matrix and $\Sigma^{0}$ is the theoretical correlation matrix. In our case, to prove the bounds for the estimated correlation matrix we just replace $\hat{S}$ with $\hat{S}^{\rho,\alpha}$ and $\hat{S}^{\tau,\alpha}$ where the bounds are proved in Theorem 4.1 and 4.2. ∎
5 Experiments
We present the empirical performance of the elliptical skew-(S)KEPTIC methods on both synthetic and real datasets, comparing the proposed method throughout this paper with the SKEPTIC estimator in Liu et al., 2012a . In particular, we consider the following methods for the estimation of graphical models. We generate data from the Normal distribution, the CSN distribution, and the SUT distribution. For each generation, we compute the SKEPTIC and the SKEW-(S)KEPTIC estimator and conduct numerical simulations on synthetic data to show the ability of graph recovery of the proposed estimators. In conclusion, we present an empirical analysis where the methods are applied to the log-returns of the S&P500 index.
5.1 Numerical Simulations
The numerical simulations follow these steps. Using the R package huge (Jiang et al., (2021)), we generate the precision matrix $\Omega=(\Sigma^{0})^{-1}$ using the ”random” structure, in which we randomly generate pairs of conditional dependencies in the off-diagonal elements. We sample $n$ observations from $p$-variate Normal, $\operatorname{CSN}$ and $\operatorname{SUT}$ distribution. We set the location parameter equal to zero for the last two, and $\Sigma^{0}=\Omega^{-1}$. We consider the following cases: $p\in\{100,100,200\}$, and $n\in\{200,100,100\}$. The cases $n=100,p=100$, and $n=100,p=200$ consider the high-dimensional scenarios. Regarding the $\operatorname{CSN}$ parameters, we set $\Delta=1$ and the skewness matrix is generated considering the off-diagonal elements equal to zero, whereas the diagonal matrix is generated from $\operatorname{Uniform}(-0.5,0.5)$. We generate the $\operatorname{SUT}$ distribution with degrees of freedom coming from $\operatorname{Uniform}(2,10)$ [^2] The range is determined by estimating the degrees of freedom of the data on stock returns. . The estimators are then plugged into the ”Glasso” algorithm in (14) to obtain $\Omega$.
To evaluate the accuracy of the methods, we use the receiver operating curve ($\operatorname{ROC}$). We define the false positive rate and true positive rate for a fixed set of tuning parameters $\lambda_{G}$. Let $\hat{G}^{\lambda}_{G}=(V,\hat{E}^{\lambda_{G}})$ be an estimated graph using the proposed procedure and a given value of $\lambda_{G}$. We define the number of false positives for a given $\lambda_{G}$, $\operatorname{FP}(\lambda_{G})$, as the number of edges in $\hat{E}^{\lambda_{G}}$ but not in $E$ and the number of false negatives, $\operatorname{FN}(\lambda_{G})$, as the number of edges in $E$ but not in $\hat{E}^{\lambda_{G}}$. We consider the rate of the number of false negative and false positive:
$\operatorname{FNR}(\lambda_{G}):=\frac{\operatorname{FN}(\lambda_{G})}{|E|},\quad\text{and}\quad\operatorname{FPR}(\lambda_{G}):=\operatorname{FP}(\lambda_{G})/\left[\left(\begin{array}[]{l}d\\ 2\end{array}\right)-|E|\right].$ (21)
To demonstrate the comprehensive performance of the analyzed methods across the entire datasets, we define the ROC curve
$(\operatorname{FNR}(\lambda_{G}),1-\operatorname{FPR}(\lambda_{G})).$ (22)
Graphically, we show averaged ROC curves for all the considered scenarios over 100 trials, plotted using the coordinates in (22). Fig. 1 shows numerical simulations of the average ROC curves of the elliptical skew-(S)KEPTIC estimators and the SKEPTIC estimators, which we employ to estimate $\Omega$ using the Glasso algorithm. The estimators exhibit similar performance, consequently yielding an optimal graph recovery in the simulation results, even when accounting for skewness and kurtosis components in the data, thus giving reliability to the elliptical skew-(S)KEPTIC estimators.
Table 2 presents the quantitative comparisons, showing that regardless of the data generation process or the estimator used, the structure of $\Omega$ can be effectively recovered. Incorporating a skewness component and accounting for heavier tails than those in the NPN distribution do not hinder the recovery of $\Omega$’s structure.
Table 2: Quantitative comparisons of the 4 estimators using different distributions to generate data, and compute the $\Omega$ matrix. The graphs are estimated using the Glasso algorithm. The values are in percentage.
SKEPTIC ($\hat{\rho}$) SKEPTIC ($\hat{\tau}$) SKEW-SKEPTIC ($\hat{\rho}$) SKEW-SKEPTIC ($\hat{\tau}$) $n$ $p$ FPR FNR FPR FNR FPR FNR FPR FNR N DATA 200 100 11.11 74.3 11.2 74.3 11.8 74.7 11.4 74.7 s.e (1.3) (3.5) (1.3) (3.5) (1.4) (3.5) (1.3) (3.5) 100 100 15.5 66.9 15.4 67.0 16.1 67.1 15.7 67.2 s.e (1.6) (2.4) (1.0) (2.4) (1.1) (2.4) (1.1) (2.4) 100 200 13.8 64.0 14.0 64.2 14.9 64.0 14.5 64.1 s.e (1.0) (2.4) (1.0) (2.4) (1.1) (2.4) (1.1) (2.4) CSN DATA 200 100 11.5 73.9 11.5 74.0 12.1 74.1 11.6 73.7 s.e (1.4) (3.5) (1.4) (3.5) (3.5) (3.5) (1.4) (3.5) 100 100 15.6 67.2 15.8 67.3 16.0 65.8 15.3 65.1 s.e (1.6) (3.4) (1.6) (3.4) (1.7) (3.5) (1.6) (3.5) 100 200 13.9 64.0 14.1 64.1 14.3 62.1 13.6 61.6 s.e (1.0) (2.4) (1.0) (2.4) (1.1) (2.5) (1.0) (2.5) SUT DATA 200 100 12.6 72.6 12.8 72.8 - - 14.9 61.0 s.e (1.5) (3.5) (1.5) (3.5) (-) (-) (1.7) (3.3) 100 100 16.8 64.4 17.0 64.5 - - 17.1 64.5 s.e (1.7) (3.3) (1.7) (3.3) (-) (-) (1.7) (3.3) 100 200 14.5 60.7 14.5 61.0 - - 14.9 61.0 s.e (1.0) (2.4) (1.0) (2.4) (-) (-) (1.1) (2.4)
Figure 1: Comparison of average ROC curves, using the Glasso algorithm, of the elliptical skew-(S)KEPTIC estimators with SKEPTIC estimators. Results are determined over 100 trials. We consider, by column, three different data generations. And, by row, three different dataset sizes.
5.2 Empirical Analysis: S&P500 Log-Returns
In the empirical analysis, we study the conditional dependence relations on daily returns of S&P500500500500 companies in the period from $04/01/2016$ to $31/12/2022$ downloaded from Yahoo! Finance, using the proposed elliptical skew-(S)KEPTIC estimator and comparing it with the SKEPTIC. After cleaning the data, the dataset consists of 454 stocks, each counting 1762 observations. The log-returns of the stocks in the S&P500 index are empirically skewed and present heavier tails than the Normal distribution (Cont, (2001)).
Figure 2: Illustration of leptokurtic distributions of daily log-returns, with a relevant skewness component. The stock data are from 04/01/2016 to 31/12/2022.
As a motivating example, we benchmark the analysis of Han and Liu, (2014), reporting the asymmetry of log-returns in Figure 2, where we plot histograms of the returns of three stocks in our sample ”eBay Inc.”, ”Coca-Cola” and ”Wells Fargo” belonging to three different sectors, respectively, Consumer Discretionary, Consumer Staples, and Financials. We appreciate the departure from normality reflected in leptokurtic distribution with a relevant skewness component. Moreover, we investigate the Gaussian assumption of the asset log-returns computing normality tests with different significance levels at 5% and 1%. Table 3 indicates a clear and complete deviation from normality in the stock data. This could be attributed to the presence of exceptional events thus a huge presence of outlier value during the analyzed period, which lead to a departure from the assumption of Normal distribution for log-returns.
Table 3: Normality tests S&P500 log-returns data. This table illustrates the number of stocks that reject the null hypothesis of normality at 1% and 5% levels.
Significance Level Kolmogorov-Smirnov Shapiro-Wilk Lilliefors
1% 454 454 454
5% 454 454 454
To estimate the graph, we determine the tuning parameter $\lambda$ using Extended Bayesian Information Criteria (EBIC) developed in Chen and Chen, (2008):
$\operatorname{EBIC}=-2L+Elog(n)+4\gamma Elog(p).$ (23)
The hyperparameter $\gamma\in\left[0,1\right]$, with greater values of $\gamma$ implying a stronger penalization, $L$ denotes the log-likelihood, $n$ the sample size, $E$ the number of non-zero edges and $p$ the number of nodes. We set $\gamma=1$. The tuning parameter is selected as $\lambda_{G}=0.54$ since it minimizes the scores of the first-order differences in EBIC.
For estimating the undirected graph conforming to the SUT distribution, using the sut Kendall estimator, we need to estimate the degrees of freedom of the log-returns distributions. We follow Arellano-Valle and Genton, 2010a using the maximum likelihood estimation method. In likelihood methods, the underlying assumption is that the observations are independent, therefore we apply the best ARIMA model to the series using the R function auto.arima to filter our returns into a white noise series, and we apply the likelihood method to the filtered series to estimate the degrees of freedom using the command mst.mle from the $R$ package sn developed by Azzalini, (2023).
Figure 3 shows the estimated graphs using npn Kendall, csn Kendall, and sut Kendall estimators.
Figure 3: The resulting graphs using the (a) csn Kendall estimator, (b) sut Kendall estimator, and (c) npn Kendall estimator for S&P 500 stock data from $04/01/2016$ to $31/12/2022$.
The 454 stocks are categorized into 11 Global Industry Classification Standard (GICS) sectors, including Consumer Discretionary (51 stocks), Consumer Staples (33 stocks), Energy (21 stocks), Financials (67 stocks), Health Care (56 stocks), Industrials (69 stocks), Information Technology (59 stocks) Telecommunications Services (17 stocks), Materials (25 stocks), Real Estate (28 stocks), and Utilities (28 stocks). The stocks that belong to the same sector tend to cluster. Nevertheless, csn Kendall and sut Kendall show in better way the grouping of the stocks, thus improving the interpretability of the visualization.
Stocks at the periphery of the network are uncorrelated to the others, during the considered period, thus they are uncorrelated with the movement of the market. Peralta and Zareei, (2016) shows that those stocks have a greater weight in constructing a portfolio using the Markowitz approach. Therefore, in high-dimension, a network approach can be valuable for identifying stocks with the least correlation to the market.
Table 4 shows that introducing the skewness and kurtosis components results in a sparser network estimation compared to the Normal case, thus the elliptical skew-(S)KEPTIC estimator improves the interpretability of the graph.
Table 4: Summary statistics of the estimated networks using Kendall’s tau statistic for the correlation matrix estimation. SK = SKEPTIC, SNK = elliptical skew-SKEPTIC, STK = elliptical skew-KEPTIC. DIM = dimension of the dataset. The notation ”$>$” indicates that one method has a greater number of edges compared to the other.
Network  Edges No. Edges diff.
DIM SK STK SNK SK$>$SNK SK$>$STK SNK$>$STK
S&P 500 454 399 303 252 96 147 51
6 Concluding remarks
We propose a rank-based correlation matrix estimator, named the elliptical skew-(S)KEPTIC estimator which extends to the class of the skew elliptical distribution in estimating the high-dimensional graphical models. The elliptical skew-(S)KEPTIC estimator results for Kendall’s tau in the class of skew-elliptical distribution, and it also encompasses Spearman’s rho when the class of distribution reduces to the skew-Normal. We expand the SKEPTIC estimator in Liu et al., 2012a to the class of semiparametric meta skew-elliptical distribution introduced throughout the paper.
Our analysis is not asymptotic, and the estimator is defined in a way in which we do not explicitly estimate the marginal transformation of the data. Theoretically, the slight reduction in the robustness of the estimator is compensated by the extension of the class of distributions over which we estimate the graphical model.
We illustrate a method for modeling data that conforms to a $\operatorname{MS}\mathcal{E}$ distribution by treating the data generation process as a mixture of Normal and half-Normal distributions. For both the multivariate random variables, the SKEPTIC estimator is defined. Thus, the estimator is a sum of two SKEPTIC estimators and a matrix accounting for skewness. This extension yields a sparser pattern compared to a Gaussian graphical model, resulting in improved interpretability of the resulting graphs. Further, the elliptical skew-(S)KEPTIC estimator shows a slower exponential concentration in $\|\cdot\|_{\max}$ rate than the SKEPTIC estimator. Nevertheless, our findings show a reliable outcome that is useful when dealing with data that exhibit high skewness when it cannot be neglected. Our approach’s effectiveness is validated through optimal graph recoveries in numerical simulations.
Appendix A Proofs of elliptical skew-(S)KEPTIC estimators
A.1 Proof of Theorem 4.1
Proof.
Let $\hat{\rho}_{ij}$ denote the Spearman’s rho estimator of $\Sigma^{0}_{ij}=Corr(X_{i},X_{j})$, the estimator is biased (Zimmerman et al., , 2003).
$\mathbb{E}(\widehat{\rho}_{ij})=\frac{6}{\pi(n+1)}\left[\arcsin\left(\Sigma_{ij}^{0}\right)+(n-2)\arcsin\left(\frac{\Sigma_{ij}^{0}}{2}\right)\right],$ (24)
we rewrite (24) as:
$\Sigma^{0}_{ij}=2\sin\left(\frac{\pi}{6}\mathbb{E}(\hat{\rho}_{ij})+a_{ij}\right)\text{where: }\quad a_{ij}=\frac{\pi\mathbb{E}(\hat{\rho}_{ij})-2\arcsin(\Sigma^{0}_{ij})}{2(n-2)}.$ (25)
The term $|a_{ij}|\leq\frac{\pi}{n-2}$, therefore for all $n>\frac{6\pi}{t}+2$, we have $|a_{ij}|<\frac{t}{6}$ and thus:
$\mathbb{P}\left(\left|\hat{S}^{\rho,\alpha}_{ij}-\Sigma^{0}_{ij}\right|>t\right)=$ (26)
$=p^{2}\mathbb{P}\left(\left|4\sin\left(\frac{\pi}{6}A_{\hat{\rho}_{ij}}\right)-2\sin\left(\frac{\pi}{6}\mathbb{E}(\hat{\rho}_{ij})+a_{ij}\right)\right|>t\right)$
$=p^{2}\mathbb{P}\left(2\left|2\sin\left(\frac{\pi}{6}A_{\hat{\rho}_{ij}}\right)-\sin\left(\frac{\pi}{6}\mathbb{E}(\hat{\rho}_{ij})+a_{ij}\right)\right|>t\right)$
$=p^{2}\mathbb{P}\left(\left|2\sin\left(\frac{\pi}{6}A_{\hat{\rho}_{ij}}\right)-\sin\left(\frac{\pi}{6}\mathbb{E}(\hat{\rho}_{ij})+a_{ij}\right)\right|>\frac{t}{2}\right)$
$\leq p^{2}\mathbb{P}\left(\left|2\left(A_{\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})+\frac{6}{\pi}a_{ij}\right|>\frac{3}{\pi}t\right)$
$\leq p^{2}\mathbb{P}\left(\left|2\left(A_{\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{3}{\pi}t-\left|\frac{6}{\pi}a_{ij}\right|\right)$
$\leq p^{2}\mathbb{P}\left(\left|2\left(A_{\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{3}{\pi}t-\frac{1}{\pi}t\right)$
$=p^{2}\mathbb{P}\left(\left|2\left(A_{\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{2}{\pi}t\right)$
The correlation matrix using Spearman rho, $\hat{\Sigma}_{\rho}$, can be decomposed, using modified Cholesky decomposition as $\hat{\Sigma}_{\rho}=\hat{L}\hat{D}_{\kappa}\hat{L}^{\top}$, where $\hat{D}_{\kappa}=\operatorname{I}_{p}$. In our case, we have: $A_{\hat{\rho}_{ij}}=\hat{L}(I-\hat{D}^{2})\hat{L}^{\top}$. By definition the $k$-th element of $(1-\hat{d}_{kk}^{2})<1$. Thus:
$\mathbb{P}\left(\left|2\left(A_{\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{2}{\pi}t\right)\leq\mathbb{P}\left(\left|2\left({\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{2}{\pi}t\right)$ (27)
We prove the bound of the estimator by rewriting Spearman’s rho in terms of the U-statistic and using the Hoeffding inequality. This implies:
$\mathbb{P}\left(\sup_{ij}\left|\hat{S}^{\rho,\alpha}-\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{2}{\pi}t\right)\leq\exp\left(-\frac{nt^{2}}{16\pi^{2}}\right)$ (28)
A.2 Proof of Theorem 4.2
Proof.
Kendall’s tau is an unbiased estimator: $\mathbb{E}(\hat{\tau}_{ij})=\tau_{ij}$, where $\tau_{ij}:=\operatorname{Corr}(\operatorname{sign}(X_{i}-\widetilde{X}_{i}),\operatorname{sign}(X_{j}-\widetilde{X}_{j}))$ is the population version.
$\mathbb{P}\left(\left|\hat{\tau}_{ij}-\tau_{ij}\right|>t\right)=$
$=\mathbb{P}\left(\left|2\sin\left(\frac{\pi}{2}A_{\hat{\tau}_{ij}}\right)-\sin\left(\frac{\pi}{2}\tau_{ij}\right)\right|>t\right)$
$\leq\mathbb{P}\left(\left|2A_{\hat{\tau}_{ij}}-\tau_{ij}\right|>\frac{2}{\pi}t\right)$
As in the Proof in Theorem 4.1 we have: $(I-\hat{D}^{2})<I$. This implies:
$\mathbb{P}\left(\sup_{ij}|\hat{S}^{\tau,\alpha}-\Sigma^{0}|>t\right)\leq\exp\left(-\frac{nt^{2}}{4\pi^{2}}\right)$ (29)
This result is obtained by exploiting the fact that we can write Kendall’s tau in terms of the U-statistic and using the Hoeffding inequality. ∎

### v2

@@
-Semiparametric Skew-Elliptical Distributions For High-Dimensional Graphical Models
+Robust Semiparametric Graphical Models with Skew-Elliptical Distributions
 Abstract
-We propose a semiparametric approach called elliptical skew-(S)KEPTIC for efficiently and robustly estimating non-Gaussian graphical models.
-Relaxing the assumption of semiparametric elliptical distributions to the family of meta skew-elliptical that accommodates a skewness component, we derive a new estimator which is an extension of the SKEPTIC estimator in Liu et al., 2012a , based on semiparametric Gaussian copula graphical models, to the case of skew-elliptical copula graphical models.
-Theoretically, we demonstrate that the elliptical skew-(S)KEPTIC estimator achieves robust parametric convergence rates in both graph recovery and parameters estimation.
-We conduct numerical simulations to prove the reliable graph recovery performance of the elliptical skew-(S)KEPTIC estimator.
-Finally, the new method is applied to the daily log-returns of the stocks of the S&P500 index and shows better interpretability compared to the Gaussian copula graphical models.
-Keywords—High-dimensional statistics, undirected graphical models, elliptical distributions, skewness, robust statistics.
+We propose semiparametric estimators called elliptical skew-(S)KEPTIC for efficiently and robustly estimating non-Gaussian graphical models, relaxing the assumption of semiparametric elliptical distributions to the family of meta skew-elliptical, which accommodates a skewness component.
+Theoretically, we demonstrate that the elliptical skew-(S)KEPTIC estimators achieve robust convergence rates in both graph recovery and parameters estimation.
+We conduct numerical simulations that prove the reliable graph recovery performance of the elliptical skew-(S)KEPTIC estimators.
+Finally, the new method is applied to the daily log-returns of the stocks of the S&P500 index and shows higher sparsity compared to the Gaussian copula graphical models.
+Keywords: Elliptical distributions, robust statistics, skewness, and undirected graphical models.
 1 Introduction
-We consider the problem of estimating high-dimensional undirected graphical models.
+We consider the problem of estimating undirected graphical models.
 Given a $p$-dimensional random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$, we aim to estimate the undirected graph $G:=(V,E)$, where the vertices set $V:=\{1,\ldots,p\}$ contains the number of nodes corresponding to the $p$ variables, and the edge set $E$ indicates the conditional independence relationships between $X_{1},\ldots,X_{p}$.
@@
 If $X_{i}$ and $X_{j}$ are conditionally independent given $X_{\setminus{\{i,j}\}}$, then there exists no edge between the corresponding nodes $i$ and $j$ in the graph G, denoted as $(i,j)\notin E$.
-One reliable approach to estimate the undirected graphical models is the semiparametric nonparanormal (NPN) method proposed in Liu et al., (2009), which relaxes the assumption of Gaussian data, considering a set of functions $\{f_{j}\}_{j=1}^{p}$ such that: $f(\boldsymbol{X})=(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim N(0,\Sigma^{0})$, where $\Omega=(\Sigma^{0})^{-1}$.
+One reliable approach to estimate the undirected graphical models is the semiparametric nonparanormal (NPN) method proposed in Liu et al., (2009), which relaxes the assumption of Gaussian data, considering a set of functions $\{f_{j}\}_{j=1}^{p}$ such that: $f(\boldsymbol{X})=(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\operatorname{N}(0,\Sigma^{0})$, where $\Omega=(\Sigma^{0})^{-1}$.
 When $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$, no edge connects $X_{i}$ and $X_{j}$ if $\Omega_{ij}=0$.
@@
 Liu et al., 2012a propose the SKEPTIC (Spearman/Kendall estimates preempt transformations to infer correlation) estimator for the nonparanormal distribution, a regularized rank-based correlation coefficient estimator based on Kendall’s tau and Spearman’s rho statistics.
-The estimated correlation matrix is then puggled into shrinkage procedures that reveal the sparsity of the graph.
-The nonparanormal distribution exploits a Gaussian-based kernel and belongs to the broad class of meta-elliptical distributions described in Han and Liu, (2014), as well as Liu et al., 2012b , where a set of functions $\{f_{j}\}_{j=1}^{p}$ such that $f(\boldsymbol{X})\sim\mathcal{E}C(0,\Sigma^{0},\xi)$ is considered.
-In particular, Liu et al., 2012b show that the graph $G$ is encoded in the latent generalized partial correlation matrix $\Xi:=-\left[\text{diag}(\Omega)\right]^{-1/2}\Omega\left[\text{diag}(\Omega)\right]^{-1/2}$, thus no edge connects $X_{i}$ and $X_{j}$ if $\Xi_{ij}=0$, which implies $\Omega_{ij}=0$.
-The main drawback of the meta-elliptical class of distributions is that, even though it embeds skewed distributions, the skewness term does not contribute to reveal the conditional independence structure.
-For this reason, we introduce the meta skew-elliptical family of distributions defined as $\boldsymbol{X}\sim MS\mathcal{E}(0,\Sigma^{0},\Lambda;f)$ considering a set of functions $f$ such that $f(\boldsymbol{X})\sim SU\mathcal{E}(0,\Sigma^{0},\Lambda,h^{(p+q)},\tau,\Gamma)$ described in Arellano-Valle and Genton, 2010b .
-In this paper, we propose a rank-based correlation matrix estimator, which we refer to as elliptical skew-(S)KEPTIC estimator, extending the SKEPTIC estimator to the $MS\mathcal{E}$ family of distributions, adding respect to nonparanormal models the possibility to rely on graphical models that take into account other distributional features such as directional skewness and heavy tails.
-The main idea is to consider the $\operatorname{SU\mathcal{E}}$ as a combination of the closed skew-Normal distribution introduced in Domìnguez-Molina et al., (2003), which belongs to the class of $\operatorname{SU\mathcal{E}}$ family of distributions, and possesses the stochastic representation defined through Normal distribution and a half-Normal distribution, for which the SKEPTIC estimator of the correlation matrix is defined.
-The estimated correlation matrix is plugged into existing shrinkage procedures.
+The estimated correlation matrix is then plugged into shrinkage procedures that reveal the sparsity of the graph.
+The nonparanormal distribution exploits a Gaussian-based kernel and belongs to the broad class of meta-elliptical distributions described in Han and Liu, (2014) and in Liu et al., 2012b , where a set of functions $\{f_{j}\}_{j=1}^{p}$ such that $f(\boldsymbol{X})\sim\mathcal{E}C(0,\Sigma^{0},\xi)$ is considered.
+In particular, Liu et al., 2012b show that the graph $G$ is encoded in the latent partial correlation matrix $\Xi:=-\left[\text{diag}(\Omega)\right]^{-1/2}\Omega\left[\text{diag}(\Omega)\right]^{-1/2}$, thus no edge connects $X_{i}$ and $X_{j}$ if $\Xi_{ij}=0$, which implies $\Omega_{ij}=0$.
+However, as shown by Rossell and Zwiernik, (2021), the dependence structure of the elliptical distribution differs from that of the Gaussian case.
+In addition, a major pitfall of the meta-elliptical class of distributions is that, even though it embeds skewed distributions, the skewness term does not contribute to reveal the conditional independence structure.
+We propose rank-based correlation matrix estimators, which we refer to as elliptical skew-(S)KEPTIC estimators.
+These estimators refer to the second moment of the meta skew-elliptical family of distributions, that we introduce in this paper considering a set of functions $\{f_{j}\}_{j=1}^{p}$ such that $f(\boldsymbol{X})\sim\operatorname{SU}\mathcal{E}(0,\Sigma^{0},\Lambda,h^{(p+q)},\tau,\Gamma)$, where the unified skew-elliptical ($\operatorname{SU}\mathcal{E}$) distribution is described in Arellano-Valle and Genton, 2010b .
+Extending the SKEPTIC estimator to the meta skew-elliptical family of distributions, our methodology allows for a semiparametric estimation of graphical models that take into account distributional features such as directional skewness and heavy tails.
+The main idea is to consider the unified skew-elliptical distribution as a combination of the closed skew-normal distribution introduced in Domìnguez-Molina et al., (2003), which belongs to the class of $\operatorname{SU\mathcal{E}}$ distributions.
+The class of unified skew-elliptical distributions, which generalizes the closed skew-normal, enjoys key closure properties under marginalization, conditioning, and linear transformations.
+These properties enable e.g., to define conjugate priors in probit, tobit, and multinomial probit regression models based on $\operatorname{SU\mathcal{E}}$ distributions, as both the prior and posterior remain within the same family (Durante,, 2019, Karling et al.,, 2024).
+The unified skew-elliptical distribution can be written in terms of the closed-skew normal, that possesses a stochastic representation defined through a normal distribution and a half-normal distribution, for which the SKEPTIC estimator of the correlation matrix is defined.
+The estimated correlation matrix is then plugged into existing shrinkage procedures.
 For instance, Meinshausen and Bühlmann, (2006) study a neighborhood selection approach with the lasso operator, Banerjee et al., (2008) and Friedman et al., (2008) propose an approach based on penalized likelihood, and Yuan, (2010) and Cai et al., (2011) introduce, respectively, graphical Dantzig selector and CLIME which are two methods based on linear programming that provide an improvement in terms of theoretical properties respect to the other shrinkage procedures.
-We generalize the results in Liu et al., 2012a extending the SKEPTIC estimator to the elliptical skew-(S)KEPTIC estimator by introducing a directional skewness term, encoded in the matrix $\Lambda$, which assigns greater importance to the asymmetry present in data distributions.
-The proposed estimator is a version of the SKEPTIC estimator conforming to the ${MS}\mathcal{E}$ distribution and embeds two new subclasses of estimators based on Kendall’s tau and Sparman’s rho statistics.
-While the nonparanormal distribution is a flexible approach to model non-Normal data, the meta skew-elliptical distribution allows us to quantify the asymmetry (and the kurtosis) presented in the data.
-When the $\operatorname{SU\mathcal{E}}$ distribution reduces to the $\operatorname{CSN}$ distribution, we consider both the statistics for estimating the unknown correlation matrix and we define the elliptical skew-SKEPTIC estimator.
-Otherwise, we only consider Kendall’s tau statistic, and we define elliptical skew-KEPTIC estimator since Spearman’s rho statistic is not invariant for the class of elliptical distributions (Hult and Lindskog, , 2002).
-The proposed framework extends to a greater variety of semiparametric graphical models, introducing the possibility of modeling the combination of directional skewness and heavy tails data and adding flexibility to the nonparanormal-type of data transformations.
-We investigate the theoretical properties of the elliptical skew-(S)KEPTIC estimator, and find bounds equal to twice the bounds of the SKEPTIC estimator.
-The increase in model flexibility justifies the slight reduction in the robustness of the proposed estimators.
-We provide numerical studies to support our theory.
-Through a backward approach, we define the true sparsity pattern given by $\Omega$ and provide the graph recovery using the elliptical skew-(S)KEPTIC estimator which is compared to the SKEPTIC estimator.
-The elliptical skew-(S)KEPTIC estimator proves an optimal graph recovery, giving statistical efficiency to the proposed method.
+The proposed estimators generalize the SKEPTIC estimator conforming to the meta skew-elliptical distribution and embed two new subclasses of estimators based on Kendall’s tau and Spearman’s rho statistics.
+While the nonparanormal distribution is a flexible approach to model non-Gaussian data, the approach based on the meta skew-elliptical distribution allows us to quantify the asymmetry and the kurtosis presented in the data.
+The elliptical skew-(S)KEPTIC estimators possess skewness terms, encoded in the matrix of skewness parameters $\Lambda$, which assigns greater importance to the asymmetry present in data distributions.
+When the unified skew-elliptical distribution reduces to the closed skew-normal distribution, we consider both the statistics for estimating the unknown correlation matrix and we define the elliptical skew-SKEPTIC estimator.
+Otherwise, we only consider Kendall’s tau statistic and we define the elliptical skew-KEPTIC estimator, since Spearman’s rho statistic is not invariant for the class of elliptical distributions (Hult and Lindskog,, 2002).
+The proposed framework extends to graphical models that incorporate both skewness and heavy tails, adding flexibility to nonparanormal-type data transformations.
+We prove that the elliptical skew-(S)KEPTIC estimators have deviation from the true correlation matrix $\Sigma^{0}$ that decreases as the sample size grows, showing a fast exponential rate of convergence in $\|\cdot\|_{\max}$ norm.
+The numerical analysis supports the theory.
+Through a backward approach, we define the true sparsity pattern given by $\Omega$ and show the graph recovery using the elliptical skew-(S)KEPTIC estimators, which are compared to the SKEPTIC estimator.
+The elliptical skew-(S)KEPTIC estimators prove an optimal graph recovery, showing statistical efficiency of the proposed method.
 Prior studies by Zareifard et al., (2016), Nghiem et al., (2022), and Sheng et al., (2023) have explored graphical models while incorporating skewness into the sparsity pattern estimation.
-Zareifard et al., (2016) exploit a multivariate $\operatorname{CSN}$ distribution to define the skew Gaussian graphical models which are estimated through a Bayesian approach.
+Zareifard et al., (2016) exploit a multivariate closed skew-normal distribution to define the skew Gaussian graphical models, which are estimated through a Bayesian approach.
 Nghiem et al., (2022) propose a novel nodewise regression approach to estimate the graph on data generated from a generalized multivariate skew-normal distribution.
 Instead, Sheng et al., (2023) include the shape parameter in the determination of the precision matrix using the skew-normal distribution of Azzalini, (1985).
-They build an algorithm that penalizes the likelihood of this distribution in order to estimate $\Omega$ and thus the graph.
-However, these studies do not emphasize the direct estimation of the correlation matrix while accounting for skewness.
-The paper is organized as follows: in Section 2 we recall the background, in Section 3 we define the ${MS}\mathcal{E}$ distribution, which we use to derive the elliptical skew-(S)KEPTIC estimator and in Section 4 we study the theoretical properties of the proposed estimator.
-In Section 5, we conduct experiments comparing elliptical skew-(S)KEPTIC estimator with the SKEPTIC estimator through numerical comparisons on synthetic data and with an empirical analysis on the S&P 500 stocks data.
-Section 6 concludes.
+They develop an algorithm that penalizes the likelihood of this distribution to estimate $\Omega$ and, consequently, the graph.
+However, these studies do not focus on the direct estimation of the correlation matrix while incorporating skewness.
+The paper is organized as follows.
+In Section 2 we describe the background, including a review of the SKEPTIC estimator and the semiparametric elliptical distribution.
+In Section 3 we introduce the meta skew-elliptical distribution, which forms the basis to derive the elliptical skew-(S)KEPTIC estimators, whereas in Section 4 we provide some shrinkage procedures to estimate the precision matrix $\Omega$.
+Theoretical properties of the proposed estimator are discussed in Section 5.
+Section 6 presents simulation analysis comparing the elliptical skew-(S)KEPTIC estimators with the SKEPTIC estimator on synthetic data, and reports an empirical application to S&P 500 stock returns.
+Section 7 concludes the paper.
 2 Background
-In this section, we briefly recall the literature preceding our article: the SKEPTIC estimator of Liu et al., 2012a , the semiparametric ellitpical distributions in Liu et al., 2012b and Han and Liu, (2014), in this work, we rely of the meta-elliptical distribution..
-The section concludes with the introduction of $\operatorname{SU\mathcal{E}}$ family of distributions in Arellano-Valle and Genton, 2010b , and its link with the undirected graphical models.
 2.1 Notation
@@
 For $1\leq q<\infty$, we define $\|\boldsymbol{a}\|_{q}=\left(\sum_{i=1}^{p}|\boldsymbol{a}|^{q}\right)^{1/q}$, when $q=\infty$: $\|\boldsymbol{a}\|_{\infty}=\max_{1\leq i\leq p}|\boldsymbol{a}_{i}|$.
-Regarding the $A$ matrix, when $q=1$: $\|A\|_{1}=\max_{1\leq j\leq q}\sum_{i=1}^{p}|A_{ij}|$, and when $q=\infty$, $|A|_{\infty}=\max_{1\leq i\leq q}\sum_{j=1}^{p}|A_{ij}|$.
+Regarding the $A$ matrix, when $q=1$: $\|A\|_{1}=\max_{1\leq j\leq q}\sum_{i=1}^{p}|A_{ij}|$, and when $q=\infty$, $\|A\|_{\infty}=\max_{1\leq i\leq q}\sum_{j=1}^{p}|A_{ij}|$.
 We denote $\|A\|_{\max}=\max_{i,j}|A_{ij}|$.
-A vector without its $j$-th element is defined: $\boldsymbol{a}_{\backslash j}=(\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{j-1},\boldsymbol{a}_{j+1},\ldots,\boldsymbol{a}_{p})^{\top}\in\mathbb{R}^{p-1}$, and we denote $A_{\backslash i,\backslash j}\in\mathbb{R}^{p-1\times p-1}$ the matrix without the $i$-$th$ row and the $j$-th column.
+A vector without its $j$-th element is defined: $\boldsymbol{a}_{\backslash j}=(\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{j-1},\boldsymbol{a}_{j+1},\ldots,\boldsymbol{a}_{p})^{\top}\in\mathbb{R}^{p-1}$, and we denote $A_{\backslash i,\backslash j}\in\mathbb{R}^{p-1\times p-1}$ the matrix without the $i$-thth row and the $j$-th column.
 We define: $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ respectively the minimum and the maximum eigenvalue of $A$.
@@
 The focal point is to find a function, $f(\boldsymbol{X})$, that transforms our data into normally distributed data.
-Suppose we have $n$ observations for the $j$-th variable, $x_{1j},...,x_{nj}$, we compute the rank of each observation, and indicate the rank with $r_{1j},...,r_{nj}$.
-Let $\Phi(\cdot)$ denote the Gaussian cumulative distribution function, the $f(\boldsymbol{X})$ that transforms the data is:
-$\hat{f}_{j}(x_{ij})=\Phi^{-1}(\hat{r}_{ij}),\quad\text{where:}\quad\hat{r}_{ij}=\frac{r_{ij}}{n+1}.$ (1)
-We consider the following statistics:
-[…]
-[…]
+Given $n$ observations of the $j$-th variable, $x_{1j},\dots,x_{nj}$, we define the empirical cumulative distribution function (ECDF) $\hat{F}_{j}(w)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(x_{ij}\leq w),$ where $\mathbb{I}(\cdot)$ is the indicator function.
+The transformation function $\hat{f}_{j}(x_{ij})$ is then defined as $\hat{f}_{j}(x_{ij})=\Phi^{-1}(\hat{F}_{j}(x_{ij}))$, where $\Phi^{-1}(\cdot)$ denotes the quantile normal distribution function.
+In practice, the ECDF at each observation $x_{ij}$ can be approximated using its normalize rank: $\hat{F}_{j}(x_{ij})\approx\hat{r}_{ij}=\frac{r_{ij}}{n+1}$, where $r_{ij}$ is the rank of $x_{ij}$ among $x_{1j},\dots,x_{nj}$.
+Thus, the transformation results in
+$\hat{f}_{j}(x_{ij})=\Phi^{-1}\left(\hat{r}_{ij}\right).$ (1)
+This approach ensures that the transformed data approximate a standard normal distribution while preserving the rank structure of the original data. We consider the following statistics:
+$\text{(Spearman's rho)}\ \hat{\rho}_{ij}=\frac{\sum_{k=1}^{n}(r^{k}_{j}-\bar{r}_{j})(r^{k}_{i}-\bar{r}_{i})}{\sqrt{\sum_{k=1}^{n}(r^{k}_{j}-\bar{r}_{j})^{2}\sum_{k=1}^{n}(r^{k}_{i}-\bar{r}_{i})^{2}}},$ (2)
+$\text{(Kendall's tau)}\ \hat{\tau}_{ij}=\frac{2}{n(n-1)}\sum_{1\leq k<k^{\prime}\leq n}\text{sign}((x^{k}_{j}-x^{k^{\prime}}_{j})(x^{k}_{i}-x^{k^{\prime}}_{i})).$ (3)
 Both $\hat{\rho}_{ij}$ and $\hat{\tau}_{ij}$ are non-parametric correlations between the empirical realizations of random variables $X_{j}$ and $X_{i}$.
-Let $\widetilde{X}_{i}$ and $\widetilde{X}_{j}$ be independent copies of the random variables $X_{i}$ and $X_{j}$, respectively.
-The cumulative distribution functions (CDFs) of $X_{i}$ and $X_{j}$ are denoted by $F_{i}$ and $F_{j}$.
-The population versions of Spearman’s rho and Kendall’s tau are defined as $\rho_{ij}:=\operatorname{Corr}(F_{i}(X_{i}),F_{j}(X_{i}))$ and $\tau_{ij}:=\operatorname{Corr}(\operatorname{sign}(X_{i}-\widetilde{X}_{i}),\operatorname{sign}(X_{j}-\widetilde{X}_{j}))$, respectively.
+Let $\widetilde{X}_{i}$ and $\widetilde{X}_{j}$ be independent copies of the random variables $X_{i}$ and $X_{j}$, $F_{i}$ and $F_{j}$ denote the cumulative distribution functions (CDFs) of $X_{i}$ and $X_{j}$, the population versions of Spearman’s rho and Kendall’s tau are defined as: $\rho_{ij}:=\operatorname{Corr}(F_{i}(X_{i}),F_{j}(X_{i}))$ and $\tau_{ij}:=\operatorname{Corr}(\operatorname{sign}(X_{i}-\widetilde{X}_{i}),\operatorname{sign}(X_{j}-\widetilde{X}_{j}))$, respectively.
 For nonparanormal distributions, there is a connection between Spearman’s rho and Kendall’s tau to the underlying Pearson’s correlation coefficient $\Sigma^{0}_{ij}$.
@@
 (Kruskal, (1958), Liu et al., 2012a ): Assuming $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$, we have $\Sigma_{ij}^{0}=2\operatorname{sin}\left(\frac{\pi}{6}\rho_{ij}\right)=\operatorname{sin}\left(\frac{\pi}{2}\tau_{ij}\right)$.
-From these results, we define two estimators for the unknown correlation matrix $(\Sigma^{0}_{ij})$ that are $\hat{S}^{\rho}=\left[\hat{S}^{\rho}_{ij}\right]$ and $\hat{S}^{\tau}=\left[\hat{S}^{\tau}_{ij}\right]$, where: $\hat{S}^{\rho}_{ij}=2\operatorname{sin}\left(\frac{\pi}{6}\hat{\rho}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ and $\hat{S}^{\tau}_{ij}=\operatorname{sin}\left(\frac{\pi}{2}\hat{\tau}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$,
+From these results two estimators of the unknown correlation matrix $(\Sigma^{0}_{ij})$ are $\hat{S}^{\rho}=\left[\hat{S}^{\rho}_{ij}\right]$ and $\hat{S}^{\tau}=\left[\hat{S}^{\tau}_{ij}\right]$, where: $\hat{S}^{\rho}_{ij}=2\operatorname{sin}\left(\frac{\pi}{6}\hat{\rho}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ and $\hat{S}^{\tau}_{ij}=\operatorname{sin}\left(\frac{\pi}{2}\hat{\tau}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$.
 Spearman’s rho and Kendall’s tau statistics are invariant under strictly increasing transformations of the underlying random variable.
-This concept will be fundamental in the definition of the elliptical skew-(S)KEPTIC estimator.
 2.3 The meta-elliptical distribution
-The meta-elliptical distribution introduced in Han and Liu, (2014) relaxes the assumption of elliptically distributed data, in the same way as the nonparanormal of Liu et al., (2009) do for Gaussian distributions.
+The meta-elliptical distribution introduced in Han and Liu, (2014) relaxes the assumption of elliptically distributed data, in a similar way as the nonparanormal of Liu et al., (2009) does for Gaussian distributions.
 The meta-elliptical family is a strict extension of the nonparanormal distribution.
 Definition 2.2.
-[…]
+(Han and Liu, (2014), Fang, (2018)) A random vector $\boldsymbol{X}$ has an elliptical distribution, $\boldsymbol{X}\sim\mathcal{E}C_{p}(\mu,\Sigma,\xi)$, if $\boldsymbol{X}$ has the following stochastic representation: $\boldsymbol{X}\overset{d}{=}\mu+\xi AU$.
+Where: $\mu\in\mathbb{R}^{p}$, $q:=rank(A)$, AA is a $p\times q$ matrix such that $\Sigma=AA^{\top}$, $\xi\geq 0$ is a random variable independent of $U$, $U\in\mathbb{S}^{q-1}$, which is uniformly distributed of the unit sphere in $\mathbb{R}^{q}$.
 Let $\mathcal{R}_{p}:=\{\Sigma\in\mathbb{R}^{p\times p}:\Sigma^{\top}=\Sigma,\text{diag}\left(\Sigma\right)=1,\Sigma\succeq 0\}$, in analogy to the nonparanormal case, the meta-elliptical distribution is defined as follows.
 Definition 2.3.
-(Han and Liu, (2014)) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has a meta-elliptical distribution, denoted by $\boldsymbol{X}\sim M\mathcal{E}_{p}(\Sigma^{0},\xi;f)$, if there exist a set of univariate monotone functions $f_{1},\ldots,f_{p}$ such that $(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\mathcal{EC}_{p}(0,\Sigma^{0},\xi)$, where $\Sigma^{0}\in\mathcal{R}_{p}$ is defined as the latent generalized correlation matrix.
+(Han and Liu, (2014)) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has a meta-elliptical distribution, denoted by $\boldsymbol{X}\sim M\mathcal{E}_{p}(\Sigma^{0},\xi;f)$, if there exist a set of univariate monotone functions $f_{1},\ldots,f_{p}$ such that $(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\mathcal{EC}_{p}(0,\Sigma^{0},\xi)$, where $\Sigma^{0}\in\mathcal{R}_{p}$ is the latent correlation matrix.
 For the distributions that belong to the meta-elliptical family, Kendall’s tau statistic in Lemma 2.1 can be exploited to estimate the unknown correlation matrix.
-Nevertheless, the meaning of a missing edge is different with respect to the nonparanormal case.
+Nevertheless, the meaning of a missing edge is different for the nonparanormal case.
 Indeed, when $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$ a missing edge implies conditional independence.
-On the other hand, when $\boldsymbol{X}\sim M\mathcal{E}(\Sigma^{0},\xi,f)$ a missing edge means conditional linear independence.
-The meta-elliptical family contains many useful distributions among which asymmetric distributions.
-However, the estimate of the correlation matrix does not directly take into account the skewness parameters.
-We extend the meta-elliptical family giving a direct contribution of the skewness parameters for the conditional independence relations, and thus considering the $\mathcal{E}$ family of distributions.
-2.4 $\mathcal{E}$ distributions
+On the other hand, when $\boldsymbol{X}\sim M\mathcal{E}(\Sigma^{0},\xi,f)$ a missing edge means uncorrelation.
+[…]
+[…]
 We recall the family of $\operatorname{SU\mathcal{E}}$ distributions described in Arellano-Valle and Genton, 2010b .
-If a random vector has a $\operatorname{SU\mathcal{E}}$ distribution, we write $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,q}(\mu,\Sigma,\Lambda,h^{(p+q)},\tau,\Gamma)$, where $h^{(p+q)}$ is the density generator.
-We focus on two specifications of the family which are: the unified skew-$t$ ($\operatorname{SUT}$) distribution (Arellano-Valle and Genton, 2010a ), and the unified skew-Normal ($\operatorname{SUN}$) distribution (Arellano-Valle and Azzalini, (2006)).
-Table 1 shows how the $\operatorname{SU\mathcal{E}}$ distribution reduces into $\operatorname{SUN}$ and $\operatorname{SUT}$ distributions.
+If a random vector $\boldsymbol{X}$ has a $\operatorname{SU\mathcal{E}}$ distribution, we write $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,p}(\mu,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$, where $h^{(2p)}$ is the density generator.
+We focus on two specifications of the family, which are: the closed skew-normal (CSN) distribution of Domìnguez-Molina et al., (2003) [^1] We build on the closed skew-normal (CSN) distribution, which represents a reparametrization of the unified skew-normal (SUN) distribution employed by Arellano-Valle and Genton, 2010b .
+and the unified skew-$t$ (SUT) distribution (Arellano-Valle and Genton, 2010a ).
+Table 1 shows how the $\operatorname{SU\mathcal{E}}$ distribution reduces into CSN and SUT distributions.
 Formally, we define the $\operatorname{SU\mathcal{E}}$ distribtuion.
 Definition 2.4.
-(Arellano-Valle and Genton, 2010b ) A continuous p-dimensional random vector $X$ has a multivariate unified skew-elliptical ($\mathcal{E}$) distribution, denoted by $\boldsymbol{X}\sim SU\mathcal{E}_{p,q}(\mu,\Sigma,\Lambda,h^{(p+q)},\tau,\Gamma)$, if its density function at $x\in\mathbb{R}^{p}$ is
-$f_{X}(x)=\frac{1}{F_{q}\left(\tau;\Gamma+\Lambda\bar{\Sigma}\Lambda^{T},h^{(q)}\right)}f_{p}\left(y;\mu,\Sigma,h^{(p)}\right)F_{q}\left(\Lambda z+\tau;\Gamma,h_{Q(z)}^{(q)}\right),$ (4)
-where $w=\sigma^{-1}(y-\mu),Q(w)=w^{\top}(\Sigma^{0})^{-1}w,\Lambda$ is a $q\times p$ real matrix controlling shape, $\tau\in\mathbb{R}^{q}$ is the extension parameter, $\Gamma$ is a $q\times q$ positive definite correlation matrix, $f_{p}\left(y;\mu,\Sigma,h^{(p)}\right)=$ $|\Sigma|^{-1/2}h^{(p)}(Q(w))$ denotes the density function of an elliptically contoured distribution with location $\mu\in\mathbb{R}^{p}$, positive definite $p\times p$ dispersion matrix $\Sigma$, with $p\times p$ scale and correlation matrices $\sigma=\operatorname{diag}(\Sigma)^{1/2}$ and $\Sigma^{0}=\sigma^{-1}\Sigma\sigma^{-1}$, respectively, and density generator $h^{(p)},F_{r}\left(y;\Sigma,h^{(r)}\right)$ denotes the $r$-dimensional centered elliptical cumulative distribution function with $r\times r$ dispersion matrix $\Sigma$ and density generator $h^{(q)}$, and $h_{Q(w)}^{(q)}(u)=h^{(p+q)}(u+Q(w))/h^{(p)}(Q(w))$.
+(Arellano-Valle and Genton, 2010b, ) A continuous p-dimensional random vector $\boldsymbol{X}$ has a multivariate unified skew-elliptical $\operatorname{(SU\mathcal{E})}$ distribution, denoted by $\boldsymbol{X}\sim SU\mathcal{E}_{p,p}(\mu,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$, if its density function at $x\in\mathbb{R}^{p}$ is
+$f_{X}(x)=2^{p}f_{p}\left(x;\mu,\Sigma,h^{(p)}\right)F_{p}\left(\Lambda z;\Gamma,h_{Q(z)}^{(p)}\right),$ (4)
+where $\mu$ is the location parameter, $\Sigma$ is a $p\times p$ dispersion matrix, $h^{(p)}$ is the density generator, $F_{r}\left(x;\Sigma,h^{(r)}\right)$ the $r$-dimensional centered elliptical cumulative distribution function with $r\times r$ dispersion matrix $\Sigma$ and density generator $h^{(r)}$.
+Additionally, $z=\sigma^{-1}(x-\mu)$, $\Lambda$ is a $p\times p$ real matrix controlling shape, $Q(z)=z^{\top}(\Sigma^{0})^{-1}z$, $\Gamma$ is a $p\times p$ positive definite symmetric matrix, $f_{p}\left(x;\mu,\Sigma,h^{(p)}\right)=$ $|\Sigma|^{-1/2}h^{(p)}(Q(z))$ denotes the density function of an elliptically contoured distribution with location $\mu\in\mathbb{R}^{p}$.
+In conclusion, $h_{Q(z)}^{(p)}(u)=h^{(2p)}(u+Q(z))/h^{(p)}(Q(z))$.
 $\operatorname{SU\mathcal{E}}$ $h^{(p)}$
-$\operatorname{SUN}$ $h^{(p)}=\phi^{(p)}$$(u)$where: $\phi^{(p)}(u)=\frac{1}{(2\pi)^{p/2}}\exp\left(-\frac{u^{2}}{2}\right)$
-$\operatorname{SUT}$ $h^{(p)}=t^{(p)}(u,\nu)$ where: $t^{(p)}(u,\nu)=\frac{\Gamma\left(\frac{\nu+p}{2}\right)}{\Gamma\left(\frac{\nu}{2}\right)\left(\pi a\right)^{p/2}}\left(1+\frac{u}{\nu}\right)^{-\frac{\nu+p}{2}}$
+CSN $h^{(p)}=\phi^{(p)}$$(u)$  where: $\phi^{(p)}(u)=\frac{1}{(2\pi)^{p/2}}\exp\left(-\frac{u^{2}}{2}\right)$
+SUT $h^{(p)}=t^{(p)}(u,\nu)$ where: $t^{(p)}(u,\nu)=\frac{\Gamma\left(\frac{\nu+p}{2}\right)}{\Gamma\left(\frac{\nu}{2}\right)\left(\pi a\right)^{p/2}}\left(1+\frac{u}{\nu}\right)^{-\frac{\nu+p}{2}}$
 Table 1: Subclasses of skew-elliptical distribution based on the density generator $h^{(p)}$.
-The relation between the $\operatorname{SUN}$ and the $\operatorname{SU\mathcal{E}}$ is defined through the following stochastic representation.
+The relation between the CSN and the $\operatorname{SU\mathcal{E}}$ is defined through the following stochastic representation.
 Proposition 2.5.
-(Arellano-Valle and Genton, 2010b ) Given $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,q}(\mu,\Sigma,\Lambda,h^{(p+q)},\tau,\Gamma)$, when the extension parameter $\tau=0$, we can write $\boldsymbol{X}$ as a mixture of a $\operatorname{SUN}$ distribution: $\boldsymbol{X}=\mu+\sigma V^{-1/2}\boldsymbol{Z}_{0}$, where $V_{0}$ is the cumulative distribution function of a uni-dimensional random variable, called mixture variable, such that: $V_{0}(0)=0$, and $V_{0}$ is independent of $\boldsymbol{Z}_{0}\sim SUN_{p,q}(0,\Sigma^{0},\Lambda,0,\Gamma)$.
-The $\operatorname{SUN}$ distribution results in a different parameterization of the closed skew-Normal (CSN) distribution in Domìnguez-Molina et al., (2003).
-The CSN distribution corresponds to the case in which the cumulative distribution functions in Definition 2.4 are the Gaussian cumulative distribution functions.
-Therefore, the $\operatorname{SU\mathcal{E}}$ distribution can be expressed in terms of the $\operatorname{CSN}$ distribution.
-This is useful for deriving the moments and for sampling random numbers from the $\operatorname{SU\mathcal{E}}$ distribution.
-The $\operatorname{CSN}$ distribution has the following stochastic representation as a combination of a Normal distribution and a half-Normal distribution.
+(Arellano-Valle and Genton, 2010b ) Given $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,p}(\mu,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$, we can write $\boldsymbol{X}$ as a mixture of a CSN distribution: $\boldsymbol{X}=\mu+\sigma V_{0}^{-1/2}\boldsymbol{Z}_{0}$, where $V_{0}$ is the cumulative distribution function of a uni-dimensional random variable, called mixture variable, such that: $V_{0}(0)=0$, and $V_{0}$ is independent of $\boldsymbol{Z}_{0}\sim\operatorname{CSN}_{p,p}(0,\Sigma^{0},\Lambda,0,\Gamma)$.
+Lemma 2.6.
+(Arellano-Valle and Genton, 2010b, ) If  $\text{ $V_{0}$}\sim Gamma(\nu/2,\nu/2)$, then
+$\boldsymbol{X}\sim\operatorname{SU}\mathcal{E}_{p,p}(\mu,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$ reduces to $\operatorname{SUT}$ distribution: $\boldsymbol{X}\sim\operatorname{SUT}_{p,p}(\mu,\Sigma,\Lambda,\nu,0,\Gamma)$.
+We construct an undirected graph on the closed skew-normal distribution, following Zareifard et al., (2016).
 Definition 2.7.
-(Domìnguez-Molina et al., (2003)) Let $\boldsymbol{Y}\sim\operatorname{CSN}_{p,q}(\mu,\Sigma,\Gamma,0,\Delta)$, the random vector $\boldsymbol{Y}$ admits the following stochastic representation:
-[…]
-where: $\boldsymbol{U}\sim\operatorname{N}_{p}(0,I_{p})$ and $\boldsymbol{V}\sim\operatorname{HN}_{q}(0,\Delta+\Gamma\Sigma\Gamma)$, the Normal and the half-Normal distribution, respectively.
-The stochastic representation in (5) allows us to estimate an undirected decomposable graph.
+Consider an undirected decomposable graph $G=(V,E)$, a random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ is called a skew decomposable graphical model concerning the graph G with mean $\mu$, the precision matrix $\Omega$ and the skewness parameter $\alpha\in\mathbb{R}^{p}$ if its density is $\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ or equivalently
+$f_{\boldsymbol{X}}(x)=2^{p}\phi_{p}\left(x-\mu;0,\Omega^{-1}\right)\Phi_{p}\left(D_{\alpha}L(x-\mu);0,D^{-1}_{\kappa}\right),$ (5)
+where $L$ and $D_{\kappa}$ are the matrix resulting from the modified Cholesky of $\Omega=LD_{\kappa}L^{\top}$, and zero in $L$ matrix determines the conditional independence relations between the $X_{1},\ldots,X_{p}$ variables, in other words: $L_{ij}=0\Leftrightarrow\Omega_{ij}=0$.
+Further, when the matrix $D_{\alpha}=0$, the random vector $\boldsymbol{X}\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ reduces to $\boldsymbol{X}\sim\operatorname{N}_{p}(0,\Omega^{-1})$.
+The graphical model for the unified skew-elliptical distribution can be constructed by utilizing the sparsity structure of the closed skew-normal distribution.
+We rely on a semiparametric approach that considers the class of unified skew-elliptical distributions to extend the SKEPTIC estimator of Liu et al., 2012a proposing a robust estimator of graphical models that accounts for skewness and heavy tails.
+3 The elliptical skew-(S)KEPTIC estimators
 Definition 3.1.
-(Zareifard et al., (2016)) Consider an undirected decomposable graph $G=(V,E)$, a random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ is called a skew decomposable graphical model concerning the graph G with mean $\mu$, the precision matrix $\Omega$ and the skewness parameter $\alpha\in\mathbb{R}^{p}$ if its density is $\operatorname{CSN}_{p,p}(\mu,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ or equivalently
-$f_{\boldsymbol{X}}(x)=2^{p}\phi_{p}\left(x;\mu,\Omega^{-1}\right)\Phi_{p}\left(D_{\alpha}L(x-\mu);0,D^{-1}_{\kappa}\right),$ (6)
-where $L$ and $D_{\kappa}$ are the matrix resulting from the modified Cholesky of $\Omega=LD_{\kappa}L^{\top}$, and zero in $L$ matrix determines the conditional independence relations between the $X_{1},\ldots,X_{p}$ variables, further, when the matrix $D_{\alpha}=0$, the random vector $\boldsymbol{X}\sim\operatorname{CSN}_{p,p}(\mu,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ reduces to $\boldsymbol{X}\sim\operatorname{N}_{p}(\mu,\Omega^{-1})$.
-We introduce an estimator for skew-Gaussian graphical models defined in Zareifard et al., (2016) that extends the SKEPTIC including a skewness component to accommodate non-symmetric data.
-3 The elliptical skew-(S)KEPTIC estimator
-Definition 3.1.
-(Meta skew-elliptical distribution) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has meta skew-elliptical distribution, denoted by $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},0,\Lambda;f)$, if there exists a set of monotone univariate functions $f_{1},\ldots,f_{p}$ and a nonnegative random variable $\xi$, such that
-$(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\operatorname{SU}\mathcal{E}_{p,p}(0,\Sigma^{0},\Lambda,0,\Gamma),$ (7)
+(Meta skew-elliptical distribution) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has meta skew-elliptical distribution, denoted by $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},\Lambda,h^{(2p)},0,\Gamma;f)$, if there exists a set of monotone univariate functions $f_{1},\ldots,f_{p}$ such that
+$(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Sigma^{0},\Lambda,h^{(2p)},0,\Gamma),$ (6)
 where $\Lambda$ is a matrix containing the skewness terms.
@@
 When $\Lambda=0$ and $h^{(p)}=\phi^{(p)}(u)$, then the meta skew-elliptical distribution reduces to the nonparanornmal distribution.
-We exploit the stochastic representation of $\operatorname{SU\mathcal{E}}$ distribution through a CSN distribution (Proposition 2.5) to define the skew-elliptical graphical model.
-That is, we propose a nonparametric rank-based estimator, which we call elliptical skew-(S)KEPTIC estimator, based on the stochastic representation (5), according to the parametrization of $\operatorname{CSN}$ distribution in (6):
-$\boldsymbol{Z}\stackrel{{\scriptstyle p}}{{=}}\boldsymbol{\mu}+L^{-1}D_{\kappa}^{-\frac{1}{2}}D_{\alpha}\left(I+D_{\alpha}^{2}\right)^{-\frac{1}{2}}\boldsymbol{U}+L^{-1}D_{\kappa}^{-\frac{1}{2}}\left(I+D_{\alpha}^{2}\right)^{-\frac{1}{2}}\boldsymbol{V},$ (8)
-where: $\boldsymbol{U}\sim N_{p}(0,I_{p})$ and $\boldsymbol{V}\sim HN_{p}(0,I_{p})$, where $N_{p}(0,I_{p})$ is a $p$-dimensional Normal distribution, and $HN_{p}(0,I_{p})$ is a $p$-dimensional half-Normal distribution.
+We employ the stochastic representation of the unified skew-elliptical distribution, characterized by a closed skew-normal distribution $\boldsymbol{X}\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ (Proposition 2.5), to introduce the elliptical skew-(S)KEPTIC estimators for constructing skew-elliptical graphical models.
+The closed skew-normal distribution in (5) can be expressed in the following stochastic form:
+$\boldsymbol{X}\stackrel{{\scriptstyle d}}{{=}}L^{-1}D_{\kappa}^{-\frac{1}{2}}D_{\alpha}\left(I+D_{\alpha}^{2}\right)^{-\frac{1}{2}}\boldsymbol{U}+L^{-1}D_{\kappa}^{-\frac{1}{2}}\left(I+D_{\alpha}^{2}\right)^{-\frac{1}{2}}\boldsymbol{V},$ (7)
+where: $\boldsymbol{U}\sim HN_{p}(0,I_{p})$ and $\boldsymbol{V}\sim N_{p}(0,I_{p})$, where $HN_{p}(0,I_{p})$ is a $p$-dimensional half-normal distribution, and $N_{p}(0,I_{p})$ is a $p$-dimensional normal distribution.
 The distributions of $\boldsymbol{U}$ and $\boldsymbol{V}$ are independent.
-The first two moments of $\boldsymbol{Z}$ are:
-$\mathbb{E}(\boldsymbol{Z})$ $=\mu+L^{\top}D_{\kappa}^{-\frac{1}{2}}D,$
-[…]
+The first two moments of $\boldsymbol{X}$ are:
+${E}(\boldsymbol{X})$ $=L^{\top}D_{\kappa}^{-\frac{1}{2}}D,$
+$\text{cov}(\boldsymbol{X})$ $=\left(L^{\top}D_{\kappa}^{\frac{1}{2}}(I-D^{2})D_{\kappa}^{\frac{1}{2}}L\right)^{-1},$ (8)
 where $D=\sqrt{\frac{2}{\pi}}D_{\alpha}(I_{p}+D^{2}_{\alpha})^{-\frac{1}{2}}$.
-Lemma 3.4.
-[…]
-$\boldsymbol{X}\sim\operatorname{SU}\mathcal{E}_{p,p}(\xi,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$ reduces to a $\operatorname{SUT}$ distribution: $\boldsymbol{X}\sim\operatorname{SUT}_{p,p}(\xi,\Sigma,\Lambda,\nu,0,\Gamma)$.
 Theorem 3.4 (elliptical skew-KEPTIC estimator).
-Assuming $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},0,\Lambda;f)$, such that $\boldsymbol{Z}=f(\boldsymbol{X})$ denotes a generic $\operatorname{SU\mathcal{E}}$ distribution: $\boldsymbol{Z}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Sigma^{0},\Lambda,h^{(2p)},0,\Gamma)$.
+Assume that $\boldsymbol{X}$ follows a meta skew-elliptical distribution: $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa};f)$, such that $\boldsymbol{Y}=f(\boldsymbol{X})$ denotes the unified skew-elliptical distribution $\boldsymbol{Y}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa})$.
 The elliptical skew-KEPTIC estimator is:
 ${S}^{\tau,\alpha}_{ij}=\begin{cases}2\operatorname{sin}\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)&i\neq j\\ 1&i=j\end{cases}.$
-As discussed in Hult and Lindskog, (2002), Spearman’s rho is not invariant in the class of elliptical distributions, and its relation with linear correlation coefficient in Lemma 2.1 holds if and only if the distribution is nonparanonrmal.
+As discussed in Hult and Lindskog, (2002), Spearman’s rho is not invariant in the class of elliptical distributions, and its relation with the Pearson correlation coefficient in Lemma 2.1 holds if and only if the distribution is nonparanormal.
+The proof of the Theorem is in Appendix A.
+Lemma 3.5.
+(elliptical skew-SKEPTIC estimator) Assume that $\boldsymbol{X}$ follows a meta skew-elliptical distribution: $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa};f)$, such that $\boldsymbol{Y}=f(\boldsymbol{X})$ denotes the unified skew-elliptical distribution.
+When $\boldsymbol{Y}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa})$ reduces to the CSN distribution in (5), the elliptical skew-SKEPTIC estimator is:
+${S}^{\rho,\alpha}_{ij}=\begin{cases}4\sin\left(\frac{\pi}{6}A_{{\rho}_{ij}}\right)&\text{if }i\neq j\\ 1&\text{if }i=j\end{cases},\quad\quad{S}^{\tau,\alpha}_{ij}=\begin{cases}2\operatorname{sin}\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)&i\neq j\\ 1&i=j\end{cases}.$
+We generally refer to the estimators in Theorem 3.4 and Lemma 3.5 as elliptical skew-(S)KEPTIC estimators.
+To obtain the correlation matrix using the elliptical skew-(S)KEPTIC estimators, we need to estimate the parameters of the CSN distribution.
+The estimation procedure is based on the EM algorithm of Abe et al., (2021) [^2] The estimated parameters come from the skew-normal distribution of Azzalini and Capitanio, (1999), with an adequate parametrization explained in Arellano-Valle and Azzalini, (2006) we conform the parameters to a CSN distribution..
+4 Methods for the Estimation of the Precision Matrix and the Undirected Graph
+4.1 The Dantzig selector
+The Dantzig selector estimator (Yuan,, 2010) exploits the relation between the coefficients $\theta$ stemming from a multivariate linear regression and the elements of the precision matrix.
+A sparse precision matrix implies sparse regression coefficients estimation that relies on an algorithm that depends on the tuning parameter $\delta$.
+The Dantzig procedure involves two steps:
+  • Estimation: for $j=1,\ldots,p$, calculate
+[…]
+[…]
+  • Symmetrization:
+    $\hat{\Omega}^{gDS}=\text{arg}\min_{\Omega=\Omega^{T}}\|\Omega-\hat{\Omega}\|_{1}.$ (10)
+4.2 The Glasso Algorithm
+Glasso algorithm is developed by Friedman et al., (2008) and it consists in adding a $L_{1}$-norm penalty term in the maximum likelihood estimation of $\Omega$ as follows:
+$\Omega^{glasso}=arg\max_{\Omega\succ 0}\left\{\text{log}\ \text{det}\Omega-\text{tr}(S\Omega)-\lambda_{G}\|\Omega\|_{1}\right\}.$ (11)
+The penalty term, $\lambda_{G}$, is the parameter that controls the sparsity of the estimated matrix, i.e., the number of zeros in the precision matrix.
+Banerjee et al., (2008) shows that the shrinking procedure based on the tuning parameter $\lambda_{G}$ is directly proportional to the input covariance matrix.
+This means that greater values of $\lambda_{G}$ imply a greater presence of zeros in the estimated $\Omega$, therefore, a minor number of interconnections are unveiled.
+4.3 The CLIME Algorithm
+CLIME stands for constrained $\ell_{1}$-minimization for inverse matrix estimation.
+This algorithm, proposed by Cai et al., (2011), is very attractive for high-dimensional data since the estimator is obtained by combining vector solutions by solving a linear program.
+Let $\Omega$ denote a $p\times p$ precision matrix and $\{\hat{\Omega}_{1}\}$ be the solution set of the optimization problem:
+$\text{min}\|\Omega\|_{1}\quad\text{subject to:}$ (12)
+$\|\Sigma\Omega-I\|_{\infty}\leq\lambda_{C},$
+where $\lambda_{C}$ is the tuning parameter. The final CLIME estimator of $\Omega$ is defined by the symmetrization of $\hat{\Omega}_{1}$:
+$\hat{\Omega}=(\hat{\omega}_{ij}),\quad\text{where:}$ (13)
+$\hat{\omega}_{ij}=\hat{\omega}_{ji}=\hat{\omega}^{1}_{ij}I\{|\hat{\omega}^{1}_{ij}\leq|\hat{\omega}^{1}_{ij}|\}+\hat{\omega}^{1}_{ji}I\{|\hat{\omega}^{1}_{ji}\leq|\hat{\omega}^{1}_{ji}|\}.$
+The CLIME estimator has the property that the optimization problem defined in (12) can be decomposed into $p$ optimization problems defined as:
+$\min\|\boldsymbol{\beta}\|_{1}\quad\text{subject to}\quad\|\Sigma\boldsymbol{\beta}-\boldsymbol{e}_{i}\|_{\infty}\leq\lambda_{n},$ (14)
+where $\boldsymbol{e}_{i}$ is defined as a standard unit vector in $\mathbb{R}^{p}$ with 1 in the $i$-th coordinate and 0 in the all other coordinates and $\boldsymbol{\beta}$ is a vector in $\mathbb{R}^{p}$.
+Solving (14) $p$ times, we obtain the CLIME estimator: $\{\hat{\Omega}_{1}\}=\{\hat{B}\}:=\{(\hat{\boldsymbol{\beta}}_{1},...,\hat{\boldsymbol{\beta}}_{p})\}$.
+5 Theoretical properties
+We investigate the theoretical properties of the elliptical skew-(S)KEPTIC estimators, establishing concentration inequalities in $\|\cdot\|_{\max}$ for our correlation matrix estimators, and the sparsistency of $\hat{\Omega}$ which describes the characteristic where any parameter that is originally zero is correctly estimated as zero, with a probability approaching one.
+For the rate of convergence of the correlation matrix estimator, we revisit the results of Liu et al., 2012a .
+For the corresponding precision matrix, we follow the framework introduced in Lam and Fan, (2009) for demonstrating the rate of convergence and the sparsistency of $\hat{\Omega}$.
+Our findings are based on the application of shrinkage procedures incorporating an $L_{1}$-penalty, with the three methods outlined in Section 4 adhering to this approach.
+Theorem 5.1.
+Let $x_{1},...,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha},h^{(2p)},0,D_{\kappa};f)$, restricted to the case $h^{(p)}=\phi^{(p)}(u)$, and let $\hat{S}^{\rho,\alpha}_{ij}=4\sin(\frac{\pi}{6}A_{\hat{\rho,}_{ij}})\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ denote the elliptical skew-SKEPTIC, then with probability at least $1-\frac{1}{p}$ we have:
+$\|\hat{S}^{\rho,\alpha}_{ij}-\Sigma^{0}_{ij}\|_{\max}\leq\sqrt{\frac{3(2\pi-3)^{2}}{(9-2\pi)^{2}}\frac{\log\left({p}\right)}{n}}.$ (15)
+Theorem 5.2.
+Let $x_{1},...,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha},h^{(2p)},0,D_{\kappa};f)$ and let $\hat{S}^{\tau,\alpha}_{ij}=2\sin(\frac{\pi}{2}A_{\hat{\tau}_{ij}})\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ denote the elliptical skew-(S)KEPTIC estimators, then with probability at least $1-\frac{1}{p}$ we have:
+$\|\hat{S}^{\tau,\alpha}_{ij}-\Sigma^{0}_{ij}\|_{\max}\leq\sqrt{{6\pi^{2}}\frac{\log\left({p}\right)}{n}}.$ (16)
+The proofs of Theorems 5.1 and 5.2 are in Appendix A.
+The elliptical skew-(S)KEPTIC estimators are consistent, so when the sample size $n$ increases, the estimator converges to the true correlation matrix.
+Let $\mathcal{S}=\{(i,j):\omega^{0}_{ij}\neq 0\}$ be the support set of off-diagonal nonzero entries, with cardinality $|\mathcal{S}|=s+p$, let $p_{\lambda}(\cdot)$ be a penalty function, and define
+$a=\max_{(i,j)\in\mathcal{S}}p^{\prime}_{\lambda}(|\ell_{ij}|),\quad b=\max_{(i,j)\in\mathcal{S}}p^{\prime\prime}_{\lambda}(|\ell_{ij}|).$
+where $\ell_{ij}$ is the $(i,j)$ element of the $L$ matrix that stems from $\Omega=L^{\top}ML$.
+In this work, we consider shrinkage procedures based on $L_{1}$-penalty, which means $a=\lambda$ and $b=0$.
+We impose the following regularity conditions:
+  • (a)
+    The eigenvalues of $\Sigma_{0}$ are bounded: $0<\tau_{1}<\lambda_{\min}(\Sigma_{0})\leq\lambda_{\max}(\Sigma_{0})<\tau_{2}<\infty$.
+  • (b)
+    Assume $a=O\left(\sqrt{\frac{(1+p/(s+1))\log p}{n}}\right)$, $b=o(1)$, and $\min_{(i,j)\in\mathcal{S}}|\omega^{0}_{ij}|/\lambda\to\infty$.
+  • (c)
+    The penalty is singular at the origin: $\lim_{t\downarrow 0}p_{\lambda}(t)/(\lambda t)=k>0$.
+  • (d)
+    For some constants $C,D>0$, when $\theta_{1},\theta_{2}<C\lambda$,
+    $|p^{\prime\prime}_{\lambda}(\theta_{1})-p^{\prime\prime}_{\lambda}(\theta_{2})|\leq D|\theta_{1}-\theta_{2}|.$
+Theorem 5.3.
+Suppose that conditions (a), (b), (c), (d) hold.
+If $\frac{\log p}{n}=O(\lambda^{2})$, then there exists a local minimizer for the maximum likelihood estimator for $\widehat{{M}}$ and an estimate of $\widehat{{L}}$ that satisfy:
+$\|\widehat{{M}}-{M}\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right),\quad\|\widehat{{L}}-{L}\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right),$
+which implies:
+$\|{\Omega}-\Omega\|_{\max}=O_{P}\left(s\sqrt{\frac{\log p}{n}}\right).$
+The proof of Theorem 5.3 is in Appendix A.
+We know from Lemma 1 in Zareifard et al., (2016) that $L_{ij}=0\Leftrightarrow\Omega_{ij}=0$, so the matrices $L$ and $\Omega$ share the same sparsity patter.
+For this reason, we provide sparsistency directly on $\Omega$.
+Theorem 5.4.
+(Sparsistency) Under the conditions (a), (b), (c), (d) for a local minimizer based on $L_{1}$-penalty that satisfies $\|\widehat{\Omega}-\Omega\|_{\max}^{2}=O_{P}\left\{s\log p/n\right\}$ and $\|\widehat{\Omega}-\Omega\|^{2}=O_{P}\left(\eta_{n}\right)$ for a sequence of $\eta_{n}\rightarrow 0$, where $\eta_{n}\equiv s\sqrt{\frac{\log p}{n}}$.
+If $\log p/n+\eta_{n}=O\left(\lambda_{n1}^{2}\right)$, then with probability tending to $1,\widehat{\omega}_{ij}=0$ for all $(i,j)\in S^{c}$.
+The proof of Theorem 5.4 is in Appendix A.
+6 Experiments
+6.1 Numerical Simulations
+We perform numerical simulations by generating synthetic data from the normal distribution, the closed skew-normal distribution, and the unified skew-$t$ distribution.
+For each dataset, we compute both the SKEPTIC and the skew-(S)KEPTIC estimators, and we evaluate the graph recovery performance of the proposed methods.
+The implementation follows these steps.
+Using the R package huge (Jiang et al., (2021)), we generate the precision matrix $\Omega=(\Sigma^{0})^{-1}$ using the ”random” structure, in which we randomly generate pairs of conditional dependencies in the off-diagonal elements.
+We sample $n$ observations from $p$-variate normal, CSN and SUT distribution.
+We set the location parameter equal to zero, and $\Sigma^{0}=\Omega^{-1}$.
+We consider the following cases: $p\in\{100,250,300\}$, and $n\in\{200,1000,500\}$.
+We generate synthetic data assuming $f(\boldsymbol{X})\sim N(0,\Sigma^{0})$, $f(\boldsymbol{X})\sim\operatorname{CSN}(0,\Sigma^{0},D_{\alpha}L,0,D_{\kappa})$, and $f(\boldsymbol{X})\sim\operatorname{SUT}(0,\Sigma^{0},D_{\alpha}L,0,D_{\kappa})\equiv\frac{W}{\sqrt{V_{0}}}$, where $W$ is distributed according to a closed skew-normal distribution, and $V_{0}\sim Gamma(\nu/2,\nu/2)$ as in Proposition 2.5.
+We set $\nu=3$ to emphasize the heavy tails in the distribution.
+To generate the data from the closed skew-normal distribution, we rely on the stochastic representation in (7).
+Moreover, we generate the skewness parameters from $\alpha=0.4\cdot\operatorname{sign}(u-0.5)\cdot\sqrt{|u-1|},\quad\text{where }u\sim\mathcal{U}(0,1)$.
+These parameters are encoded in the diagonal of the matrix $D_{\alpha}$.
+After the data generation, we compute the skew-(S)KEPTIC and the SKEPTIC estimators, which are then plugged into the glasso algorithm in (11) to obtain $\Omega$.
+The glasso algorithm depends on a tuning parameter $\lambda_{G}$ whose the choice is performed applying the StARS method in Liu et al., (2010), and when the value is found for each dataset we consider a set of 30 tuning parameters in the range $(\lambda_{G}-0.1,\lambda_{G}+0.1)$.
+To evaluate the accuracy of the methods, we use the receiver operating curve ($\operatorname{ROC}$).
+We define the false positive rate and true positive rate for a fixed set of tuning parameters $\lambda_{G}$.
+When the tuning parameter is determined we estimate the graph: $\hat{G}^{\lambda_{G}}=(V,\hat{E}^{\lambda_{G}})$, to evaluate it, we define the number of false positives for a given $\lambda_{G}$, $\operatorname{FP}(\lambda_{G})$, as the number of edges in $\hat{E}^{\lambda_{G}}$ but not in $E$ and the number of false negatives, $\operatorname{FN}(\lambda_{G})$, as the number of edges in $E$ but not in $\hat{E}^{\lambda_{G}}$.
+We consider the rate of the number of false negatives and false positives:
+$\operatorname{FNR}(\lambda_{G}):=\frac{\operatorname{FN}(\lambda_{G})}{|E|},\quad\text{and}\quad\operatorname{FPR}(\lambda_{G}):=\operatorname{FP}(\lambda_{G})/\left[\left(\begin{array}[]{l}d\\ 2\end{array}\right)-|E|\right].$ (17)
+To demonstrate the comprehensive performance of the analyzed methods across the entire datasets, we define the ROC curve
+$(\operatorname{FNR}(\lambda_{G}),1-\operatorname{FPR}(\lambda_{G})).$ (18)
+Graphically, we show averaged ROC curves for all the considered scenarios over 100 trials, plotted using the coordinates in (18).
+Fig.
+1 shows that the skew-(S)KEPTIC estimators generally outperform the Pearson correlation approach in non-Gaussian and improve upon the standard SKEPTIC estimator.
+Adjusting for skewness, also for excess kurtosis in the case of the SUT distribution, enhances the estimation of $\Omega$, especially in settings where the distribution presents heavy tails.
+In some cases, this leads to better recovery compared to the original SKEPTIC estimator, which has been shown to be optimal in Liu et al., 2012a .
+Table 2 provides a comprehensive overview of the simulation results, presenting quantitative comparisons of various estimators.
+The average AUC values of the skew-(S)KEPTIC estimators highlight strong performance in graph recovery under non-Gaussian conditions while still maintaining reliability when applied to normally distributed data.
+These findings suggest that considering estimators that include the skewness and accommodate for heavier tails, extending the nonparanormal distribution, do not hinder the recovery of the $\Omega$ structure.
+On the contrary, they may offer a practical advantage due to the robustness across different data distributions.
+Figure 1: Comparison of average ROC curves, using the glasso algorithm, of the elliptical skew-(S)KEPTIC estimators with SKEPTIC estimators.
+Results are determined over 50 trials.
+We consider, by row, three different data generations and, by column, three different dataset sizes.
+Table 2: Quantitative comparisons of the 5 correlation matrix estimators using different distributions to generate data, and compute the $\Omega$ matrix.
+The graphs are estimated using the glasso algorithm.
+The values are in percentages.
+SKEPTIC ($\hat{\rho}$) SKEPTIC ($\hat{\tau}$) SKEW-SKEPTIC ($\hat{\rho}$) SKEW-SKEPTIC ($\hat{\tau}$) PEARSON
+$n$ $p$ AUC FPR FNR AUC FPR FNR AUC FPR FNR AUC FPR FNR AUC FPR FNR
+N DATA 200 100 85.0 2.06 27.4 85.0 2.07 27.5 92.0 11.7 4.8 92.0 12.8 3.7 86.0 1.8 25.7
+s.e (10.0) (2.6) (21.6) (10.0) (2.6) (21.6) (2.0) (7.7) (3.2) (3.0) (7.8) (2.2) (10.0) (2.3) (21.7)
+1000 250 98.3 2.7 0.7 98.0 2.7 0.7 93.0 13.2 6.3 90.5 16.7 11.8 98.5 2.4 0.6
+s.e (3.0) (5.5) (1.3) (3.0) (5.5) (1.4) (12.5) (24.6) (24.3) (16.0) (31.4) (32.3) (2.4) (4.9) (1.1)
+500 300 94.0 1.3 9.8 94.0 1.3 9.7 96.0 7.8 0.3 96.0 8.8 0.1 95.0 1.1 9.4
+s.e (7.0) (2.4) (14.4) (7.0) (2.5) (14.2) (4.0) (8.3) (0.5) (4.0) (8.7) (0.2) (7.0) (2.1) (14.1)
+CSN DATA 200 100 86.0 5.0 22.8 86.0 5.0 22.2 88.0 17.3 6.3 88.0 18.5 5.7 83.0 4.4 28.9
+s.e (10.0) (4.9) (21.1) (10.0) (4.9) (20.8) (4.0) (10.2) (2.1) (5.0) (10.2) (2.1) (10.0) (4.7) (24.1)
+1000 250 98.0 1.6 3.2 98.0 1.6 3.0 96.0 8.3 0.7 94.2 11.0 4.5 96.0 1.3 6.7
+s.e (3.0) (2.5) (6.6) (3.0) (2.5) (6.2) (5.7) (11.3) (8.1) (10.0) (20.6) (20.7) (6.0) (2.2) (12.5)
+500 300 92.0 1.5 14.7 92.1 1.6 14.2 95.0 8.3 1.7 95.0 9.4 0.9 89.5 1.2 19.7
+s.e (9.1) (2.2) (19.3) (8.9) (2.2) (18.8) (3.4) (7.8) (1.6) (4.0) (8.3) (0.6) (11.7) (1.8) (24.4)
+SUT DATA 200 100 80.0 2.9 37.4 80.0 3.0 36.7 (-) (-) (-) 89.0 15.6 7.1 71.0 2.2 54.9
+s.e (10.0) (3.8) (24.2) (10.0) (3.9) (37.4) (-) (-) (-) (2.0) (6.3) (1.7) (12.0) (3.2) (26.8)
+1000 250 97.0 2.1 4.7 97.0 2.1 4.5 (-) (-) (-) 96.0 7.7 0.1 89.0 1.7 20.3
+s.e (4.0) (4.5) (8.6) (4.0) (4.5) (8.3) (-) (-) (-) (3.7) (7.4) (0.01) (12.0) (4.2) (25.2)
+500 300 89.3 1.9 19.8 89.6 1.9 18.9 (-) (-) (-) 94.4 9.7 1.5 80.0 1.6 37.6
+s.e (9.1) (3.7) (19.8) (9.0) (3.7) (19.5) (-) (-) (-) (3.3) (6.9) (0.7) (14.4) (3.3) (30.5)
+AUC = Area Under Curve, FPR = False Positive Rate, and FNR = False Negative Rate.
+6.2 Empirical Analysis: S&P500 Log-Returns
+In the empirical analysis, we study the conditional dependence relations on daily returns of S&P500500 companies in the period from $04/01/2016$ to $31/12/2022$ downloaded from Yahoo!
+Finance, using the proposed elliptical skew-(S)KEPTIC estimators and comparing it with the SKEPTIC.
+After cleaning the data, the dataset consists of 454 stocks, each counting 1762 observations.
+The log-returns of the stocks in the S&P500 index are empirically skewed and present heavier tails than the normal distribution (Cont, (2001)).
+Figure 2: Illustration of leptokurtic distributions of daily log-returns, with a relevant skewness component. The stock data are from 04/01/2016 to 31/12/2022.
+As a motivating example, we benchmark the analysis of Han and Liu, (2014), reporting the asymmetry of log-returns in Figure 2, where we plot histograms of the returns of three stocks in our sample ”eBay Inc.”, ”Coca-Cola” and ”Wells Fargo” belonging to three different sectors, respectively, Consumer Discretionary, Consumer Staples, and Financials.
+We appreciate the departure from normality reflected in leptokurtic distribution with a relevant skewness component.
+Moreover, we investigate the Gaussian assumption of the asset log-returns by computing normality tests with different significance levels at 5% and 1%.
+Table 3 indicates a clear and complete deviation from normality in the stock data.
+This could be attributed to the presence of exceptional events, thus a huge presence of outlier values during the analyzed period, which lead to a departure from the assumption of normal distribution for log-returns.
+Table 3: Normality tests S&P500 log-returns data. This table illustrates the number of stocks that reject the null hypothesis of normality at 1% and 5% levels.
+Significance Level Kolmogorov-Smirnov Shapiro-Wilk Lilliefors
+1% 454 454 454
+5% 454 454 454
+To estimate the graphs, given the large number of variables relative to observations we determine the tuning parameter $\lambda_{G}$ using the Extended Bayesian Information Criteria (EBIC) of Chen and Chen, (2008) to ensure a balanced trade-off between model fit and complexity.
+$\operatorname{EBIC}=-2L+E\log(n)+4\gamma E\log(p).$ (19)
+The hyperparameter $\gamma\in\left[0,1\right]$, with greater values of $\gamma$ implying a stronger penalization, $L$ denotes the log-likelihood, $n$ the sample size, $E$ the number of non-zero edges and $p$ the number of nodes.
+We set $\gamma=1$, ensuring a strong penalization, as desired.
+The tuning parameter $\lambda_{G}$ is set equal to $0.54$ the value that minimizes the scores of the first-order differences in EBIC.
+To align the log-returns to the unified skew-$t$ distribution for computing the elliptical skew-KEPTIC estimator, we estimate the degrees of freedom of the distribution.
+We follow Arellano-Valle and Genton, 2010a using the maximum likelihood estimation method.
+In likelihood methods, the underlying assumption is that the observations are independent, therefore, we apply the best ARIMA model to the series using the R function auto.arima to filter our returns into a white noise series, and we apply the likelihood method to the filtered series to estimate the degrees of freedom using the command mst.mle from the R package sn developed by Azzalini, (2023).
+Figure 3 shows the estimated graphs using SKEPTIC, elliptical skew-SKEPTIC, and elliptical skew-KEPITC estimators; all the correlation matrix estimators are based on Kendall’s tau statistic.
+Figure 3: The resulting graphs of the log-returns of the S&P 500 stocks from $04/01/2016$ to $31/12/2022$.
+Kendall’s tau statistic is employed to estimate correlation.
+In the graph: (a) SKEPTIC estimator, (b) elliptical skew-SKEPTIC estimator, and (c) elliptical skew-KEPTIC estimator.
+The 454 stocks are categorized into 11 Global Industry Classification Standard (GICS) sectors, including Consumer Discretionary (51 stocks), Consumer Staples (33 stocks), Energy (21 stocks), Financials (67 stocks), Health Care (56 stocks), Industrials (69 stocks), Information Technology (59 stocks) Telecommunications Services (17 stocks), Materials (25 stocks), Real Estate (28 stocks), and Utilities (28 stocks).
+The stocks that belong to the same sector tend to cluster, and those at the periphery of the network are uncorrelated with the others during the considered period, thus, they are uncorrelated with the movement of the market.
+Peralta and Zareei, (2016) show that those stocks have a greater weight in constructing a portfolio using the Markowitz approach.
+Therefore, in large-dimensions, a network approach can be valuable for identifying stocks with the least correlation to the market.
+Table 4 demonstrates that incorporating skewness and kurtosis components leads to a sparser network estimation compared to the normal case.
+Table 4: Summary statistics of the estimated networks using Kendall’s tau statistic for the correlation matrix estimation.
+SK == SKEPTIC, SNK == elliptical skew-SKEPTIC, STK == elliptical skew-KEPTIC.
+DIM = dimension of the dataset.
+The notation ”>>” indicates that one method has a greater number of edges compared to the other.
+Network  Edges No. Edges diff.
+DIM SK STK SNK SK>>SNK SK>>STK SNK>>STK
+S&P 500 454 384 284 231 100 153 53
+7 Concluding remarks
+We propose rank-based correlation matrix estimators, named the elliptical skew-(S)KEPTIC estimators, to provide semiparametric estimation of graphical models built on meta skew-elliptical distributions.
+The proposed framework includes two classes of estimators: the elliptical skew-KEPTIC estimator which employs Kendall’s tau statistic, and the elliptical skew-SKEPTIC estimator which also uses Spearman’s rho statistic to estimate the correlation matrix.
+We study the theoretical properties of the rank-based correlation matrix estimators and establish concentration bounds.
+These bounds yield statistical rates of convergence under the elementwise maximum norm for both the correlation matrix estimation and the resulting estimation of the precision matrix.
+We also study the sparsistency of the precision matrix, evaluating the performance of the elliptical skew-(S)KEPTIC estimators in correctly recovering the edges of the graphical model.
+The empirical analysis and numerical simulations demonstrate that the elliptical skew-(S)KEPTIC estimators yield the optimal performance when the data deviates from normality.
+Applied to real data, specifically the log-returns of S&P500 constituent stocks, the resulting graphs are sparser compared to those estimated under normality.
+Our findings show a reliable framework that is useful when dealing with data that exhibit high skewness which cannot be neglected, and its effectiveness is validated through the optimal graph recoveries of the numerical simulations.
+Appendix A Proofs of elliptical skew-(S)KEPTIC estimators
+A.1 Proof of Theorem 3.4
 Proof.
-We consider $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},0,\Lambda;f)$, where $f(\boldsymbol{X})\sim V_{0}^{-1/2}Z_{0}$, and $Z_{0}\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ where $D_{\kappa}=I_{p}$, and $V_{0}=1$.
-We compute $\mathbb{C}orr(X)$ using Kendall’s tau statistic, $\Sigma^{0}_{\tau}=\left[\tau_{ij}\right]$, and we decompose $\Sigma^{0}_{\tau}=L^{\top}D_{\kappa}L$, aligning it $\Sigma^{0}_{\tau}$ to the assumption of $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},0,\Lambda;f)$ exploiting (9).
-[…]
-$=(L^{\top}ML)^{-1}\quad\text{where:}\quad M=I-D^{2},$
-$=\frac{Adj(L^{\top}ML)}{det(L^{\top}ML)},$
-$=\frac{Adj(L^{\top}ML)}{det(L)^{2}det(M)}.$
-Where $Adj$ is the adjoint matrix, and $det$ is the determinant of the matrix.
+We consider $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa};f)$, where $f(\boldsymbol{X})\sim V_{0}^{-1/2}Z_{0}$, and
+$Z_{0}\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ where $D_{\kappa}=I_{p}$, and $V_{0}=1$.
+We compute the correlation matrix of $\boldsymbol{X}$ using Kendall’s tau statistic, $\Sigma^{0}_{\tau}=\left[\tau_{ij}\right]$, and we decompose $\Sigma^{0}_{\tau}=L^{\top}D_{\kappa}L$, aligning it $\Sigma^{0}_{\tau}$ to the assumption of $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa};f)$ exploiting (8).
+$\text{cov}(\boldsymbol{X})$ $=\left(L^{\top}(I-D^{2})L\right)^{-1},$
+$=(L^{\top}ML)^{-1}\quad\text{where:}\quad M=I-D^{2}.$
 The matrices $I$ and $D^{2}$ are both diagonal matrices, and the matrix $M$ is.
@@
 $m_{kk}=I_{kk}-D^{2}_{kk}=1-\frac{2}{\pi}\frac{\alpha_{kk}^{2}}{1+\alpha_{kk}^{2}}.$
-Let $A_{{\tau}}$ denote the final correlation matrix of $X$ with Kendall’s tau statistic.
-We recall that the stochastic representation (8) of the $\operatorname{CSN}$ is a sum of two comonotic copula functions that have the invariance property, so the SKEPTIC estimator is the same for both the distributions, thus the elliptical skew-KEPTIC estimator is the sum of two SKEPTIC estimators and an estimated matrix, $A_{\tau}$, from $f(\boldsymbol{X})\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$.
-When $V_{0}\sim Gamma(\nu/2,\nu/2)$, we have that $f(\boldsymbol{X})\sim\operatorname{SUT}_{p,p}(\xi,\Sigma,\Lambda,\nu,0,\Gamma)$ or equivalently: $f(\boldsymbol{X})\sim\frac{Z_{0}}{\sqrt{V_{0}}}$.
-Since the positive support of Gamma random variable, and the fact that Kendall’s tau is invariant under monotone transformations, we state that the skew-KEPTIC estimator maintains the same ordinal measures of dependencies of CSN distribution even when $f(\boldsymbol{X})\sim\operatorname{SUT}_{p,p}(0,\Sigma^{0},\Lambda,\nu,0,\Gamma)$, where $\Gamma=I_{p}$.
+Let $A_{{\tau}}\equiv L^{\top}ML$ denote the rescaled final correlation matrix of $\boldsymbol{X}$, estimated using Kendall’s tau statistic.
+We recall that the stochastic representation (7) of the CSN is a sum of two comonotic copula functions that have the invariance property, so the SKEPTIC estimator is the same for both the distributions, thus the elliptical skew-KEPTIC estimator is the sum of two SKEPTIC estimators and an estimated matrix, $A_{\tau}$, from $f(\boldsymbol{X})\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$.
+When $V_{0}\sim Gamma(\nu/2,\nu/2)$, we have that $f(\boldsymbol{X})\sim\operatorname{SUT}_{p,p}(0,\Sigma,D_{\alpha}L,\nu,0,D_{\kappa})$ or equivalently: $f(\boldsymbol{X})\sim\frac{Z_{0}}{\sqrt{V_{0}}}$.
+Since the positive support of Gamma random variable, and the fact that Kendall’s tau is invariant under monotone transformations, we state that the skew-KEPTIC estimator maintains the same ordinal measures of dependencies of the closed skew-normal distribution even when $f(\boldsymbol{X})\sim\operatorname{SUT}_{p,p}(0,\Omega^{-1},D_{\alpha}L,\nu,0,D_{\kappa})$.
 ∎
-Lemma 3.6.
-(skew-Normal SKEPTIC estimator) Assuming $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},0,\Lambda;f)$, such that $\boldsymbol{Z}=f(\boldsymbol{X})$ denotes a generic $\operatorname{SU\mathcal{E}}$ distribution, when $\boldsymbol{Z}\sim\operatorname{SU}\mathcal{E}_{p,p}(0,\Sigma^{0},\Lambda,h^{(2p)},0,\Gamma)$ reduces to a $\operatorname{CSN}$ distribution, the skew-Normal SKEPTIC estimator is:
-${S}^{\rho,\alpha}_{ij}=\begin{cases}4\sin\left(\frac{\pi}{6}A_{{\rho}_{ij}}\right)&\text{if }i\neq j\\ 1&\text{if }i=j\end{cases},\quad\quad{S}^{\tau,\alpha}_{ij}=\begin{cases}2\operatorname{sin}\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)&i\neq j\\ 1&i=j\end{cases}.$
+A.2 Proof of Lemma 3.5
 Proof.
 We recall Lemma 2.1, where for the normal distribution, and for $i\neq j$, we have: ${S}_{ij}=2\sin\left(\frac{\pi}{6}{\rho}_{ij}\right)=\sin\left(\frac{\pi}{2}{\tau}_{ij}\right)$.
-Let $A_{\rho_{ij}}$ denote the final correlation matrix of $\boldsymbol{X}$ with Spearman’s rho statistic, as in the Normal distribution, we have: ${S}^{\rho,\alpha}_{ij}\equiv 4\sin\left(\frac{\pi}{6}A_{{\rho}_{ij}}\right)=2\sin\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)$.
+Let $A_{\rho}$ denote the final correlation matrix of $\boldsymbol{X}$ with Spearman’s rho statistic, as in the normal distribution, we have: ${S}^{\rho,\alpha}_{ij}\equiv 4\sin\left(\frac{\pi}{6}A_{{\rho}_{t}{ij}}\right)=2\sin\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)$.
 ∎
-We generally refer to the estimators in Theorem 3.5 and Lemma 3.6 as elliptical skew-(S)KEPTIC estimators.
-The specification of the inclusion of the Spearman’s rho will be clear in the application.
-To obtain the correlation matrix using the elliptical skew-(S)KEPTIC estimator, we need to estimate the parameters of the $\operatorname{CSN}$ distribution.
-The estimation procedure is based on the EM algorithm of Abe et al., (2021) [^1] The estimated parameters come from the skew-Normal distribution of Azzalini and Capitanio, (1999), with an adequate parametrization explained in Arellano-Valle and Azzalini, (2006) we conform the parameters to a CSN distribution..
-3.1 Estimating Marginal Distributions
-We find estimators $\{f_{j}\}_{j=1}^{p}$ such that the transformed data are distributed according to a $\operatorname{CSN}$.
-Let $\Psi$ denote the cumulative distribution function of the $\operatorname{CSN}$ distribution, then the estimators are:
-$\hat{f}_{csn,j}(x_{ij})=\Psi^{-1}(\hat{r}_{ij}),\quad\text{where: }\hat{r}_{ij}=\frac{r_{ij}}{n+1}.$ (11)
-The functions $\{f_{csn,j}(x_{ij})\}_{j=1}^{p}$ are monotone and differentiable, and Gupta et al., (2004) show that $\operatorname{CSN}$ distribution is closed under conditioning therefore, according to Sklar’s theorem, (6) is a $\operatorname{CSN}$ copula.
-3.2 Methods for the Estimation of the Precision Matrix and the Undirected Graph
-Upon completing the estimation of the correlation matrix, we follow Liu et al., 2012a and estimate the precision matrix capturing the interrelationships among variables within an undirected graph.
-3.2.1 The Dantzig selector
-The Dantzig selector estimator (Yuan, , 2010) exploits the relation between the coefficients $\theta$ stemming from a multivariate linear regression and the elements of the precision matrix.
-Sparse precision matrix implies sparse regression coefficients estimation that relies on an algorithm that depends on the tuning parameter $\delta$.
-Dantzig procedure involves two steps:
-  • Estimation: for $j=1,\ldots,p$, calculate
-    $\hat{\theta}^{j}$ $=\text{arg}\min_{\theta\in\mathbb{R}^{p-1}}\|\theta\|_{1}\quad\text{subject to }\quad\|\hat{S}_{\backslash j,j}-\hat{S}_{\backslash j,\backslash j}\theta\|_{\infty}\leq\delta,$ (12)
-    $\hat{\Omega}_{jj}$ $=\left[1-2(\hat{\theta}^{j})\hat{S}_{\backslash j,j}+(\hat{\theta}^{j})^{\top}\hat{S}_{\backslash j,\backslash j}\hat{\theta}^{j}\right]\quad\text{and }\quad\hat{\Omega}_{\backslash j.j}=-\hat{\Omega}_{\backslash jj}\hat{\omega}^{j}.$
-  • Symmetrization:
-    $\hat{\Omega}^{gDS}=\text{arg}\min_{\Omega=\Omega^{T}}\|\Omega-\hat{\Omega}\|_{1}.$ (13)
-3.2.2 The Glasso Algorithm
-Glasso algorithm is developed by Friedman et al., (2008) and it consists in adding a $L_{1}$-norm penalty term in the maximum likelihood estimation of $\Omega$ as follows:
-$\Omega^{glasso}=arg\max_{\Omega\succ 0}\left\{\text{log}\ \text{det}\Omega-\text{tr}(S\Omega)-\lambda_{G}\|\Omega\|_{1}\right\}.$ (14)
-The penalty term, $\lambda_{G}$, is the parameter that controls the sparsity of the estimated matrix, i.e., the number of zeros in the precision matrix.
-Banerjee et al., (2008) shows that the shrinking procedure based on the tuning parameter $\lambda_{G}$ is directly proportional to the input covariance matrix.
-This means that greater values of $\lambda_{G}$ imply a greater presence of zeros in the estimated $\Omega$, therefore a minor number of interconnections are unveiled.
-3.2.3 The CLIME Algorithm
-CLIME stands for constrained $\ell_{1}$-minimization for inverse matrix estimation.
-This algorithm, proposed by Cai et al., (2011), is very attractive for high-dimensional data since the estimator is obtained by combining vector solutions solving a linear program.
-Let $\Omega$ denote a $p\times p$ precision matrix and $\{\hat{\Omega}_{1}\}$ be the solution set of the optimization problem:
-$\text{min}\|\Omega\|_{1}\quad\text{subject to:}$: (15)
-$\|\Sigma_{n}\Omega-I\|_{\infty}\leq\lambda_{C},$
-where $\lambda_{C}$ is the tuning parameter. The final CLIME estimator of $\Omega_{0}$ is defined by the symmetrization of $\hat{\Omega}_{1}$:
-$\hat{\Omega}=(\hat{\omega}_{ij}),\quad\text{where:}$: (16)
-$\hat{\omega}_{ij}=\hat{\omega}_{ji}=\hat{\omega}^{1}_{ij}I\{|\hat{\omega}^{1}_{ij}\leq|\hat{\omega}^{1}_{ij}|\}+\hat{\omega}^{1}_{ji}I\{|\hat{\omega}^{1}_{ji}\leq|\hat{\omega}^{1}_{ji}|\}.$
-The CLIME estimator has the property that the optimization problem defined in (15) can be decomposed into $p$ optimization problems defined as:
-$\min\|\boldsymbol{\beta}\|_{1}\quad\text{subject to}\quad\|\Sigma\boldsymbol{\beta}-\boldsymbol{e}_{i}\|_{\infty}\leq\lambda_{n},$ (17)
-where $\boldsymbol{e}_{i}$ is defined as a standard unit vector in $\mathbb{R}^{p}$ with 1 in the $i$-th coordinate and 0 in the all other coordinates and $\boldsymbol{\beta}$ is a vector in $\mathbb{R}^{p}$.
-Solving (17) $p$ times, we obtain the CLIME estimator: $\{\hat{\Omega}_{1}\}=\{\hat{B}\}:=\{(\hat{\boldsymbol{\beta}}_{1},...,\hat{\boldsymbol{\beta}}_{p})\}$.
-4 Theoretical properties
-We analyze the theoretical properties of the elliptical skew-(S)KEPTIC estimator.
-The results suggest that our estimator is reliable when dealing with highly skewed data.
-As Liu et al., 2012a , our results show an exponential concentration rate in $\|\cdot\|_{\max}$.
-Moreover, we prove optimal bounds for shrinkage procedures.
-We show the results for the Dantzig selector, however, they can be easily extended to the CLIME and the Glasso procedures.
-Theorem 4.1.
-Let $x_{1},...,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim\operatorname{MS\mathcal{E}}_{p}(\Sigma^{0},\xi,\Lambda;f)$, restricted to the case $h^{(p)}=\phi^{(p)}(u)$, and let $\hat{S}^{\rho,\alpha}_{ij}=4\sin(\frac{\pi}{6}A_{\hat{\rho,}_{ij}})\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ denote the elliptical skew-SKEPTIC, then with probability at least $1-1/p^{2}$ we have:
-$\|\hat{S}^{\rho,\alpha}_{ij}-\Sigma^{0}_{ij}\|_{\max}\leq 16\pi\sqrt{\frac{logp}{n}}.$ (18)
-Theorem 4.2.
-Let $x_{1},...,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim\operatorname{MS\mathcal{E}}_{p}(\Sigma^{0},\xi,\Lambda;f)$ and let $\hat{S}^{\tau,\alpha}_{ij}=2\sin(\frac{\pi}{2}A_{\hat{\tau}_{ij}})\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ denote the elliptical skew-(S)KEPTIC estimator, then with probability at least $1-1/p^{2}$ we have:
-$\|\hat{S}^{\tau,\alpha}_{ij}-\Sigma^{0}_{ij}\|_{\max}\leq 4\pi\sqrt{\frac{logp}{n}}.$ (19)
-Proofs of Theorems 4.1 and 4.2 are in Appendix A. The elliptical skew-(S)KEPTIC estimator achieves convergence in high-dimension, thus it is possible to define the following result.
-Theorem 4.3.
-When we substitute the estimated matrices $\hat{S}^{\rho,\alpha}$ and $\hat{S}^{\tau,\alpha}$ into the parametric graphical lasso (or the CLIME or graphical Dantzig selector), the estimators provide a slower convergence than the Gaussian case ensuring consistency and graph recovery as a Gaussian model.
+A.3 Proof of Theorem 5.1
 Proof.
-The proof is based on the observation that the sample correlation matrix $\hat{S}$ is a sufficient statistic for all three shrinkage procedures: the graphical lasso, graphical Dantzig selector, and CLIME.
-As Yuan, (2010) points out, a sufficient condition to ensure that an estimator is consistent is if there exists some constant $c$, such that $\mathbb{P}\left(\|\hat{S}-\Sigma^{0}\|_{\max}>c\sqrt{\frac{\log p}{n}}\right)\leq 1-\frac{1}{p}$, which can be replaced by (18) and (19) in Theorems 4.1 and 4.2.
-∎
-The high-dimensional convergence of the elliptical skew-(S)KEPTIC estimator also means that the precision matrix is bounded.
-Similar results are found by Cai et al., (2011) and Yuan, (2010).
-In particular, let $\hat{\Omega}^{mse-s}$ denote the inverse correlation matrix using the elliptical skew-(S)KEPTIC estimator and consider the Dantzig selector as the shrinkage procedure to find $\hat{\Omega}$.
-We introduce a class of inverse matrix defined as $\mathcal{M}_{1}(\kappa,\tau,M):=\{\Omega:\Omega\succ 0,\text{diag}(\Omega^{-1})=1,\|\Omega\|_{1}\leq\kappa,\frac{1}{\tau}\leq\lambda_{\min}(\Omega)\leq\lambda_{\max}(\Omega)\leq\tau,\deg(\Omega)\leq M\}$, where $\kappa,\tau>1$.
-Given Theorem 4.3, we obtain the following result:
-Theorem 4.4.
-For $1\leq q<\infty$, there exists a constant $c$ that depends on $\kappa$, $\tau$, $\lambda_{\min}(\Omega)$ and $\lambda_{\max}(\Omega)$, such that
-$\sup_{\Omega\in\mathcal{M}_{1}(\kappa,\tau,M)}\|\hat{\Omega}^{mse-s}-\Omega\|_{q}=O_{P}\left(M\sqrt{\frac{\log p}{n}}\right).$ (20)
-provided that $\lim_{n\rightarrow\infty}\frac{n}{M^{2}\log d}=\infty$ and $\delta=c\sqrt{\frac{\log d}{n}}$, for sufficiently large $c$, where $\delta$ is the parameter in (12).
+We have: $\hat{S}^{\rho,\alpha}_{ij}=$ $4\sin\left(\frac{\pi}{6}A_{\hat{\rho}_{ij}}\right)$ and we introduce $\Sigma^{\rho}$, the correlation matrix using Spearman’s rho statistic.
+We know that $A_{\rho}=L^{\top}ML$, where $M=I-D^{2}$.
+By Definition: $\Sigma^{\rho}=L^{\top}D_{\kappa}L$.
+This implies:
+$\left\lVert\widehat{D}_{\kappa}\right\rVert_{\max}\geq\left\lVert\widehat{M}\right\rVert_{\max},$
+$\left\lVert L^{\top}\widehat{D}_{\kappa}L\right\rVert_{\max}\geq\left\lVert L^{\top}\widehat{M}L\right\rVert_{\max},$
+$\left\lVert\widehat{\Sigma}^{\rho}\right\rVert_{\max}\geq\left\lVert A_{\widehat{\rho}}\right\rVert_{\max},$
+$\left\lVert 4\sin\left(\frac{\pi}{6}\widehat{\Sigma}^{\rho}\right)\right\rVert_{\max}\geq\left\lVert 4\sin\left(\frac{\pi}{6}A_{\widehat{\rho}}\right)\right\rVert_{\max}.$
+[…]
+$E(\widehat{\rho}_{ij})=\frac{6}{\pi(n+1)}\left[\arcsin\left(\Sigma_{ij}^{0}\right)+(n-2)\arcsin\left(\frac{\Sigma_{ij}^{0}}{2}\right)\right].$ (20)
+We rewrite (20) as:
+$\Sigma^{0}_{ij}=2\sin\left(\frac{\pi}{6}E(\hat{\rho}_{ij})+a_{ij}\right),\quad\text{where }a_{ij}=\frac{\pi E(\hat{\rho}_{ij})-2\arcsin(\Sigma^{0}_{ij})}{2(n-2)}.$
+The term $|a_{ij}|\leq\frac{\pi}{n-2}$. Therefore, for all $n>\frac{6\pi}{t}+2$, we have $|a_{ij}|<\frac{t}{6}$, and thus:
+$\text{pr}\left(\left\lVert 4\sin\left(\frac{\pi}{6}\hat{\Sigma}^{\rho}\right)-\Sigma^{0}\right\rVert_{\max}>t\right)$ $=\text{pr}\left(\max_{ij}\left|4\sin\left(\frac{\pi}{6}\hat{\rho}_{ij}\right)-2\sin\left(\frac{\pi}{6}E(\hat{\rho}_{ij})+a_{ij}\right)\right|>t\right)$
+$=\text{pr}\left(\max_{ij}\left|2\sin\left(\frac{\pi}{6}\hat{\rho}_{ij}\right)-\sin\left(\frac{\pi}{6}E(\hat{\rho}_{ij})+a_{ij}\right)\right|>\frac{t}{2}\right)$
+$\leq\text{pr}\left(\max_{ij}\frac{2\pi}{3}\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})-a_{ij}\right|>\frac{t}{2}\right)\quad\text{(Lipshitz function)}$
+$\leq\text{pr}\left(\max_{ij}\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})\right|>\frac{3t}{4\pi}-|a_{ij}|\right)$
+$\leq\text{pr}\left(\max_{ij}\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})\right|>\frac{3t}{4\pi}-\frac{t}{6}\right)$
+$=\text{pr}\left(\max_{ij}\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})\right|>\frac{(9-2\pi)t}{(2\pi-3)}\right)$
+$\leq\sum_{ij}\text{pr}\left(\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})\right|>\frac{(9-2\pi)t}{(2\pi-3)}\right)\quad\text{(Boole's inequality)}$
+$\leq p^{2}\exp\left(-\frac{nt^{2}(9-2\pi)^{2}}{(2\pi-3)^{2}}\right)\quad\text{(Hoeffding's inequality)}$
+Setting:
+$\beta=p^{2}\exp\left(\frac{(9-2\pi)^{2}nt^{2}}{(2\pi-3)^{2}}\right)$
+and solving for $t$, we get:
+$t=\sqrt{\frac{(2\pi-3)^{2}}{(9-2\pi)^{2}n}\log\left(\frac{p^{2}}{\beta}\right)}$
+When $\beta=\frac{1}{p}$, we have $\sqrt{\frac{(2\pi-3)^{2}}{(9-2\pi)^{2}n}\log\left({p^{3}}\right)}$. ∎
+A.4 Proof of Theorem 5.2
 Proof.
-This result follows Yuan, (2010), which proves that the difference between the estimated precision matrix and the theoretical precision matrix is bounded by $\|\hat{S}-\Sigma^{0}\|_{\max}$, where $\hat{S}$ is the estimated correlation matrix and $\Sigma^{0}$ is the theoretical correlation matrix.
-In our case, to prove the bounds for the estimated correlation matrix we just replace $\hat{S}$ with $\hat{S}^{\rho,\alpha}$ and $\hat{S}^{\tau,\alpha}$ where the bounds are proved in Theorem 4.1 and 4.2.
-∎
-5 Experiments
-We present the empirical performance of the elliptical skew-(S)KEPTIC methods on both synthetic and real datasets, comparing the proposed method throughout this paper with the SKEPTIC estimator in Liu et al., 2012a .
-In particular, we consider the following methods for the estimation of graphical models.
-We generate data from the Normal distribution, the CSN distribution, and the SUT distribution.
-For each generation, we compute the SKEPTIC and the SKEW-(S)KEPTIC estimator and conduct numerical simulations on synthetic data to show the ability of graph recovery of the proposed estimators.
-In conclusion, we present an empirical analysis where the methods are applied to the log-returns of the S&P500 index.
-5.1 Numerical Simulations
-The numerical simulations follow these steps.
-Using the R package huge (Jiang et al., (2021)), we generate the precision matrix $\Omega=(\Sigma^{0})^{-1}$ using the ”random” structure, in which we randomly generate pairs of conditional dependencies in the off-diagonal elements.
-We sample $n$ observations from $p$-variate Normal, $\operatorname{CSN}$ and $\operatorname{SUT}$ distribution.
-We set the location parameter equal to zero for the last two, and $\Sigma^{0}=\Omega^{-1}$.
-We consider the following cases: $p\in\{100,100,200\}$, and $n\in\{200,100,100\}$.
-The cases $n=100,p=100$, and $n=100,p=200$ consider the high-dimensional scenarios.
-Regarding the $\operatorname{CSN}$ parameters, we set $\Delta=1$ and the skewness matrix is generated considering the off-diagonal elements equal to zero, whereas the diagonal matrix is generated from $\operatorname{Uniform}(-0.5,0.5)$.
-We generate the $\operatorname{SUT}$ distribution with degrees of freedom coming from $\operatorname{Uniform}(2,10)$ [^2] The range is determined by estimating the degrees of freedom of the data on stock returns.
-.
-The estimators are then plugged into the ”Glasso” algorithm in (14) to obtain $\Omega$.
-To evaluate the accuracy of the methods, we use the receiver operating curve ($\operatorname{ROC}$).
-We define the false positive rate and true positive rate for a fixed set of tuning parameters $\lambda_{G}$.
-Let $\hat{G}^{\lambda}_{G}=(V,\hat{E}^{\lambda_{G}})$ be an estimated graph using the proposed procedure and a given value of $\lambda_{G}$.
-We define the number of false positives for a given $\lambda_{G}$, $\operatorname{FP}(\lambda_{G})$, as the number of edges in $\hat{E}^{\lambda_{G}}$ but not in $E$ and the number of false negatives, $\operatorname{FN}(\lambda_{G})$, as the number of edges in $E$ but not in $\hat{E}^{\lambda_{G}}$.
-We consider the rate of the number of false negative and false positive:
-$\operatorname{FNR}(\lambda_{G}):=\frac{\operatorname{FN}(\lambda_{G})}{|E|},\quad\text{and}\quad\operatorname{FPR}(\lambda_{G}):=\operatorname{FP}(\lambda_{G})/\left[\left(\begin{array}[]{l}d\\ 2\end{array}\right)-|E|\right].$ (21)
-To demonstrate the comprehensive performance of the analyzed methods across the entire datasets, we define the ROC curve
-$(\operatorname{FNR}(\lambda_{G}),1-\operatorname{FPR}(\lambda_{G})).$ (22)
-Graphically, we show averaged ROC curves for all the considered scenarios over 100 trials, plotted using the coordinates in (22).
-Fig.
-1 shows numerical simulations of the average ROC curves of the elliptical skew-(S)KEPTIC estimators and the SKEPTIC estimators, which we employ to estimate $\Omega$ using the Glasso algorithm.
-The estimators exhibit similar performance, consequently yielding an optimal graph recovery in the simulation results, even when accounting for skewness and kurtosis components in the data, thus giving reliability to the elliptical skew-(S)KEPTIC estimators.
-Table 2 presents the quantitative comparisons, showing that regardless of the data generation process or the estimator used, the structure of $\Omega$ can be effectively recovered.
-Incorporating a skewness component and accounting for heavier tails than those in the NPN distribution do not hinder the recovery of $\Omega$’s structure.
-Table 2: Quantitative comparisons of the 4 estimators using different distributions to generate data, and compute the $\Omega$ matrix.
-The graphs are estimated using the Glasso algorithm.
-The values are in percentage.
-SKEPTIC ($\hat{\rho}$) SKEPTIC ($\hat{\tau}$) SKEW-SKEPTIC ($\hat{\rho}$) SKEW-SKEPTIC ($\hat{\tau}$) $n$ $p$ FPR FNR FPR FNR FPR FNR FPR FNR N DATA 200 100 11.11 74.3 11.2 74.3 11.8 74.7 11.4 74.7 s.e (1.3) (3.5) (1.3) (3.5) (1.4) (3.5) (1.3) (3.5) 100 100 15.5 66.9 15.4 67.0 16.1 67.1 15.7 67.2 s.e (1.6) (2.4) (1.0) (2.4) (1.1) (2.4) (1.1) (2.4) 100 200 13.8 64.0 14.0 64.2 14.9 64.0 14.5 64.1 s.e (1.0) (2.4) (1.0) (2.4) (1.1) (2.4) (1.1) (2.4) CSN DATA 200 100 11.5 73.9 11.5 74.0 12.1 74.1 11.6 73.7 s.e (1.4) (3.5) (1.4) (3.5) (3.5) (3.5) (1.4) (3.5) 100 100 15.6 67.2 15.8 67.3 16.0 65.8 15.3 65.1 s.e (1.6) (3.4) (1.6) (3.4) (1.7) (3.5) (1.6) (3.5) 100 200 13.9 64.0 14.1 64.1 14.3 62.1 13.6 61.6 s.e (1.0) (2.4) (1.0) (2.4) (1.1) (2.5) (1.0) (2.5) SUT DATA 200 100 12.6 72.6 12.8 72.8 - - 14.9 61.0 s.e (1.5) (3.5) (1.5) (3.5) (-) (-) (1.7) (3.3) 100 100 16.8 64.4 17.0 64.5 - - 17.1 64.5 s.e (1.7) (3.3) (1.7) (3.3) (-) (-) (1.7) (3.3) 100 200 14.5 60.7 14.5 61.0 - - 14.9 61.0 s.e (1.0) (2.4) (1.0) (2.4) (-) (-) (1.1) (2.4)
-Figure 1: Comparison of average ROC curves, using the Glasso algorithm, of the elliptical skew-(S)KEPTIC estimators with SKEPTIC estimators.
-Results are determined over 100 trials.
-We consider, by column, three different data generations.
-And, by row, three different dataset sizes.
-5.2 Empirical Analysis: S&P500 Log-Returns
-In the empirical analysis, we study the conditional dependence relations on daily returns of S&P500500500500 companies in the period from $04/01/2016$ to $31/12/2022$ downloaded from Yahoo!
-Finance, using the proposed elliptical skew-(S)KEPTIC estimator and comparing it with the SKEPTIC.
-After cleaning the data, the dataset consists of 454 stocks, each counting 1762 observations.
-The log-returns of the stocks in the S&P500 index are empirically skewed and present heavier tails than the Normal distribution (Cont, (2001)).
-Figure 2: Illustration of leptokurtic distributions of daily log-returns, with a relevant skewness component. The stock data are from 04/01/2016 to 31/12/2022.
-As a motivating example, we benchmark the analysis of Han and Liu, (2014), reporting the asymmetry of log-returns in Figure 2, where we plot histograms of the returns of three stocks in our sample ”eBay Inc.”, ”Coca-Cola” and ”Wells Fargo” belonging to three different sectors, respectively, Consumer Discretionary, Consumer Staples, and Financials.
-We appreciate the departure from normality reflected in leptokurtic distribution with a relevant skewness component.
-Moreover, we investigate the Gaussian assumption of the asset log-returns computing normality tests with different significance levels at 5% and 1%.
-Table 3 indicates a clear and complete deviation from normality in the stock data.
-This could be attributed to the presence of exceptional events thus a huge presence of outlier value during the analyzed period, which lead to a departure from the assumption of Normal distribution for log-returns.
-Table 3: Normality tests S&P500 log-returns data. This table illustrates the number of stocks that reject the null hypothesis of normality at 1% and 5% levels.
-Significance Level Kolmogorov-Smirnov Shapiro-Wilk Lilliefors
-1% 454 454 454
-5% 454 454 454
-To estimate the graph, we determine the tuning parameter $\lambda$ using Extended Bayesian Information Criteria (EBIC) developed in Chen and Chen, (2008):
-$\operatorname{EBIC}=-2L+Elog(n)+4\gamma Elog(p).$ (23)
-The hyperparameter $\gamma\in\left[0,1\right]$, with greater values of $\gamma$ implying a stronger penalization, $L$ denotes the log-likelihood, $n$ the sample size, $E$ the number of non-zero edges and $p$ the number of nodes.
-We set $\gamma=1$.
-The tuning parameter is selected as $\lambda_{G}=0.54$ since it minimizes the scores of the first-order differences in EBIC.
-For estimating the undirected graph conforming to the SUT distribution, using the sut Kendall estimator, we need to estimate the degrees of freedom of the log-returns distributions.
-We follow Arellano-Valle and Genton, 2010a using the maximum likelihood estimation method.
-In likelihood methods, the underlying assumption is that the observations are independent, therefore we apply the best ARIMA model to the series using the R function auto.arima to filter our returns into a white noise series, and we apply the likelihood method to the filtered series to estimate the degrees of freedom using the command mst.mle from the $R$ package sn developed by Azzalini, (2023).
-Figure 3 shows the estimated graphs using npn Kendall, csn Kendall, and sut Kendall estimators.
-Figure 3: The resulting graphs using the (a) csn Kendall estimator, (b) sut Kendall estimator, and (c) npn Kendall estimator for S&P 500 stock data from $04/01/2016$ to $31/12/2022$.
-The 454 stocks are categorized into 11 Global Industry Classification Standard (GICS) sectors, including Consumer Discretionary (51 stocks), Consumer Staples (33 stocks), Energy (21 stocks), Financials (67 stocks), Health Care (56 stocks), Industrials (69 stocks), Information Technology (59 stocks) Telecommunications Services (17 stocks), Materials (25 stocks), Real Estate (28 stocks), and Utilities (28 stocks).
-The stocks that belong to the same sector tend to cluster.
-Nevertheless, csn Kendall and sut Kendall show in better way the grouping of the stocks, thus improving the interpretability of the visualization.
-Stocks at the periphery of the network are uncorrelated to the others, during the considered period, thus they are uncorrelated with the movement of the market.
-Peralta and Zareei, (2016) shows that those stocks have a greater weight in constructing a portfolio using the Markowitz approach.
-Therefore, in high-dimension, a network approach can be valuable for identifying stocks with the least correlation to the market.
-Table 4 shows that introducing the skewness and kurtosis components results in a sparser network estimation compared to the Normal case, thus the elliptical skew-(S)KEPTIC estimator improves the interpretability of the graph.
-Table 4: Summary statistics of the estimated networks using Kendall’s tau statistic for the correlation matrix estimation.
-SK = SKEPTIC, SNK = elliptical skew-SKEPTIC, STK = elliptical skew-KEPTIC.
-DIM = dimension of the dataset.
-The notation ”$>$” indicates that one method has a greater number of edges compared to the other.
-Network  Edges No. Edges diff.
-DIM SK STK SNK SK$>$SNK SK$>$STK SNK$>$STK
-S&P 500 454 399 303 252 96 147 51
-6 Concluding remarks
-We propose a rank-based correlation matrix estimator, named the elliptical skew-(S)KEPTIC estimator which extends to the class of the skew elliptical distribution in estimating the high-dimensional graphical models.
-The elliptical skew-(S)KEPTIC estimator results for Kendall’s tau in the class of skew-elliptical distribution, and it also encompasses Spearman’s rho when the class of distribution reduces to the skew-Normal.
-We expand the SKEPTIC estimator in Liu et al., 2012a to the class of semiparametric meta skew-elliptical distribution introduced throughout the paper.
-Our analysis is not asymptotic, and the estimator is defined in a way in which we do not explicitly estimate the marginal transformation of the data.
-Theoretically, the slight reduction in the robustness of the estimator is compensated by the extension of the class of distributions over which we estimate the graphical model.
-We illustrate a method for modeling data that conforms to a $\operatorname{MS}\mathcal{E}$ distribution by treating the data generation process as a mixture of Normal and half-Normal distributions.
-For both the multivariate random variables, the SKEPTIC estimator is defined.
-Thus, the estimator is a sum of two SKEPTIC estimators and a matrix accounting for skewness.
-This extension yields a sparser pattern compared to a Gaussian graphical model, resulting in improved interpretability of the resulting graphs.
-Further, the elliptical skew-(S)KEPTIC estimator shows a slower exponential concentration in $\|\cdot\|_{\max}$ rate than the SKEPTIC estimator.
-Nevertheless, our findings show a reliable outcome that is useful when dealing with data that exhibit high skewness when it cannot be neglected.
-Our approach’s effectiveness is validated through optimal graph recoveries in numerical simulations.
-Appendix A Proofs of elliptical skew-(S)KEPTIC estimators
-A.1 Proof of Theorem 4.1
-Proof.
-Let $\hat{\rho}_{ij}$ denote the Spearman’s rho estimator of $\Sigma^{0}_{ij}=Corr(X_{i},X_{j})$, the estimator is biased (Zimmerman et al., , 2003).
-$\mathbb{E}(\widehat{\rho}_{ij})=\frac{6}{\pi(n+1)}\left[\arcsin\left(\Sigma_{ij}^{0}\right)+(n-2)\arcsin\left(\frac{\Sigma_{ij}^{0}}{2}\right)\right],$ (24)
-we rewrite (24) as:
-$\Sigma^{0}_{ij}=2\sin\left(\frac{\pi}{6}\mathbb{E}(\hat{\rho}_{ij})+a_{ij}\right)\text{where: }\quad a_{ij}=\frac{\pi\mathbb{E}(\hat{\rho}_{ij})-2\arcsin(\Sigma^{0}_{ij})}{2(n-2)}.$ (25)
-The term $|a_{ij}|\leq\frac{\pi}{n-2}$, therefore for all $n>\frac{6\pi}{t}+2$, we have $|a_{ij}|<\frac{t}{6}$ and thus:
-$\mathbb{P}\left(\left|\hat{S}^{\rho,\alpha}_{ij}-\Sigma^{0}_{ij}\right|>t\right)=$ (26)
-$=p^{2}\mathbb{P}\left(\left|4\sin\left(\frac{\pi}{6}A_{\hat{\rho}_{ij}}\right)-2\sin\left(\frac{\pi}{6}\mathbb{E}(\hat{\rho}_{ij})+a_{ij}\right)\right|>t\right)$
-$=p^{2}\mathbb{P}\left(2\left|2\sin\left(\frac{\pi}{6}A_{\hat{\rho}_{ij}}\right)-\sin\left(\frac{\pi}{6}\mathbb{E}(\hat{\rho}_{ij})+a_{ij}\right)\right|>t\right)$
-$=p^{2}\mathbb{P}\left(\left|2\sin\left(\frac{\pi}{6}A_{\hat{\rho}_{ij}}\right)-\sin\left(\frac{\pi}{6}\mathbb{E}(\hat{\rho}_{ij})+a_{ij}\right)\right|>\frac{t}{2}\right)$
-$\leq p^{2}\mathbb{P}\left(\left|2\left(A_{\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})+\frac{6}{\pi}a_{ij}\right|>\frac{3}{\pi}t\right)$
-$\leq p^{2}\mathbb{P}\left(\left|2\left(A_{\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{3}{\pi}t-\left|\frac{6}{\pi}a_{ij}\right|\right)$
-$\leq p^{2}\mathbb{P}\left(\left|2\left(A_{\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{3}{\pi}t-\frac{1}{\pi}t\right)$
-$=p^{2}\mathbb{P}\left(\left|2\left(A_{\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{2}{\pi}t\right)$
-The correlation matrix using Spearman rho, $\hat{\Sigma}_{\rho}$, can be decomposed, using modified Cholesky decomposition as $\hat{\Sigma}_{\rho}=\hat{L}\hat{D}_{\kappa}\hat{L}^{\top}$, where $\hat{D}_{\kappa}=\operatorname{I}_{p}$.
-In our case, we have: $A_{\hat{\rho}_{ij}}=\hat{L}(I-\hat{D}^{2})\hat{L}^{\top}$.
-By definition the $k$-th element of $(1-\hat{d}_{kk}^{2})<1$.
-Thus:
-$\mathbb{P}\left(\left|2\left(A_{\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{2}{\pi}t\right)\leq\mathbb{P}\left(\left|2\left({\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{2}{\pi}t\right)$ (27)
-We prove the bound of the estimator by rewriting Spearman’s rho in terms of the U-statistic and using the Hoeffding inequality. This implies:
-$\mathbb{P}\left(\sup_{ij}\left|\hat{S}^{\rho,\alpha}-\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{2}{\pi}t\right)\leq\exp\left(-\frac{nt^{2}}{16\pi^{2}}\right)$ (28)
-A.2 Proof of Theorem 4.2
-Proof.
-Kendall’s tau is an unbiased estimator: $\mathbb{E}(\hat{\tau}_{ij})=\tau_{ij}$, where $\tau_{ij}:=\operatorname{Corr}(\operatorname{sign}(X_{i}-\widetilde{X}_{i}),\operatorname{sign}(X_{j}-\widetilde{X}_{j}))$ is the population version.
-$\mathbb{P}\left(\left|\hat{\tau}_{ij}-\tau_{ij}\right|>t\right)=$
-$=\mathbb{P}\left(\left|2\sin\left(\frac{\pi}{2}A_{\hat{\tau}_{ij}}\right)-\sin\left(\frac{\pi}{2}\tau_{ij}\right)\right|>t\right)$
-$\leq\mathbb{P}\left(\left|2A_{\hat{\tau}_{ij}}-\tau_{ij}\right|>\frac{2}{\pi}t\right)$
-As in the Proof in Theorem 4.1 we have: $(I-\hat{D}^{2})<I$. This implies:
-$\mathbb{P}\left(\sup_{ij}|\hat{S}^{\tau,\alpha}-\Sigma^{0}|>t\right)\leq\exp\left(-\frac{nt^{2}}{4\pi^{2}}\right)$ (29)
-This result is obtained by exploiting the fact that we can write Kendall’s tau in terms of the U-statistic and using the Hoeffding inequality. ∎
+Similarly to the case of Spearman’s rho, we have that $\hat{S}^{\tau,\alpha}_{ij}=$ $2\sin\left(\frac{\pi}{2}A_{\hat{\tau}_{ij}}\right)$.
+We introduce $\Sigma^{\tau}$, the correlation matrix using Kendall’s tau statistic.
+We know that $A_{\tau}=L^{\top}ML$, where $M=I-D^{2}$.
+By Definition: $\Sigma^{\tau}=L^{\top}D_{\kappa}L$.
+This implies that:
+$\left\lVert\widehat{D}_{\kappa}\right\rVert_{\max}\geq\left\lVert\widehat{M}\right\rVert_{\max},$
+$\left\lVert L^{\top}\widehat{D}_{\kappa}L\right\rVert_{\max}\geq\left\lVert L^{\top}\widehat{M}L\right\rVert_{\max},$
+$\left\lVert\widehat{\Sigma}^{\tau}\right\rVert_{\max}\geq\left\lVert A_{\widehat{\tau}}\right\rVert_{\max},$
+$\left\lVert 2\sin\left(\frac{\pi}{2}\widehat{\Sigma}^{\tau}\right)\right\rVert_{\max}\geq\left\lVert 2\sin\left(\frac{\pi}{2}A_{\widehat{\tau}}\right)\right\rVert_{\max}.$
+[…]
+$E(\hat{\tau}_{ij})=\tau_{ij},$
+where
+$\tau_{ij}:=\operatorname{Corr}(\operatorname{sign}(X_{i}-\widetilde{X}_{i}),\operatorname{sign}(X_{j}-\widetilde{X}_{j}))$
+is the population version.
+$\text{pr}\left(\left\lVert 2\sin\left(\frac{\pi}{2}\hat{\Sigma}^{\tau}\right)-\Sigma^{0}\right\rVert_{\max}>t\right)\quad$ $=\text{pr}\left(\max_{ij}\ \left|2\sin\left(\frac{\pi}{2}\hat{\tau}_{ij}\right)-\sin\left(\frac{\pi}{2}\tau_{ij}\right)\right|>t\right)$
+$\leq\text{pr}\left(\max_{ij}\pi\left|\hat{\tau}_{ij}-\tau_{ij}\right|>t\right)\quad\text{(Lipshitz function)}$
+$\leq\sum_{ij}\text{pr}\left(\left|\hat{\tau}_{ij}-\tau_{ij}\right|>\frac{t}{\pi}\right)\quad\text{(Boole's inequality)}$
+$\leq p^{2}\exp\left(-\frac{nt^{2}}{2\pi^{2}}\right)\quad\text{(Hoeffding's inequality)}$
+Setting
+$\beta=p^{2}\exp\left(-\frac{nt^{2}}{2\pi^{2}}\right).$
+And solving for $t$, we get:
+$t=\sqrt{\frac{2\pi^{2}\log\left(\frac{2p^{2}}{\beta}\right)}{n}}.$
+When $\beta=\frac{1}{p}$, we have $\sqrt{\frac{2\pi^{2}}{n}\log\left({p^{3}}\right)}.$
+A.5 Proof of Theorem 5.3
+We start by proving $\|\hat{L}-L\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$ and $\|\hat{M}-M\|_{\max}O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$.
+Kendall’s tau and Spearman’s rho statistics can be rewritten in terms of U-statistics, and their estimates of correlation matrix achieve $O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$ as rate of convergence.
+Under assumption (a), we know that $\Sigma^{0}$ has bounded eigenvalues, which implies:
+$\|\hat{L}-L\|_{\max}\leq\tau_{2}\|\hat{\Sigma}-\Sigma^{0}\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right).$
+Regarding the convergence of the estimated $\hat{M}$ matrix, we know that it has diagonal elements defined as: $\hat{m}_{kk}=1-\frac{2}{\pi}\frac{\hat{\alpha}_{kk}}{1+\hat{\alpha}_{kk}}$ where $\hat{\alpha}$ parameters are estimated via the EM algorithm introduced in Abe et al., (2021), for which we assume that the regulation conditions of Wu, (1983) hold.
+Then $\|\hat{M}-M\|_{\max}O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$.
+We find bounds for $\|\hat{\Omega}-\Omega\|_{\max}$.
+$\|\hat{\Omega}-\Omega\|_{\max}$ $=\|2\sin\left(\frac{\pi}{2}(\hat{L}^{\top}\hat{M}\hat{L}^{\top})\right)^{-1}-2\sin\left(\frac{\pi}{2}({L}^{\top}{M}{L}^{\top})\right)^{-1}\|_{\max}$
+$\leq\|(\hat{L}^{\top}\hat{M}\hat{L})^{-1}-({L}^{\top}{M}{L})^{-1}\|_{\max}$
+$=\|(\hat{L}\hat{M}^{-1}\hat{L}^{\top})-({L}{M}^{-1}{L}^{\top})\|_{\max}$
+$=\|(\hat{L}-L)\hat{M}^{-1}\hat{L}^{\top}+L(\hat{M}^{-1}-M^{-1})\hat{L}^{\top}+LM^{-1}(\hat{L}-L)\|_{\max}$
+We define $A_{1}=(\hat{L}-L)\hat{M}^{-1}\hat{L}^{\top}$, $A_{2}=L(\hat{M}^{-1}-M^{-1})\hat{L}^{\top}$, and $A_{3}=LM^{-1}(\hat{L}^{\top}-L^{\top}_{0})$.
+$\|A_{1}\|_{\max}=\|(\hat{L}-L)\hat{M}^{-1}\hat{L}^{\top}\|_{\max}\leq\|\hat{L}-L\|_{\max}\|\hat{M}^{-1}\|_{\max}\|\hat{L}^{\top}\|_{1}$
+$\|\hat{L}-L\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right),\quad\|\hat{M}^{-1}\|_{\max}=O(1),\quad\|\hat{L}^{\top}\|_{1}=O(s).$
+Where $s$ are the non-zero elements in the $L$ matrix.
+$A_{2}=\|L(\hat{M}^{-1}-M^{-1})\hat{L}^{\top}\|_{\max}\leq\|L\|_{\infty}\|\hat{M}^{-1}-M^{-1}\|_{\max}\|\hat{L}^{\top}\|_{1}$
+Where $\|\hat{M}^{-1}-M^{-1}\|_{\max}=\max_{j}\left|\frac{1}{\hat{M}_{jj}}-\frac{1}{M_{0,jj}}\right|=\left|\frac{\hat{M}_{jj}-M_{0,jj}}{\hat{M}_{jj}M_{0jj}}\right|$.
+This means that: $\|\hat{M}^{-1}-M^{-1}\||_{\max}=O_{P}(\|\hat{M}-M\|_{\max})=O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$.
+$\|A_{3}\|_{\max}$ $=\|LM^{-1}(\hat{L}^{\top}-L^{\top}_{0})\|_{\max}$
+$\leq\|L\|_{\infty}\|M^{-1}\|_{\infty}\|\hat{L}^{\top}-L^{\top}_{0}\|_{\max}=O_{P}\left(s\sqrt{\frac{\log p}{n}}\right)$
+We conclude that:
+$\|\hat{\Omega}-\Omega\|_{\max}=O_{P}\left(s\sqrt{\frac{\log p}{n}}\right).$
+A.6 Proof of Theorem 5.4
+As discussed in Lam and Fan, (2009), we only need to prove a bound in $\|\hat{\Omega}-\Omega\|_{\max}$ for the three shrinkage procedures proposed to recovery $\Omega$.
+For the glasso algorithm, the proof follows the proof of Theorem 2, in Lam and Fan, (2009).
+For the CLIME, the bounds are provided in Theorem 4 in Cai et al., (2011), under the Frobenius norm, which implies $\|\cdot\|_{\max}$.
+And for the graphical Dantzig selector, the bounds are provided in Section 3 of Yuan, (2010) work.

### v3

@@
 Abstract
-We propose semiparametric estimators called elliptical skew-(S)KEPTIC for efficiently and robustly estimating non-Gaussian graphical models, relaxing the assumption of semiparametric elliptical distributions to the family of meta skew-elliptical, which accommodates a skewness component.
-Theoretically, we demonstrate that the elliptical skew-(S)KEPTIC estimators achieve robust convergence rates in both graph recovery and parameters estimation.
-We conduct numerical simulations that prove the reliable graph recovery performance of the elliptical skew-(S)KEPTIC estimators.
-Finally, the new method is applied to the daily log-returns of the stocks of the S&P500 index and shows higher sparsity compared to the Gaussian copula graphical models.
-Keywords: Elliptical distributions, robust statistics, skewness, and undirected graphical models.
+We propose semiparametric estimators, called elliptical skew-(S)KEPTIC, for efficiently and robustly estimating non-Gaussian graphical models.
+Our approach extends the semiparametric elliptical framework to the meta skew-elliptical family, which accommodates skewness.
+Theoretically, we show that the elliptical skew-(S)KEPTIC estimators achieve robust convergence rates for both graph recovery and parameter estimation.
+Through numerical simulations, we illustrate the reliable graph recovery performance of the elliptical skew-(S)KEPTIC estimators.
+Finally, we apply the new method to the daily log-returns of the stocks in the S&P 500 index and obtain a sparser graph than with Gaussian copula graphical models.
+Keywords: Elliptical distributions; Robust statistics; Skewness; Undirected graphical models.
 1 Introduction
@@
 If $X_{i}$ and $X_{j}$ are conditionally independent given $X_{\setminus{\{i,j}\}}$, then there exists no edge between the corresponding nodes $i$ and $j$ in the graph G, denoted as $(i,j)\notin E$.
-One reliable approach to estimate the undirected graphical models is the semiparametric nonparanormal (NPN) method proposed in Liu et al., (2009), which relaxes the assumption of Gaussian data, considering a set of functions $\{f_{j}\}_{j=1}^{p}$ such that: $f(\boldsymbol{X})=(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\operatorname{N}(0,\Sigma^{0})$, where $\Omega=(\Sigma^{0})^{-1}$.
+One reliable approach to estimate the undirected graphical models is the semiparametric nonparanormal (NPN) method proposed in liu2009nonparanormal, which relaxes the assumption of Gaussian data, considering a set of functions $\{f_{j}\}_{j=1}^{p}$ such that: $f(\boldsymbol{X})=(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\operatorname{N}(0,\Sigma^{0})$, where $\Omega=(\Sigma^{0})^{-1}$.
 When $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$, no edge connects $X_{i}$ and $X_{j}$ if $\Omega_{ij}=0$.
 Uncovering sparsity is subject to the estimation of the unknown correlation matrix.
-Liu et al., 2012a propose the SKEPTIC (Spearman/Kendall estimates preempt transformations to infer correlation) estimator for the nonparanormal distribution, a regularized rank-based correlation coefficient estimator based on Kendall’s tau and Spearman’s rho statistics.
+liu2012high propose the SKEPTIC (Spearman/Kendall estimates preempt transformations to infer correlation) estimator for the nonparanormal distribution, a regularized rank-based correlation coefficient estimator based on Kendall’s tau and Spearman’s rho statistics.
 The estimated correlation matrix is then plugged into shrinkage procedures that reveal the sparsity of the graph.
-The nonparanormal distribution exploits a Gaussian-based kernel and belongs to the broad class of meta-elliptical distributions described in Han and Liu, (2014) and in Liu et al., 2012b , where a set of functions $\{f_{j}\}_{j=1}^{p}$ such that $f(\boldsymbol{X})\sim\mathcal{E}C(0,\Sigma^{0},\xi)$ is considered.
-In particular, Liu et al., 2012b show that the graph $G$ is encoded in the latent partial correlation matrix $\Xi:=-\left[\text{diag}(\Omega)\right]^{-1/2}\Omega\left[\text{diag}(\Omega)\right]^{-1/2}$, thus no edge connects $X_{i}$ and $X_{j}$ if $\Xi_{ij}=0$, which implies $\Omega_{ij}=0$.
-However, as shown by Rossell and Zwiernik, (2021), the dependence structure of the elliptical distribution differs from that of the Gaussian case.
-In addition, a major pitfall of the meta-elliptical class of distributions is that, even though it embeds skewed distributions, the skewness term does not contribute to reveal the conditional independence structure.
+The nonparanormal distribution exploits a Gaussian-based kernel and belongs to the broad class of meta-elliptical distributions described in han2014scale and in liu2012transelliptical, where a set of functions $\{f_{j}\}_{j=1}^{p}$ such that $f(\boldsymbol{X})\sim\mathcal{E}C(0,\Sigma^{0},\xi)$ is considered.
+In particular, liu2012transelliptical show that the graph $G$ is encoded in the latent partial correlation matrix $\Xi:=-\left[\text{diag}(\Omega)\right]^{-1/2}\Omega\left[\text{diag}(\Omega)\right]^{-1/2}$, thus no edge connects $X_{i}$ and $X_{j}$ if $\Xi_{ij}=0$, which implies $\Omega_{ij}=0$.
+A major drawback of the meta-elliptical class is that, while it includes skewed distributions, the dependence structure of its copula does not contain the skewness parameters, preventing them from playing a role in identifying the conditional independence structure.
 We propose rank-based correlation matrix estimators, which we refer to as elliptical skew-(S)KEPTIC estimators.
-These estimators refer to the second moment of the meta skew-elliptical family of distributions, that we introduce in this paper considering a set of functions $\{f_{j}\}_{j=1}^{p}$ such that $f(\boldsymbol{X})\sim\operatorname{SU}\mathcal{E}(0,\Sigma^{0},\Lambda,h^{(p+q)},\tau,\Gamma)$, where the unified skew-elliptical ($\operatorname{SU}\mathcal{E}$) distribution is described in Arellano-Valle and Genton, 2010b .
-Extending the SKEPTIC estimator to the meta skew-elliptical family of distributions, our methodology allows for a semiparametric estimation of graphical models that take into account distributional features such as directional skewness and heavy tails.
-The main idea is to consider the unified skew-elliptical distribution as a combination of the closed skew-normal distribution introduced in Domìnguez-Molina et al., (2003), which belongs to the class of $\operatorname{SU\mathcal{E}}$ distributions.
-The class of unified skew-elliptical distributions, which generalizes the closed skew-normal, enjoys key closure properties under marginalization, conditioning, and linear transformations.
-These properties enable e.g., to define conjugate priors in probit, tobit, and multinomial probit regression models based on $\operatorname{SU\mathcal{E}}$ distributions, as both the prior and posterior remain within the same family (Durante,, 2019, Karling et al.,, 2024).
-The unified skew-elliptical distribution can be written in terms of the closed-skew normal, that possesses a stochastic representation defined through a normal distribution and a half-normal distribution, for which the SKEPTIC estimator of the correlation matrix is defined.
-The estimated correlation matrix is then plugged into existing shrinkage procedures.
-For instance, Meinshausen and Bühlmann, (2006) study a neighborhood selection approach with the lasso operator, Banerjee et al., (2008) and Friedman et al., (2008) propose an approach based on penalized likelihood, and Yuan, (2010) and Cai et al., (2011) introduce, respectively, graphical Dantzig selector and CLIME which are two methods based on linear programming that provide an improvement in terms of theoretical properties respect to the other shrinkage procedures.
-The proposed estimators generalize the SKEPTIC estimator conforming to the meta skew-elliptical distribution and embed two new subclasses of estimators based on Kendall’s tau and Spearman’s rho statistics.
-While the nonparanormal distribution is a flexible approach to model non-Gaussian data, the approach based on the meta skew-elliptical distribution allows us to quantify the asymmetry and the kurtosis presented in the data.
-The elliptical skew-(S)KEPTIC estimators possess skewness terms, encoded in the matrix of skewness parameters $\Lambda$, which assigns greater importance to the asymmetry present in data distributions.
+These estimators refer to the second moment of the meta skew-elliptical family of distributions, that we introduce in this paper considering a set of functions $\{f_{j}\}_{j=1}^{p}$ such that $f(\boldsymbol{X})\sim\operatorname{SU}\mathcal{E}(0,\Sigma^{0},\Lambda,h^{(p+q)},\tau,\Gamma)$, where the unified skew-elliptical ($\operatorname{SU}\mathcal{E}$) distribution is described in arellano2010skew-e.
+Extending the SKEPTIC estimator to the meta skew-elliptical family of distributions, our methodology allows for constructing graphical models by using a copula approach to estimate the second moment of the distribution taking into account the univariate directional skewness of the data.
+To do so, we consider the unified skew-elliptical distribution as a combination of the closed skew-normal distribution introduced in dominguez2003multivariate, which belongs to the class of $\operatorname{SU\mathcal{E}}$ distributions.
+The closed skew-normal distribution admits a stochastic representation involving both a normal and a half-normal distribution, for which the SKEPTIC estimator of the correlation matrix is defined.
+The elliptical skew-(S)KEPTIC extends this estimator by incorporating a correction term that accounts for the directional skewness of the closed skew-normal distribution.
+Remarkably, we show that, given a bivariate relationship, concordant skewness directions strengthen the dependence between the variables, whereas discordant skewness directions weaken it.
+The estimated correlation matrix is then plugged into existing shrinkage procedures, see, e.g., (meinshausen2006high; banerjee2008model; friedman2008sparse; yuan2010high; cai2011constrained).
+The elliptical skew-(S)KEPTIC estimators that generalize the SKEPTIC estimator are based on Kendall’s tau and Spearman’s rho statistics.
 When the unified skew-elliptical distribution reduces to the closed skew-normal distribution, we consider both the statistics for estimating the unknown correlation matrix and we define the elliptical skew-SKEPTIC estimator.
-Otherwise, we only consider Kendall’s tau statistic and we define the elliptical skew-KEPTIC estimator, since Spearman’s rho statistic is not invariant for the class of elliptical distributions (Hult and Lindskog,, 2002).
-The proposed framework extends to graphical models that incorporate both skewness and heavy tails, adding flexibility to nonparanormal-type data transformations.
-We prove that the elliptical skew-(S)KEPTIC estimators have deviation from the true correlation matrix $\Sigma^{0}$ that decreases as the sample size grows, showing a fast exponential rate of convergence in $\|\cdot\|_{\max}$ norm.
-The numerical analysis supports the theory.
+Otherwise, we only consider Kendall’s tau statistic and we define the elliptical skew-KEPTIC estimator, since Spearman’s rho statistic is not invariant for the class of elliptical distributions (hult2002multivariate).
+We prove that the elliptical skew-(S)KEPTIC estimators have the same rate of convergence of the SKEPTIC estimator of liu2012high plus a error rate from the estimate of the skewness parameters.
+The rates of convergence are studied in $\|\cdot\|_{\max}$ norm.
+We provide a numerical analysis to support the theory.
 Through a backward approach, we define the true sparsity pattern given by $\Omega$ and show the graph recovery using the elliptical skew-(S)KEPTIC estimators, which are compared to the SKEPTIC estimator.
-The elliptical skew-(S)KEPTIC estimators prove an optimal graph recovery, showing statistical efficiency of the proposed method.
-Prior studies by Zareifard et al., (2016), Nghiem et al., (2022), and Sheng et al., (2023) have explored graphical models while incorporating skewness into the sparsity pattern estimation.
-Zareifard et al., (2016) exploit a multivariate closed skew-normal distribution to define the skew Gaussian graphical models, which are estimated through a Bayesian approach.
-Nghiem et al., (2022) propose a novel nodewise regression approach to estimate the graph on data generated from a generalized multivariate skew-normal distribution.
-Instead, Sheng et al., (2023) include the shape parameter in the determination of the precision matrix using the skew-normal distribution of Azzalini, (1985).
-They develop an algorithm that penalizes the likelihood of this distribution to estimate $\Omega$ and, consequently, the graph.
-However, these studies do not focus on the direct estimation of the correlation matrix while incorporating skewness.
+The elliptical skew-(S)KEPTIC estimators results in similar graph recovery as the SKEPTIC, which is proven to be optimal in liu2012high, showing statistical efficiency of the proposed method.
+Prior studies by zareifard2016skew, nghiem2022estimation, and sheng2023skewed have explored graphical models while incorporating skewness into the sparsity pattern estimation.
+zareifard2016skew exploit a multivariate closed skew-normal distribution to define the skew Gaussian graphical models, which are estimated through a Bayesian approach.
+nghiem2022estimation propose a novel nodewise regression approach to estimate the graph on data generated from a generalized multivariate skew-normal distribution.
+Instead, sheng2023skewed include the shape parameter in the determination of the precision matrix using the skew-normal distribution of azzalini1985class and developing an algorithm that penalizes the likelihood of this distribution to estimate $\Omega$ and, consequently, the graph.
+However, these studies do not address the direct estimation of the correlation matrix.
+By contrast, we propose an estimator that explicitly incorporates the skewness parameters in the dependence structure through a copula-based approach.
 The paper is organized as follows.
-In Section 2 we describe the background, including a review of the SKEPTIC estimator and the semiparametric elliptical distribution.
-In Section 3 we introduce the meta skew-elliptical distribution, which forms the basis to derive the elliptical skew-(S)KEPTIC estimators, whereas in Section 4 we provide some shrinkage procedures to estimate the precision matrix $\Omega$.
+In Section 2, we describe the background, including a review of the SKEPTIC estimator and the semiparametric elliptical distribution.
+In Section 3, we introduce the meta skew-elliptical distribution, which forms the basis to derive the elliptical skew-(S)KEPTIC estimators whereas in Section 4, we provide some shrinkage procedures to estimate the precision matrix $\Omega$.
 Theoretical properties of the proposed estimator are discussed in Section 5.
-Section 6 presents simulation analysis comparing the elliptical skew-(S)KEPTIC estimators with the SKEPTIC estimator on synthetic data, and reports an empirical application to S&P 500 stock returns.
+Section 6, presents simulation analysis comparing the elliptical skew-(S)KEPTIC estimators with the SKEPTIC estimator on synthetic data, and reports an empirical application to S&P 500 stock returns.
 Section 7 concludes the paper.
 2 Background
+In this section, we briefly recall the literature preceding our article: the SKEPTIC estimator of liu2012high, the semiparametric elliptical distributions in liu2012transelliptical and han2014scale.
+The section concludes with the introduction of the unified skew-elliptical family of distributions in arellano2010skew-e, and its link with the undirected graphical models.
 2.1 Notation
@@
 A vector without its $j$-th element is defined: $\boldsymbol{a}_{\backslash j}=(\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{j-1},\boldsymbol{a}_{j+1},\ldots,\boldsymbol{a}_{p})^{\top}\in\mathbb{R}^{p-1}$, and we denote $A_{\backslash i,\backslash j}\in\mathbb{R}^{p-1\times p-1}$ the matrix without the $i$-thth row and the $j$-th column.
-We define: $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ respectively the minimum and the maximum eigenvalue of $A$.
-$\Omega$ denotes a sparse matrix, and we define the degree of $\Omega$: $\operatorname{deg}(\Omega)=\max_{1\leq i\leq p}\sum_{j=1}^{p}\operatorname{I}(|\Omega_{ij}|\neq 0)$.
+We define: $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$, respectively, the minimum and the maximum eigenvalue of $A$.
+$\Omega$ denotes a sparse matrix, and we define the degree of $\Omega$: $\operatorname{deg}(\Omega)=\max_{1\leq i\leq p}\sum_{j=1}^{p}\operatorname{I}(|\Omega_{ij}|\neq 0)$, where $i\neq j$.
 2.2 The Nonparanormal SKEPTIC
@@
 The focal point is to find a function, $f(\boldsymbol{X})$, that transforms our data into normally distributed data.
-Given $n$ observations of the $j$-th variable, $x_{1j},\dots,x_{nj}$, we define the empirical cumulative distribution function (ECDF) $\hat{F}_{j}(w)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(x_{ij}\leq w),$ where $\mathbb{I}(\cdot)$ is the indicator function.
-The transformation function $\hat{f}_{j}(x_{ij})$ is then defined as $\hat{f}_{j}(x_{ij})=\Phi^{-1}(\hat{F}_{j}(x_{ij}))$, where $\Phi^{-1}(\cdot)$ denotes the quantile normal distribution function.
-In practice, the ECDF at each observation $x_{ij}$ can be approximated using its normalize rank: $\hat{F}_{j}(x_{ij})\approx\hat{r}_{ij}=\frac{r_{ij}}{n+1}$, where $r_{ij}$ is the rank of $x_{ij}$ among $x_{1j},\dots,x_{nj}$.
-Thus, the transformation results in
-$\hat{f}_{j}(x_{ij})=\Phi^{-1}\left(\hat{r}_{ij}\right).$ (1)
+Given $n$ observations of the $j$-th variable, $x_{1j},\dots,x_{nj}$, we define the empirical cumulative distribution function (ECDF) $\widehat{F}_{j}(w)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(x_{ij}\leq w),$ where $\mathbb{I}(\cdot)$ is the indicator function.
+The transformation function $\widehat{f}_{j}(x_{ij})$ is then defined as $\widehat{f}_{j}(x_{ij})=\Phi^{-1}(\widehat{F}_{j}(x_{ij}))$, where $\Phi^{-1}(\cdot)$ denotes the quantile normal distribution function.
+In practice, the ECDF at each observation $x_{ij}$ can be approximated using its normalize rank: $\widehat{F}_{j}(x_{ij})\approx\widehat{r}_{ij}=\frac{r_{ij}}{n+1}$, where $r_{ij}$ is the rank of $x_{ij}$ among $x_{1j},\dots,x_{nj}$.
+Thus, the transformation results in:
+$\widehat{f}_{j}(x_{ij})=\Phi^{-1}\left(\widehat{r}_{ij}\right).$ (1)
 This approach ensures that the transformed data approximate a standard normal distribution while preserving the rank structure of the original data. We consider the following statistics:
-$\text{(Spearman's rho)}\ \hat{\rho}_{ij}=\frac{\sum_{k=1}^{n}(r^{k}_{j}-\bar{r}_{j})(r^{k}_{i}-\bar{r}_{i})}{\sqrt{\sum_{k=1}^{n}(r^{k}_{j}-\bar{r}_{j})^{2}\sum_{k=1}^{n}(r^{k}_{i}-\bar{r}_{i})^{2}}},$ (2)
-$\text{(Kendall's tau)}\ \hat{\tau}_{ij}=\frac{2}{n(n-1)}\sum_{1\leq k<k^{\prime}\leq n}\text{sign}((x^{k}_{j}-x^{k^{\prime}}_{j})(x^{k}_{i}-x^{k^{\prime}}_{i})).$ (3)
-Both $\hat{\rho}_{ij}$ and $\hat{\tau}_{ij}$ are non-parametric correlations between the empirical realizations of random variables $X_{j}$ and $X_{i}$.
+$\text{(Spearman's rho)}\ \widehat{\rho}_{ij}=\frac{\sum_{k=1}^{n}(r^{k}_{j}-\bar{r}_{j})(r^{k}_{i}-\bar{r}_{i})}{\sqrt{\sum_{k=1}^{n}(r^{k}_{j}-\bar{r}_{j})^{2}\sum_{k=1}^{n}(r^{k}_{i}-\bar{r}_{i})^{2}}},$ (2)
+$\text{(Kendall's tau)}\ \widehat{\tau}_{ij}=\frac{2}{n(n-1)}\sum_{1\leq k<k^{\prime}\leq n}\text{sign}((x^{k}_{j}-x^{k^{\prime}}_{j})(x^{k}_{i}-x^{k^{\prime}}_{i})).$ (3)
+Both $\widehat{\rho}_{ij}$ and $\widehat{\tau}_{ij}$ are non-parametric correlations between the empirical realizations of random variables $X_{j}$ and $X_{i}$.
 Let $\widetilde{X}_{i}$ and $\widetilde{X}_{j}$ be independent copies of the random variables $X_{i}$ and $X_{j}$, $F_{i}$ and $F_{j}$ denote the cumulative distribution functions (CDFs) of $X_{i}$ and $X_{j}$, the population versions of Spearman’s rho and Kendall’s tau are defined as: $\rho_{ij}:=\operatorname{Corr}(F_{i}(X_{i}),F_{j}(X_{i}))$ and $\tau_{ij}:=\operatorname{Corr}(\operatorname{sign}(X_{i}-\widetilde{X}_{i}),\operatorname{sign}(X_{j}-\widetilde{X}_{j}))$, respectively.
-For nonparanormal distributions, there is a connection between Spearman’s rho and Kendall’s tau to the underlying Pearson’s correlation coefficient $\Sigma^{0}_{ij}$.
-We recall the SKEPTIC estimator.
+For nonparanormal distributions, there is a connection between Spearman’s rho and Kendall’s tau to estimate the underlying Pearson’s correlation coefficient $\Sigma^{0}_{ij}$, that we recall.
 Lemma 2.1.
-(Kruskal, (1958), Liu et al., 2012a ): Assuming $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$, we have $\Sigma_{ij}^{0}=2\operatorname{sin}\left(\frac{\pi}{6}\rho_{ij}\right)=\operatorname{sin}\left(\frac{\pi}{2}\tau_{ij}\right)$.
-From these results two estimators of the unknown correlation matrix $(\Sigma^{0}_{ij})$ are $\hat{S}^{\rho}=\left[\hat{S}^{\rho}_{ij}\right]$ and $\hat{S}^{\tau}=\left[\hat{S}^{\tau}_{ij}\right]$, where: $\hat{S}^{\rho}_{ij}=2\operatorname{sin}\left(\frac{\pi}{6}\hat{\rho}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ and $\hat{S}^{\tau}_{ij}=\operatorname{sin}\left(\frac{\pi}{2}\hat{\tau}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$.
+(kruskal1958ordinal, liu2012high): Suppose that there exists a set of $f=\left\{f_{1},\ldots,f_{p}\right\}$ monotonic functions such that $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$, then we have $\Sigma_{ij}^{0}=2\operatorname{sin}\left(\frac{\pi}{6}\rho_{ij}\right)=\operatorname{sin}\left(\frac{\pi}{2}\tau_{ij}\right)$.
+From these results two estimators of the unknown correlation matrix $(\Sigma^{0}_{ij})$ are $\widehat{S}^{\rho}=\left[\widehat{S}^{\rho}_{ij}\right]$ and $\widehat{S}^{\tau}=\left[\widehat{S}^{\tau}_{ij}\right]$, where: $\widehat{S}^{\rho}_{ij}=2\operatorname{sin}\left(\frac{\pi}{6}\widehat{\rho}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ and $\widehat{S}^{\tau}_{ij}=\operatorname{sin}\left(\frac{\pi}{2}\widehat{\tau}_{ij}\right)\cdot\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$.
 Spearman’s rho and Kendall’s tau statistics are invariant under strictly increasing transformations of the underlying random variable.
 2.3 The meta-elliptical distribution
-The meta-elliptical distribution introduced in Han and Liu, (2014) relaxes the assumption of elliptically distributed data, in a similar way as the nonparanormal of Liu et al., (2009) does for Gaussian distributions.
+The meta-elliptical distribution introduced in han2014scale relaxes the assumption of elliptically distributed data, in a similar way as the nonparanormal of liu2009nonparanormal does for Gaussian distributions.
 The meta-elliptical family is a strict extension of the nonparanormal distribution.
 Definition 2.2.
-(Han and Liu, (2014), Fang, (2018)) A random vector $\boldsymbol{X}$ has an elliptical distribution, $\boldsymbol{X}\sim\mathcal{E}C_{p}(\mu,\Sigma,\xi)$, if $\boldsymbol{X}$ has the following stochastic representation: $\boldsymbol{X}\overset{d}{=}\mu+\xi AU$.
+(han2014scale, fang2018symmetric) A random vector $\boldsymbol{X}$ has an elliptical distribution, $\boldsymbol{X}\sim\mathcal{E}C_{p}(\mu,\Sigma,\xi)$, if $\boldsymbol{X}$ has the following stochastic representation: $\boldsymbol{X}\overset{d}{=}\mu+\xi AU$.
 Where: $\mu\in\mathbb{R}^{p}$, $q:=rank(A)$, AA is a $p\times q$ matrix such that $\Sigma=AA^{\top}$, $\xi\geq 0$ is a random variable independent of $U$, $U\in\mathbb{S}^{q-1}$, which is uniformly distributed of the unit sphere in $\mathbb{R}^{q}$.
+If $\xi$ has a density, then $\boldsymbol{X}$ has a density of the form:
+$f_{p}\left(x;\mu,\Sigma,h^{(p)}\right)=|\Sigma|^{-1/2}h^{(p)}((x-\mu)^{\top}\Sigma^{-1}(x-\mu)),$ (4)
+where $h^{(p)}$ is a scale function uniquely determined by the distribution of $\xi$, and it is called density generator.
 Let $\mathcal{R}_{p}:=\{\Sigma\in\mathbb{R}^{p\times p}:\Sigma^{\top}=\Sigma,\text{diag}\left(\Sigma\right)=1,\Sigma\succeq 0\}$, in analogy to the nonparanormal case, the meta-elliptical distribution is defined as follows.
 Definition 2.3.
-(Han and Liu, (2014)) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has a meta-elliptical distribution, denoted by $\boldsymbol{X}\sim M\mathcal{E}_{p}(\Sigma^{0},\xi;f)$, if there exist a set of univariate monotone functions $f_{1},\ldots,f_{p}$ such that $(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\mathcal{EC}_{p}(0,\Sigma^{0},\xi)$, where $\Sigma^{0}\in\mathcal{R}_{p}$ is the latent correlation matrix.
+(han2014scale) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has a meta-elliptical distribution, denoted by $\boldsymbol{X}\sim M\mathcal{E}_{p}(\Sigma^{0},\xi;f)$, if there exist a set of univariate monotone functions $f_{1},\ldots,f_{p}$ such that $(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\mathcal{EC}_{p}(0,\Sigma^{0},\xi)$, where $\Sigma^{0}\in\mathcal{R}_{p}$ is the latent correlation matrix.
 For the distributions that belong to the meta-elliptical family, Kendall’s tau statistic in Lemma 2.1 can be exploited to estimate the unknown correlation matrix.
@@
 Indeed, when $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$ a missing edge implies conditional independence.
-On the other hand, when $\boldsymbol{X}\sim M\mathcal{E}(\Sigma^{0},\xi,f)$ a missing edge means uncorrelation.
+On the other hand, when $\boldsymbol{X}\sim M\mathcal{E}(\Sigma^{0},\xi,f)$, if $\Omega_{ij}=0$, then $X_{i}$ and $X_{j}$ are conditionally uncorrelated (rossell2021dependence).
 […]
 […]
-We recall the family of $\operatorname{SU\mathcal{E}}$ distributions described in Arellano-Valle and Genton, 2010b .
-If a random vector $\boldsymbol{X}$ has a $\operatorname{SU\mathcal{E}}$ distribution, we write $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,p}(\mu,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$, where $h^{(2p)}$ is the density generator.
-We focus on two specifications of the family, which are: the closed skew-normal (CSN) distribution of Domìnguez-Molina et al., (2003) [^1] We build on the closed skew-normal (CSN) distribution, which represents a reparametrization of the unified skew-normal (SUN) distribution employed by Arellano-Valle and Genton, 2010b .
-and the unified skew-$t$ (SUT) distribution (Arellano-Valle and Genton, 2010a ).
+We recall the family of $\operatorname{SU\mathcal{E}}$ distributions described in arellano2010skew-e.
+If a random vector $\boldsymbol{X}$ has a $\operatorname{SU\mathcal{E}}$ distribution, we write $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,p}(\mu,\Sigma,\Lambda,h^{(p)},0,\Gamma)$.
+We focus on two specifications of the family, which are: the closed skew-normal (CSN) distribution of dominguez2003multivariate [^1] We build on the closed skew-normal (CSN) distribution, which represents a reparametrization of the unified skew-normal (SUN) distribution employed by arellano2010skew-e.
+and the unified skew-$t$ (SUT) distribution (arellano2010skew-t).
 Table 1 shows how the $\operatorname{SU\mathcal{E}}$ distribution reduces into CSN and SUT distributions.
@@
 Definition 2.4.
-(Arellano-Valle and Genton, 2010b, ) A continuous p-dimensional random vector $\boldsymbol{X}$ has a multivariate unified skew-elliptical $\operatorname{(SU\mathcal{E})}$ distribution, denoted by $\boldsymbol{X}\sim SU\mathcal{E}_{p,p}(\mu,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$, if its density function at $x\in\mathbb{R}^{p}$ is
-$f_{X}(x)=2^{p}f_{p}\left(x;\mu,\Sigma,h^{(p)}\right)F_{p}\left(\Lambda z;\Gamma,h_{Q(z)}^{(p)}\right),$ (4)
-where $\mu$ is the location parameter, $\Sigma$ is a $p\times p$ dispersion matrix, $h^{(p)}$ is the density generator, $F_{r}\left(x;\Sigma,h^{(r)}\right)$ the $r$-dimensional centered elliptical cumulative distribution function with $r\times r$ dispersion matrix $\Sigma$ and density generator $h^{(r)}$.
-Additionally, $z=\sigma^{-1}(x-\mu)$, $\Lambda$ is a $p\times p$ real matrix controlling shape, $Q(z)=z^{\top}(\Sigma^{0})^{-1}z$, $\Gamma$ is a $p\times p$ positive definite symmetric matrix, $f_{p}\left(x;\mu,\Sigma,h^{(p)}\right)=$ $|\Sigma|^{-1/2}h^{(p)}(Q(z))$ denotes the density function of an elliptically contoured distribution with location $\mu\in\mathbb{R}^{p}$.
-In conclusion, $h_{Q(z)}^{(p)}(u)=h^{(2p)}(u+Q(z))/h^{(p)}(Q(z))$.
+(arellano2010skew-e) A continuous p-dimensional random vector $\boldsymbol{X}$ has a multivariate unified skew-elliptical $\operatorname{(SU\mathcal{E})}$ distribution, denoted by $\boldsymbol{X}\sim SU\mathcal{E}_{p,p}(\mu,\Sigma,\Lambda,h^{(p)},0,\Gamma)$, if its density function at $x\in\mathbb{R}^{p}$ is:
+$f_{X}(x)=2^{p}f_{p}\left(x;\mu,\Sigma,h^{(p)}\right)F_{p}\left(\Lambda\sigma^{-1}(x-\mu);\Gamma,h_{Q(z)}^{(p)}\right),$ (5)
+where $f_{p}\left(x;\mu,\Sigma,h^{(p)}\right)$ is the elliptical distribution defined in (4), and $F_{p}\left(\Lambda\sigma^{-1}(x-\mu);\Gamma,h^{(p)}_{Q(x)}\right)$ the $p$-dimensional centered elliptical cumulative distribution function with $p\times p$ dispersion matrix $\Gamma$, $\Lambda$ is a $p\times p$ real matrix controlling shape.
+In conclusion, $Q(\mathbf{x})$ is a quadratic form defined as $Q(\mathbf{x})=(\mathbf{x}-\mu)^{\top}\boldsymbol{\Sigma}^{-1}(\mathbf{x}-\mu)\in\mathbb{R}^{+}$, whereas $h_{Q(\mathbf{x})}^{(p)}(u)=h^{(2p)}[u+Q(\mathbf{x})]/g^{(p)}[Q(\mathbf{x})]$ denotes the elliptical conditional density generator.
 $\operatorname{SU\mathcal{E}}$ $h^{(p)}$
@@
 Proposition 2.5.
-(Arellano-Valle and Genton, 2010b ) Given $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,p}(\mu,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$, we can write $\boldsymbol{X}$ as a mixture of a CSN distribution: $\boldsymbol{X}=\mu+\sigma V_{0}^{-1/2}\boldsymbol{Z}_{0}$, where $V_{0}$ is the cumulative distribution function of a uni-dimensional random variable, called mixture variable, such that: $V_{0}(0)=0$, and $V_{0}$ is independent of $\boldsymbol{Z}_{0}\sim\operatorname{CSN}_{p,p}(0,\Sigma^{0},\Lambda,0,\Gamma)$.
+(arellano2010skew-e) Given $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,p}(\mu,\Sigma,\Lambda,h^{(p)},0,\Gamma)$, we can write $\boldsymbol{X}$ as a mixture of a CSN distribution: $\boldsymbol{X}=\mu+\sigma V_{0}^{-1/2}\boldsymbol{Z}_{0}$, where $V_{0}$ is the cumulative distribution function of a uni-dimensional random variable, called mixture variable, such that: $V_{0}(0)=0$, and $V_{0}$ is independent of $\boldsymbol{Z}_{0}\sim\operatorname{CSN}_{p,p}(0,\Sigma^{0},\Lambda,0,\Gamma)$.
 Lemma 2.6.
-(Arellano-Valle and Genton, 2010b, ) If  $\text{ $V_{0}$}\sim Gamma(\nu/2,\nu/2)$, then
-$\boldsymbol{X}\sim\operatorname{SU}\mathcal{E}_{p,p}(\mu,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$ reduces to $\operatorname{SUT}$ distribution: $\boldsymbol{X}\sim\operatorname{SUT}_{p,p}(\mu,\Sigma,\Lambda,\nu,0,\Gamma)$.
-We construct an undirected graph on the closed skew-normal distribution, following Zareifard et al., (2016).
+(arellano2010skew-e) If  $\text{ $V_{0}$}\sim Gamma(\nu/2,\nu/2)$, then the unified skew-elliptical distribution $\boldsymbol{X}\sim\operatorname{SU}\mathcal{E}_{p,p}(\mu,\Sigma,\Lambda,h^{(p)},0,\Gamma)$ reduces to a unified skew-$t$ $\operatorname{(SUT)}$ distribution: $\boldsymbol{X}\sim\operatorname{SUT}_{p,p}(\mu,\Sigma,\Lambda,\nu,0,\Gamma)$.
+We construct an undirected graph on the closed skew-normal distribution, following zareifard2016skew.
 Definition 2.7.
-Consider an undirected decomposable graph $G=(V,E)$, a random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ is called a skew decomposable graphical model concerning the graph G with mean $\mu$, the precision matrix $\Omega$ and the skewness parameter $\alpha\in\mathbb{R}^{p}$ if its density is $\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ or equivalently
+Consider an undirected graph $G=(V,E)$, a random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ is called a skew graphical model concerning the graph G with mean $\mu$, the precision matrix $\Omega$ and the skewness parameter $\alpha\in\mathbb{R}^{p}$ if its density is $\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ or equivalently
 $f_{\boldsymbol{X}}(x)=2^{p}\phi_{p}\left(x-\mu;0,\Omega^{-1}\right)\Phi_{p}\left(D_{\alpha}L(x-\mu);0,D^{-1}_{\kappa}\right),$ (6)
-where $L$ and $D_{\kappa}$ are the matrix resulting from the modified Cholesky of $\Omega=LD_{\kappa}L^{\top}$, and zero in $L$ matrix determines the conditional independence relations between the $X_{1},\ldots,X_{p}$ variables, in other words: $L_{ij}=0\Leftrightarrow\Omega_{ij}=0$.
+where $L$ and $D_{\kappa}$ are the matrix resulting from the modified Cholesky of $\Omega=LD_{\kappa}L^{\top}$, and zero in $L$ matrix determines the conditional independence relations between the $X_{1},\ldots,X_{p}$ variables.
 Further, when the matrix $D_{\alpha}=0$, the random vector $\boldsymbol{X}\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ reduces to $\boldsymbol{X}\sim\operatorname{N}_{p}(0,\Omega^{-1})$.
 The graphical model for the unified skew-elliptical distribution can be constructed by utilizing the sparsity structure of the closed skew-normal distribution.
-We rely on a semiparametric approach that considers the class of unified skew-elliptical distributions to extend the SKEPTIC estimator of Liu et al., 2012a proposing a robust estimator of graphical models that accounts for skewness and heavy tails.
+We employ a semiparametric approach, utilizing the class of unified skew-elliptical distributions, to extend the SKEPTIC estimator proposed by liu2012high.
+This approach enables the development of robust estimators for graphical models that quantify the skewness present in the data distributions.
 3 The elliptical skew-(S)KEPTIC estimators
 Definition 3.1.
-(Meta skew-elliptical distribution) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has meta skew-elliptical distribution, denoted by $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},\Lambda,h^{(2p)},0,\Gamma;f)$, if there exists a set of monotone univariate functions $f_{1},\ldots,f_{p}$ such that
-$(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Sigma^{0},\Lambda,h^{(2p)},0,\Gamma),$ (6)
+(Meta skew-elliptical distribution) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has meta skew-elliptical distribution, denoted by $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},\Lambda,h^{(p)},0,\Gamma;f)$, if there exists a set of monotone univariate functions $f_{1},\ldots,f_{p}$ such that
+$(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Sigma^{0},\Lambda,h^{(p)},0,\Gamma),$ (7)
 where $\Lambda$ is a matrix containing the skewness terms.
@@
 $\boldsymbol{X}\stackrel{{\scriptstyle d}}{{=}}L^{-1}D_{\kappa}^{-\frac{1}{2}}D_{\alpha}\left(I+D_{\alpha}^{2}\right)^{-\frac{1}{2}}\boldsymbol{U}+L^{-1}D_{\kappa}^{-\frac{1}{2}}\left(I+D_{\alpha}^{2}\right)^{-\frac{1}{2}}\boldsymbol{V},$ (8)
-where: $\boldsymbol{U}\sim HN_{p}(0,I_{p})$ and $\boldsymbol{V}\sim N_{p}(0,I_{p})$, where $HN_{p}(0,I_{p})$ is a $p$-dimensional half-normal distribution, and $N_{p}(0,I_{p})$ is a $p$-dimensional normal distribution.
-The distributions of $\boldsymbol{U}$ and $\boldsymbol{V}$ are independent.
-The first two moments of $\boldsymbol{X}$ are:
-${E}(\boldsymbol{X})$ $=L^{\top}D_{\kappa}^{-\frac{1}{2}}D,$
-$\text{cov}(\boldsymbol{X})$ $=\left(L^{\top}D_{\kappa}^{\frac{1}{2}}(I-D^{2})D_{\kappa}^{\frac{1}{2}}L\right)^{-1},$ (8)
-where $D=\sqrt{\frac{2}{\pi}}D_{\alpha}(I_{p}+D^{2}_{\alpha})^{-\frac{1}{2}}$.
+where: $\boldsymbol{U}\sim HN_{p}(0,I_{p})$ and $\boldsymbol{V}\sim N_{p}(0,I_{p})$, where $HN_{p}(0,I_{p})$ is a $p$-dimensional half-normal distribution, and $N_{p}(0,I_{p})$ is a $p$-dimensional normal distribution, the distributions of $\boldsymbol{U}$ and $\boldsymbol{V}$ are independent.
+We introduce a correlation matrix estimator that explicitly accounts for the cross effects of univariate skewness parameters on the linear dependence among variables.
+The robustness of the proposed estimators follows from the direct estimation of the skewness parameters without the explicit computation of the marginal distributions, in the spirit of the SKEPTIC estimator of liu2012high.
 Theorem 3.4 (elliptical skew-KEPTIC estimator).
-Assume that $\boldsymbol{X}$ follows a meta skew-elliptical distribution: $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa};f)$, such that $\boldsymbol{Y}=f(\boldsymbol{X})$ denotes the unified skew-elliptical distribution $\boldsymbol{Y}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa})$.
+Assume that $\boldsymbol{X}$ follows a meta skew-elliptical distribution: $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(p)},0,D_{\kappa};f)$, such that $\boldsymbol{Y}=f(\boldsymbol{X})$ denotes the unified skew-elliptical distribution $\boldsymbol{Y}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Omega^{-1},D_{\alpha}L,h^{(p)},0,D_{\kappa})$.
 The elliptical skew-KEPTIC estimator is:
-${S}^{\tau,\alpha}_{ij}=\begin{cases}2\operatorname{sin}\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)&i\neq j\\ 1&i=j\end{cases}.$
-As discussed in Hult and Lindskog, (2002), Spearman’s rho is not invariant in the class of elliptical distributions, and its relation with the Pearson correlation coefficient in Lemma 2.1 holds if and only if the distribution is nonparanormal.
-The proof of the Theorem is in Appendix A.
+${S}^{\tau,\alpha}_{ij}=\begin{cases}\sin\Big(\frac{\pi}{2}\tau_{ij}\Big)\frac{1+\alpha_{i}\alpha_{j}}{\sqrt{(1+\alpha_{i}^{2})(1+\alpha_{j}^{2})}}&i\neq j\\ 1&i=j\end{cases}.$
+As discussed in hult2002multivariate, Spearman’s rho is not invariant in the class of elliptical distributions, and its relation with the Pearson correlation coefficient in Lemma 2.1 holds if and only if the distribution is nonparanormal.
+The proof of the Theorem is in A.
 Lemma 3.5.
-(elliptical skew-SKEPTIC estimator) Assume that $\boldsymbol{X}$ follows a meta skew-elliptical distribution: $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa};f)$, such that $\boldsymbol{Y}=f(\boldsymbol{X})$ denotes the unified skew-elliptical distribution.
-When $\boldsymbol{Y}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa})$ reduces to the CSN distribution in (5), the elliptical skew-SKEPTIC estimator is:
-${S}^{\rho,\alpha}_{ij}=\begin{cases}4\sin\left(\frac{\pi}{6}A_{{\rho}_{ij}}\right)&\text{if }i\neq j\\ 1&\text{if }i=j\end{cases},\quad\quad{S}^{\tau,\alpha}_{ij}=\begin{cases}2\operatorname{sin}\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)&i\neq j\\ 1&i=j\end{cases}.$
+(elliptical skew-SKEPTIC estimator) Assume that $\boldsymbol{X}$ follows a meta skew-elliptical distribution: $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(p)},0,D_{\kappa};f)$, such that $\boldsymbol{Y}=f(\boldsymbol{X})$ denotes the unified skew-elliptical distribution.
+When $\boldsymbol{Y}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Omega^{-1},D_{\alpha}L,h^{(p)},0,D_{\kappa})$ reduces to the CSN distribution in (6), the elliptical skew-SKEPTIC estimator is:
+${S}^{\rho,\alpha}_{ij}=\begin{cases}2\sin\Big(\frac{\pi}{6}\rho_{ij}\Big)\frac{1+\alpha_{i}\alpha_{j}}{\sqrt{(1+\alpha_{i}^{2})(1+\alpha_{j}^{2})}}&i\neq j\\ 1&i=j\end{cases},\quad{S}^{\tau,\alpha}_{ij}=\begin{cases}\sin\Big(\frac{\pi}{2}\tau_{ij}\Big)\frac{1+\alpha_{i}\alpha_{j}}{\sqrt{(1+\alpha_{i}^{2})(1+\alpha_{j}^{2})}}&i\neq j\\ 1&i=j\end{cases}.$
 We generally refer to the estimators in Theorem 3.4 and Lemma 3.5 as elliptical skew-(S)KEPTIC estimators.
-To obtain the correlation matrix using the elliptical skew-(S)KEPTIC estimators, we need to estimate the parameters of the CSN distribution.
-The estimation procedure is based on the EM algorithm of Abe et al., (2021) [^2] The estimated parameters come from the skew-normal distribution of Azzalini and Capitanio, (1999), with an adequate parametrization explained in Arellano-Valle and Azzalini, (2006) we conform the parameters to a CSN distribution..
+By construction, the estimated dependence measures are amplified when the skewness components exhibit concordance, whereas they are attenuated when the skewness components are discordant.
+To obtain the correlation matrix using the elliptical skew-(S)KEPTIC estimators, we need to estimate the parameters of the distributions, which are encoded in the diagonal elements of $D_{\alpha}$, for closed skew-normal and unified skew-$t$ distributions.
 4 Methods for the Estimation of the Precision Matrix and the Undirected Graph
 4.1 The Dantzig selector
-The Dantzig selector estimator (Yuan,, 2010) exploits the relation between the coefficients $\theta$ stemming from a multivariate linear regression and the elements of the precision matrix.
+The Dantzig selector estimator (yuan2010high) exploits the relation between the coefficients $\theta$ stemming from a multivariate linear regression and the elements of the precision matrix.
 A sparse precision matrix implies sparse regression coefficients estimation that relies on an algorithm that depends on the tuning parameter $\delta$.
@@
   • Symmetrization:
-    $\hat{\Omega}^{gDS}=\text{arg}\min_{\Omega=\Omega^{T}}\|\Omega-\hat{\Omega}\|_{1}.$ (10)
+    $\widehat{\Omega}^{gDS}=\text{arg}\min_{\Omega=\Omega^{T}}\|\Omega-\widehat{\Omega}\|_{1}.$ (10)
 4.2 The Glasso Algorithm
-Glasso algorithm is developed by Friedman et al., (2008) and it consists in adding a $L_{1}$-norm penalty term in the maximum likelihood estimation of $\Omega$ as follows:
+Glasso algorithm is developed by friedman2008sparse and it consists in adding a $L_{1}$-norm penalty term in the maximum likelihood estimation of $\Omega$ as follows:
 $\Omega^{glasso}=arg\max_{\Omega\succ 0}\left\{\text{log}\ \text{det}\Omega-\text{tr}(S\Omega)-\lambda_{G}\|\Omega\|_{1}\right\}.$ (11)
 The penalty term, $\lambda_{G}$, is the parameter that controls the sparsity of the estimated matrix, i.e., the number of zeros in the precision matrix.
-Banerjee et al., (2008) shows that the shrinking procedure based on the tuning parameter $\lambda_{G}$ is directly proportional to the input covariance matrix.
+banerjee2008model shows that the shrinking procedure based on the tuning parameter $\lambda_{G}$ is directly proportional to the input covariance matrix.
 This means that greater values of $\lambda_{G}$ imply a greater presence of zeros in the estimated $\Omega$, therefore, a minor number of interconnections are unveiled.
@@
 CLIME stands for constrained $\ell_{1}$-minimization for inverse matrix estimation.
-This algorithm, proposed by Cai et al., (2011), is very attractive for high-dimensional data since the estimator is obtained by combining vector solutions by solving a linear program.
-Let $\Omega$ denote a $p\times p$ precision matrix and $\{\hat{\Omega}_{1}\}$ be the solution set of the optimization problem:
+This algorithm, proposed by cai2011constrained, is very attractive for high-dimensional data since the estimator is obtained by combining vector solutions by solving a linear program.
+Let $\Omega$ denote a $p\times p$ precision matrix and $\{\widehat{\Omega}_{1}\}$ be the solution set of the optimization problem:
 $\text{min}\|\Omega\|_{1}\quad\text{subject to:}$ (12)
 $\|\Sigma\Omega-I\|_{\infty}\leq\lambda_{C},$
-where $\lambda_{C}$ is the tuning parameter. The final CLIME estimator of $\Omega$ is defined by the symmetrization of $\hat{\Omega}_{1}$:
-$\hat{\Omega}=(\hat{\omega}_{ij}),\quad\text{where:}$ (13)
-$\hat{\omega}_{ij}=\hat{\omega}_{ji}=\hat{\omega}^{1}_{ij}I\{|\hat{\omega}^{1}_{ij}\leq|\hat{\omega}^{1}_{ij}|\}+\hat{\omega}^{1}_{ji}I\{|\hat{\omega}^{1}_{ji}\leq|\hat{\omega}^{1}_{ji}|\}.$
+where $\lambda_{C}$ is the tuning parameter. The final CLIME estimator of $\Omega$ is defined by the symmetrization of $\widehat{\Omega}_{1}$:
+$\widehat{\Omega}=(\widehat{\omega}_{ij}),\quad\text{where:}$ (13)
+$\widehat{\omega}_{ij}=\widehat{\omega}_{ji}=\widehat{\omega}^{1}_{ij}I\{|\widehat{\omega}^{1}_{ij}\leq|\widehat{\omega}^{1}_{ij}|\}+\widehat{\omega}^{1}_{ji}I\{|\widehat{\omega}^{1}_{ji}\leq|\widehat{\omega}^{1}_{ji}|\}.$
 The CLIME estimator has the property that the optimization problem defined in (12) can be decomposed into $p$ optimization problems defined as:
@@
 where $\boldsymbol{e}_{i}$ is defined as a standard unit vector in $\mathbb{R}^{p}$ with 1 in the $i$-th coordinate and 0 in the all other coordinates and $\boldsymbol{\beta}$ is a vector in $\mathbb{R}^{p}$.
-Solving (14) $p$ times, we obtain the CLIME estimator: $\{\hat{\Omega}_{1}\}=\{\hat{B}\}:=\{(\hat{\boldsymbol{\beta}}_{1},...,\hat{\boldsymbol{\beta}}_{p})\}$.
+Solving (14) $p$ times, we obtain the CLIME estimator: $\{\widehat{\Omega}_{1}\}=\{\widehat{B}\}:=\{(\widehat{\boldsymbol{\beta}}_{1},...,\widehat{\boldsymbol{\beta}}_{p})\}$.
 5 Theoretical properties
-We investigate the theoretical properties of the elliptical skew-(S)KEPTIC estimators, establishing concentration inequalities in $\|\cdot\|_{\max}$ for our correlation matrix estimators, and the sparsistency of $\hat{\Omega}$ which describes the characteristic where any parameter that is originally zero is correctly estimated as zero, with a probability approaching one.
-For the rate of convergence of the correlation matrix estimator, we revisit the results of Liu et al., 2012a .
-For the corresponding precision matrix, we follow the framework introduced in Lam and Fan, (2009) for demonstrating the rate of convergence and the sparsistency of $\hat{\Omega}$.
-Our findings are based on the application of shrinkage procedures incorporating an $L_{1}$-penalty, with the three methods outlined in Section 4 adhering to this approach.
+We analyze the theoretical properties of the elliptical skew-(S)KEPTIC estimator.
+The results suggest that our estimator is reliable when dealing with highly skewed data.
+Our results show an exponential concentration rate in $\|\cdot\|_{\max}$.
+Moreover, we prove optimal bounds for shrinkage procedures.
+We show the results for the Dantzig selector, however, they can be easily extended to the CLIME and the Glasso procedures.
 Theorem 5.1.
-Let $x_{1},...,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha},h^{(2p)},0,D_{\kappa};f)$, restricted to the case $h^{(p)}=\phi^{(p)}(u)$, and let $\hat{S}^{\rho,\alpha}_{ij}=4\sin(\frac{\pi}{6}A_{\hat{\rho,}_{ij}})\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ denote the elliptical skew-SKEPTIC, then with probability at least $1-\frac{1}{p}$ we have:
-$\|\hat{S}^{\rho,\alpha}_{ij}-\Sigma^{0}_{ij}\|_{\max}\leq\sqrt{\frac{3(2\pi-3)^{2}}{(9-2\pi)^{2}}\frac{\log\left({p}\right)}{n}}.$ (15)
+Let $x_{1},\dots,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},\Lambda,h^{(p)},0,\Gamma;f)$ restricted to the case $h^{(p)}=\phi^{(p)}(u)$, and let
+$\widehat{S}^{\rho,\alpha}_{ij}=2\sin\!\left(\tfrac{\pi}{6}\widehat{\rho}_{ij}\right)\,\frac{1+\widehat{\alpha}_{i}\widehat{\alpha}_{j}}{\sqrt{(1+\widehat{\alpha}_{i}^{2})(1+\widehat{\alpha}_{j}^{2})}}\,\operatorname{I}(i\neq j)\;+\;\operatorname{I}(i=j)$
+denote the elliptical skew-SKEPTIC estimator. Then, with probability at least $1-2/p^{2}$, we have
+$\|\widehat{S}^{\rho,\alpha}-\Sigma^{0}\|_{\max}\leq(8\pi+2C_{\alpha})\sqrt{\frac{\log p}{n}}.$ (15)
 Theorem 5.2.
-Let $x_{1},...,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha},h^{(2p)},0,D_{\kappa};f)$ and let $\hat{S}^{\tau,\alpha}_{ij}=2\sin(\frac{\pi}{2}A_{\hat{\tau}_{ij}})\operatorname{I}(i\neq j)+\operatorname{I}(i=j)$ denote the elliptical skew-(S)KEPTIC estimators, then with probability at least $1-\frac{1}{p}$ we have:
-$\|\hat{S}^{\tau,\alpha}_{ij}-\Sigma^{0}_{ij}\|_{\max}\leq\sqrt{{6\pi^{2}}\frac{\log\left({p}\right)}{n}}.$ (16)
-The proofs of Theorems 5.1 and 5.2 are in Appendix A.
-The elliptical skew-(S)KEPTIC estimators are consistent, so when the sample size $n$ increases, the estimator converges to the true correlation matrix.
-Let $\mathcal{S}=\{(i,j):\omega^{0}_{ij}\neq 0\}$ be the support set of off-diagonal nonzero entries, with cardinality $|\mathcal{S}|=s+p$, let $p_{\lambda}(\cdot)$ be a penalty function, and define
-$a=\max_{(i,j)\in\mathcal{S}}p^{\prime}_{\lambda}(|\ell_{ij}|),\quad b=\max_{(i,j)\in\mathcal{S}}p^{\prime\prime}_{\lambda}(|\ell_{ij}|).$
-where $\ell_{ij}$ is the $(i,j)$ element of the $L$ matrix that stems from $\Omega=L^{\top}ML$.
-In this work, we consider shrinkage procedures based on $L_{1}$-penalty, which means $a=\lambda$ and $b=0$.
-We impose the following regularity conditions:
-  • (a)
-    The eigenvalues of $\Sigma_{0}$ are bounded: $0<\tau_{1}<\lambda_{\min}(\Sigma_{0})\leq\lambda_{\max}(\Sigma_{0})<\tau_{2}<\infty$.
-  • (b)
-    Assume $a=O\left(\sqrt{\frac{(1+p/(s+1))\log p}{n}}\right)$, $b=o(1)$, and $\min_{(i,j)\in\mathcal{S}}|\omega^{0}_{ij}|/\lambda\to\infty$.
-  • (c)
-    The penalty is singular at the origin: $\lim_{t\downarrow 0}p_{\lambda}(t)/(\lambda t)=k>0$.
-  • (d)
-    For some constants $C,D>0$, when $\theta_{1},\theta_{2}<C\lambda$,
-    $|p^{\prime\prime}_{\lambda}(\theta_{1})-p^{\prime\prime}_{\lambda}(\theta_{2})|\leq D|\theta_{1}-\theta_{2}|.$
+Let $x_{1},...,x_{n}$ be $n$ observations from $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},\Lambda,h^{(p)},0,\Gamma;f)$ and let
+$\widehat{S}^{\tau,\alpha}_{ij}=\sin\!\left(\tfrac{\pi}{2}\widehat{\tau}_{ij}\right)\,\frac{1+\widehat{\alpha}_{i}\widehat{\alpha}_{j}}{\sqrt{(1+\widehat{\alpha}_{i}^{2})(1+\widehat{\alpha}_{j}^{2})}}\,\operatorname{I}(i\neq j)\;+\;\operatorname{I}(i=j),$
+denote the elliptical skew-(S)KEPTIC estimator, then with probability at least $1-2/p^{2}$ we have:
+$\|\widehat{S}^{\tau,\alpha}_{ij}-\Sigma^{0}_{ij}\|_{\max}\leq(2.45\pi+2C_{\alpha})\sqrt{\frac{\log p}{n}}.$ (16)
+Proofs of Theorems 5.1 and 5.2 are in A. The elliptical skew-(S)KEPTIC estimator achieves convergence, thus it is possible to define the following result.
 Theorem 5.3.
-Suppose that conditions (a), (b), (c), (d) hold.
-If $\frac{\log p}{n}=O(\lambda^{2})$, then there exists a local minimizer for the maximum likelihood estimator for $\widehat{{M}}$ and an estimate of $\widehat{{L}}$ that satisfy:
-$\|\widehat{{M}}-{M}\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right),\quad\|\widehat{{L}}-{L}\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right),$
-which implies:
-$\|{\Omega}-\Omega\|_{\max}=O_{P}\left(s\sqrt{\frac{\log p}{n}}\right).$
-The proof of Theorem 5.3 is in Appendix A.
-We know from Lemma 1 in Zareifard et al., (2016) that $L_{ij}=0\Leftrightarrow\Omega_{ij}=0$, so the matrices $L$ and $\Omega$ share the same sparsity patter.
-For this reason, we provide sparsistency directly on $\Omega$.
+When we substitute the estimated matrices $\hat{S}^{\rho,\alpha}$ and $\hat{S}^{\tau,\alpha}$ into the parametric Graphical Lasso (or the CLIME or graphical Dantzig selector), the estimators provide convergence, ensuring consistency and graph recovery.
+Proof.
+The proof is based on the observation that the sample correlation matrix $\widehat{S}$ is a sufficient statistic for all three shrinkage procedures: the graphical lasso, graphical Dantzig selector, and CLIME.
+As yuan2010high points out, a sufficient condition to ensure that an estimator is consistent is if there exists some constant $c$, such that $\mathbb{P}\left(\|\widehat{S}-\Sigma^{0}\|_{\max}>c\sqrt{\frac{\log p}{n}}\right)\leq 1-\frac{1}{p}$, which can be replaced by (15) and (16) in Theorems 5.1 and 5.2.
+∎
+The convergence of the elliptical skew-(S)KEPTIC estimator also means that the precision matrix is bounded.
+Similar results are found by cai2011constrained and yuan2010high.
+In particular, let $\widehat{\Omega}^{mse-s}$ denote the inverse correlation matrix using the elliptical skew-(S)KEPTIC estimator and consider the Dantzig selector as the shrinkage procedure to find $\widehat{\Omega}$.
+We introduce a class of inverse matrix defined as $\mathcal{M}_{1}(\kappa,\tau,M):=\{\Omega:\Omega\succ 0,\text{diag}(\Omega^{-1})=1,\|\Omega\|_{1}\leq\kappa,\frac{1}{\tau}\leq\lambda_{\min}(\Omega)\leq\lambda_{\max}(\Omega)\leq\tau,\deg(\Omega)\leq M\}$, where $\kappa,\tau>1$.
+Given Theorem 5.3, we obtain the following result:
 Theorem 5.4.
-(Sparsistency) Under the conditions (a), (b), (c), (d) for a local minimizer based on $L_{1}$-penalty that satisfies $\|\widehat{\Omega}-\Omega\|_{\max}^{2}=O_{P}\left\{s\log p/n\right\}$ and $\|\widehat{\Omega}-\Omega\|^{2}=O_{P}\left(\eta_{n}\right)$ for a sequence of $\eta_{n}\rightarrow 0$, where $\eta_{n}\equiv s\sqrt{\frac{\log p}{n}}$.
-If $\log p/n+\eta_{n}=O\left(\lambda_{n1}^{2}\right)$, then with probability tending to $1,\widehat{\omega}_{ij}=0$ for all $(i,j)\in S^{c}$.
-The proof of Theorem 5.4 is in Appendix A.
+For $1\leq q<\infty$, there exists a constant $c$ that depends on $\kappa$, $\tau$, $\lambda_{\min}(\Omega)$ and $\lambda_{\max}(\Omega)$, such that
+$\sup_{\Omega\in\mathcal{M}_{1}(\kappa,\tau,M)}\|\widehat{\Omega}^{mse-s}-\Omega\|_{q}=O_{P}\left(M\sqrt{\frac{\log p}{n}}\right).$ (17)
+provided that $\lim_{n\rightarrow\infty}\frac{n}{M^{2}\log d}=\infty$ and $\delta=c\sqrt{\frac{\log d}{n}}$, for sufficiently large $c$, where $\delta$ is the parameter in (9).
+Proof.
+This result follows yuan2010high, which proves that the difference between the estimated precision matrix and the theoretical precision matrix is bounded by $\|\widehat{S}-\Sigma^{0}\|_{\max}$, where $\widehat{S}$ is the estimated correlation matrix and $\Sigma^{0}$ is the theoretical correlation matrix.
+In our case, to prove the bounds for the estimated correlation matrix, we just replace $\widehat{S}$ with $\hat{S}^{\rho,\alpha}$ and $\hat{S}^{\tau,\alpha}$ where the bounds are proved in Theorem 5.1 and 5.2.
+∎
 6 Experiments
+In this section, we conduct an empirical analysis based on both synthetic and real data.
+The study of synthetic data is performed through numerical simulations, whereas the analysis of real data involves the estimation of undirected graphs for the stocks that compose the S&P500500 index.
 6.1 Numerical Simulations
-We perform numerical simulations by generating synthetic data from the normal distribution, the closed skew-normal distribution, and the unified skew-$t$ distribution.
-For each dataset, we compute both the SKEPTIC and the skew-(S)KEPTIC estimators, and we evaluate the graph recovery performance of the proposed methods.
-The implementation follows these steps.
-Using the R package huge (Jiang et al., (2021)), we generate the precision matrix $\Omega=(\Sigma^{0})^{-1}$ using the ”random” structure, in which we randomly generate pairs of conditional dependencies in the off-diagonal elements.
-We sample $n$ observations from $p$-variate normal, CSN and SUT distribution.
+The implementation of the numerical simulations is carried out following these steps.
+Using the R package huge (jiang2020package), we generate the precision matrix $\Omega=(\Sigma^{0})^{-1}$ using the ”random” structure, in which we randomly generate pairs of conditional dependencies in the off-diagonal elements, the sparsity of $\Omega$ is set to 0.02 to ensure the positive defitivness of the matrix so that the matrix can be inverted.
+We consider the case of $p=100$, and $n=200$, and we sample $n$ observations from $p$-variate normal distribution.
 We set the location parameter equal to zero, and $\Sigma^{0}=\Omega^{-1}$.
-We consider the following cases: $p\in\{100,250,300\}$, and $n\in\{200,1000,500\}$.
-We generate synthetic data assuming $f(\boldsymbol{X})\sim N(0,\Sigma^{0})$, $f(\boldsymbol{X})\sim\operatorname{CSN}(0,\Sigma^{0},D_{\alpha}L,0,D_{\kappa})$, and $f(\boldsymbol{X})\sim\operatorname{SUT}(0,\Sigma^{0},D_{\alpha}L,0,D_{\kappa})\equiv\frac{W}{\sqrt{V_{0}}}$, where $W$ is distributed according to a closed skew-normal distribution, and $V_{0}\sim Gamma(\nu/2,\nu/2)$ as in Proposition 2.5.
-We set $\nu=3$ to emphasize the heavy tails in the distribution.
-To generate the data from the closed skew-normal distribution, we rely on the stochastic representation in (7).
-Moreover, we generate the skewness parameters from $\alpha=0.4\cdot\operatorname{sign}(u-0.5)\cdot\sqrt{|u-1|},\quad\text{where }u\sim\mathcal{U}(0,1)$.
-These parameters are encoded in the diagonal of the matrix $D_{\alpha}$.
+To evaluate the robustness of our estimators, we introduce a random contamination mechanism.
+For each marginal distribution, let $r\in(0,1)$ denote the contamination level.
+We randomly select $\lfloor nr\rfloor$ observations and replace them with values drawn from $\mathcal{Y}\sim N(0,5)$.
+Moreover, we apply the power transformation to the contaminated data, which is the following.
+Let $g_{0}(t):=\operatorname{sign}(t)|t|^{\gamma}$ where $\gamma>0$ is a parameter.
+The power transformation for the $j$th dimension is defined as
+$g_{j}^{0}\left(z_{j}\right):=\frac{g_{0}\left(z_{j}-\mu_{j}\right)}{\sqrt{\int g_{0}^{2}\left(t-\mu_{j}\right)\phi\left(\left(t-\mu_{j}\right)/\sigma_{j}\right)dt}},$
+where $\phi(\cdot)$ is the standard Gaussian density function.
+These two procedures applied to the data induce a departure from normality, resulting in marginal distributions characterized by asymmetry and heavier tails compared to the Gaussian case.
+We consider three contamination levels (r = 0.05, 0.1, and 0.2), and three different parameters of the power function ($\gamma$ = 1.5, 2.5, and 3), considering a total of nine scenarios.
+In this analysis, we want to estimate the asymmetry of the data estimating the skewness of the data encoded in the matrix $D_{\alpha}$ of the distributions $f(\boldsymbol{X})\sim\operatorname{CSN}_{p,p}(0,\Sigma^{0},D_{\alpha}L,0,D_{\kappa})$, and $f(\boldsymbol{X})\sim\operatorname{SUT}_{p,p}(0,\Sigma^{0},D_{\alpha}L,0,D_{\kappa})$.
+Regarding the closed skew-normal, we use the EM algorithm described in abe2021algorithm, for the unified skew-$t$ we employ the likelihood-based method from the R package sn developed by azzalini_pkg_sn.
 After the data generation, we compute the skew-(S)KEPTIC and the SKEPTIC estimators, which are then plugged into the glasso algorithm in (11) to obtain $\Omega$.
-The glasso algorithm depends on a tuning parameter $\lambda_{G}$ whose the choice is performed applying the StARS method in Liu et al., (2010), and when the value is found for each dataset we consider a set of 30 tuning parameters in the range $(\lambda_{G}-0.1,\lambda_{G}+0.1)$.
+The glasso algorithm depends on a tuning parameter $\lambda_{G}$ whose choice is performed by applying the StARS method in liu2010stability, which selects the regularization parameter based on the stability of the estimated graph under subsampling, and when the value is found for each dataset, we consider a set of 30 tuning parameters in the range $(\lambda_{G}-0.1,\lambda_{G}+0.1)$.
 To evaluate the accuracy of the methods, we use the receiver operating curve ($\operatorname{ROC}$).
 We define the false positive rate and true positive rate for a fixed set of tuning parameters $\lambda_{G}$.
-When the tuning parameter is determined we estimate the graph: $\hat{G}^{\lambda_{G}}=(V,\hat{E}^{\lambda_{G}})$, to evaluate it, we define the number of false positives for a given $\lambda_{G}$, $\operatorname{FP}(\lambda_{G})$, as the number of edges in $\hat{E}^{\lambda_{G}}$ but not in $E$ and the number of false negatives, $\operatorname{FN}(\lambda_{G})$, as the number of edges in $E$ but not in $\hat{E}^{\lambda_{G}}$.
+When the tuning parameter is determined we estimate the graph: $\widehat{G}^{\lambda_{G}}=(V,\widehat{E}^{\lambda_{G}})$, to evaluate it, we define the number of false positives for a given $\lambda_{G}$, $\operatorname{FP}(\lambda_{G})$, as the number of edges in $\widehat{E}^{\lambda_{G}}$ but not in $E$ and the number of false negatives, $\operatorname{FN}(\lambda_{G})$, as the number of edges in $E$ but not in $\widehat{E}^{\lambda_{G}}$.
 We consider the rate of the number of false negatives and false positives:
@@
 $(\operatorname{FNR}(\lambda_{G}),1-\operatorname{FPR}(\lambda_{G})).$ (19)
-Graphically, we show averaged ROC curves for all the considered scenarios over 100 trials, plotted using the coordinates in (18).
+Graphically, we show averaged ROC curves for all the considered scenarios over 50 trials, plotted using the coordinates in (19).
 Fig.
-1 shows that the skew-(S)KEPTIC estimators generally outperform the Pearson correlation approach in non-Gaussian and improve upon the standard SKEPTIC estimator.
-Adjusting for skewness, also for excess kurtosis in the case of the SUT distribution, enhances the estimation of $\Omega$, especially in settings where the distribution presents heavy tails.
-In some cases, this leads to better recovery compared to the original SKEPTIC estimator, which has been shown to be optimal in Liu et al., 2012a .
-Table 2 provides a comprehensive overview of the simulation results, presenting quantitative comparisons of various estimators.
-The average AUC values of the skew-(S)KEPTIC estimators highlight strong performance in graph recovery under non-Gaussian conditions while still maintaining reliability when applied to normally distributed data.
-These findings suggest that considering estimators that include the skewness and accommodate for heavier tails, extending the nonparanormal distribution, do not hinder the recovery of the $\Omega$ structure.
+1 illustrates that the skew-(S)KEPTIC estimators yield results comparable to those of the SKEPTIC estimator, which has been shown in liu2012high to be optimal for a similar class of data.
+Importantly, incorporating skewness does not deteriorate the recovery of the true precision matrix $\Omega$.
+A more detailed comparison is provided in Table 2, which reports the values (in percentage) of the Area Under the Curve (AUC), False Positive Rate (FPR), and False Negative Rate (FNR).
+Across all scenarios, taking into account both data contamination and power-function transformations, the elliptical skew-(S)KEPTIC estimators exhibit lower false positive rates and higher false negative rates, reflecting the smaller off-diagonal entries relative to those of the SKEPTIC estimator.
+Given that the tuning parameter $\lambda_{G}$ is fixed across all methods, it is not expected that the elliptical skew-(S)KEPTIC estimators outperform the SKEPTIC estimator in terms of overall graph recovery.
+Nevertheless, it is noteworthy that the skew-(S)KEPTIC estimators consistently exhibit superior performance compared to the Pearson correlation coefficient in the recovery of graphical structures.
+These findings suggest that considering estimators that include the skewness of a skew-normal or a skew-t distribution and extend the nonparanormal SKEPTIC estimator, do not hinder the recovery of the $\Omega$ structure.
 On the contrary, they may offer a practical advantage due to the robustness across different data distributions.
-Figure 1: Comparison of average ROC curves, using the glasso algorithm, of the elliptical skew-(S)KEPTIC estimators with SKEPTIC estimators.
+Figure 1: Comparison of average ROC curves, using the glasso algorithm, of the elliptical skew-(S)KEPTIC estimators with the SKEPTIC estimator.
 Results are determined over 50 trials.
-We consider, by row, three different data generations and, by column, three different dataset sizes.
-Table 2: Quantitative comparisons of the 5 correlation matrix estimators using different distributions to generate data, and compute the $\Omega$ matrix.
+We consider, by row, three different data contamination and, by column, three different values of the parameter $\gamma$ that drives the power transformation.
+Table 2: Quantitative comparisons of the 6 correlation matrix estimators using different distributions to generate data, and compute the $\Omega$ matrix.
 The graphs are estimated using the glasso algorithm.
 The values are in percentages.
-SKEPTIC ($\hat{\rho}$) SKEPTIC ($\hat{\tau}$) SKEW-SKEPTIC ($\hat{\rho}$) SKEW-SKEPTIC ($\hat{\tau}$) PEARSON
-$n$ $p$ AUC FPR FNR AUC FPR FNR AUC FPR FNR AUC FPR FNR AUC FPR FNR
-N DATA 200 100 85.0 2.06 27.4 85.0 2.07 27.5 92.0 11.7 4.8 92.0 12.8 3.7 86.0 1.8 25.7
-s.e (10.0) (2.6) (21.6) (10.0) (2.6) (21.6) (2.0) (7.7) (3.2) (3.0) (7.8) (2.2) (10.0) (2.3) (21.7)
-1000 250 98.3 2.7 0.7 98.0 2.7 0.7 93.0 13.2 6.3 90.5 16.7 11.8 98.5 2.4 0.6
-s.e (3.0) (5.5) (1.3) (3.0) (5.5) (1.4) (12.5) (24.6) (24.3) (16.0) (31.4) (32.3) (2.4) (4.9) (1.1)
-500 300 94.0 1.3 9.8 94.0 1.3 9.7 96.0 7.8 0.3 96.0 8.8 0.1 95.0 1.1 9.4
-s.e (7.0) (2.4) (14.4) (7.0) (2.5) (14.2) (4.0) (8.3) (0.5) (4.0) (8.7) (0.2) (7.0) (2.1) (14.1)
-CSN DATA 200 100 86.0 5.0 22.8 86.0 5.0 22.2 88.0 17.3 6.3 88.0 18.5 5.7 83.0 4.4 28.9
-s.e (10.0) (4.9) (21.1) (10.0) (4.9) (20.8) (4.0) (10.2) (2.1) (5.0) (10.2) (2.1) (10.0) (4.7) (24.1)
-1000 250 98.0 1.6 3.2 98.0 1.6 3.0 96.0 8.3 0.7 94.2 11.0 4.5 96.0 1.3 6.7
-s.e (3.0) (2.5) (6.6) (3.0) (2.5) (6.2) (5.7) (11.3) (8.1) (10.0) (20.6) (20.7) (6.0) (2.2) (12.5)
-500 300 92.0 1.5 14.7 92.1 1.6 14.2 95.0 8.3 1.7 95.0 9.4 0.9 89.5 1.2 19.7
-s.e (9.1) (2.2) (19.3) (8.9) (2.2) (18.8) (3.4) (7.8) (1.6) (4.0) (8.3) (0.6) (11.7) (1.8) (24.4)
-SUT DATA 200 100 80.0 2.9 37.4 80.0 3.0 36.7 (-) (-) (-) 89.0 15.6 7.1 71.0 2.2 54.9
-s.e (10.0) (3.8) (24.2) (10.0) (3.9) (37.4) (-) (-) (-) (2.0) (6.3) (1.7) (12.0) (3.2) (26.8)
-1000 250 97.0 2.1 4.7 97.0 2.1 4.5 (-) (-) (-) 96.0 7.7 0.1 89.0 1.7 20.3
-s.e (4.0) (4.5) (8.6) (4.0) (4.5) (8.3) (-) (-) (-) (3.7) (7.4) (0.01) (12.0) (4.2) (25.2)
-500 300 89.3 1.9 19.8 89.6 1.9 18.9 (-) (-) (-) 94.4 9.7 1.5 80.0 1.6 37.6
-s.e (9.1) (3.7) (19.8) (9.0) (3.7) (19.5) (-) (-) (-) (3.3) (6.9) (0.7) (14.4) (3.3) (30.5)
+SKEPTIC ($\widehat{\rho}$) SKEPTIC ($\widehat{\tau}$) SKEW-SKEPTIC ($\widehat{\rho}$) SKEW-SKEPTIC ($\widehat{\tau}$) SKEW-KEPTIC PEARSON
+AUC FPR FNR AUC FPR FNR AUC FPR FNR AUC FPR FNR AUC FPR FNR AUC FPR FNR
+r = 0.05 $\gamma$ = 1.5 91.0 6.8 11.8 91.0 6.9 11.5 91.0 6.6 12.2 91.0 6.7 11.9 90.0 6.4 13.6 54.0 14.6 78.4
+s.e (1.3) (2.1) (3.5) (1.4) (2.1) (3.5) (1.4) (2.0) (3.6) (1.4) (2.0) (3.6) (1.7) (2.0) (4.4) (1.4) (11.7) (9.9)
+r = 0.05 $\gamma$ = 2.5 87.8 17.8 6.5 88.8 17.9 6.4 88.8 17.6 6.8 88.8 17.7 6.6 88.7 17.5 7.2 50.1 30.9 68.3
+s.e (2.7) (6.0) (3.2) (2.8) (6.0) (3.2) (2.7) (5.9) (3.2) (2.7) (6.0) (3.4) (2.6) (6.0) (3.3) (0.9) (15.4) (14.3)
+r = 0.05 $\gamma$ = 3 84.8 20.8 9.5 84.8 21.0 9.4 84.8 20.7 9.7 84.8 20.8 9.6 84.7 20.6 10.0 49.6 41.5 58.1
+s.e (12.3) (10.5) (14.3) (12.3) (10.5) (14.4) (12.3) (10.6) (14.3) (12.3) (10.6) (14.3) (12.3) (10.6) (14.2) (0.7) (13.4) (12.8)
+r = 0.1 $\gamma$ = 1.5 81.0 3.4 34.6 81.4 3.5 33.7 80.6 3.3 35.5 81.0 3.4 34.6 80.1 3.2 36.7 50.4 24.2 74.8
+s.e (3.0) (2.0) (6.9) (3.0) (2.0) (6.9) (3.0) (1.9) (6.8) (3.0) (1.9) (7.0) (3.0) (1.9) (6.8) (0.5) (17.9) (17.2)
+r = 0.1 $\gamma$ = 2.5 84.8 7.8 22.8 85.1 7.7 22.2 84.6 7.4 23.4 84.8 7.5 22.9 84.5 7.3 23.8 49.5 39.8 59.9
+s.e (1.8) (1.8) (4.9) (1.7) (1.8) (4.6) (1.8) (1.8) (5.0) (1.8) (1.8) (4.9) (1.8) (1.8) (4.9) (0.6) (15.6) (15.1)
+r = 0.1 $\gamma$ = 3 86.6 12.8 14.0 86.7 13.0 13.6 86.4 12.7 14.4 86.6 12.8 14.0 86.3 12.6 14.8 50.0 31.6 67.9
+s.e (2.2) (3.4) (6.0) (2.2) (3.4) (5.7) (2.3) (3.6) (6.1) (2.2) (3.4) (5.8) (2.3) (3.4) (6.0) (0.7) (15.0) (14.5)
+r = 0.2 $\gamma$ = 1.5 68.1 2.9 61.0 68.5 2.9 60.1 67.8 2.8 61.6 68.2 2.8 60.9 67.5 2.7 62.4 50.2 31.8 67.7
+s.e (4.2) (1.5) (9.5) (4.3) (1.5) (9.6) (4.2) (1.5) (9.5) (4.3) (1.5) (9.6) (4.3) (1.4) (9.6) (0.4) (12.3) (11.8)
+r = 0.2 $\gamma$ = 2.5 70.5 5.3 53.7 70.8 5.4 53.0 70.2 5.2 54.4 70.5 5.3 53.7 70.3 5.2 54.3 50.1 31.1 68.4
+s.e (3.1) (1.8) (7.8) (3.3) (1.8) (8.0) (3.2) (1.8) (7.9) (3.3) (1.8) (8.1) (3.3) (1.8) (8.1) (0.6) (15.6) (14.9)
+r = 0.2 $\gamma$ = 3 71.9 6.6 49.6 72.3 6.6 48.8 71.7 6.4 50.2 72.0 6.5 49.5 71.5 6.3 50.7 50.1 30.9 68.5
+s.e (4.0) (3.3) (9.9) (3.8) (3.4) (9.7) (3.9) (3.3) (9.8) (3.8) (3.2) (9.7) (3.9) (3.2) (9.5) (0.6) (14.5) (14.0)
 AUC = Area Under Curve, FPR = False Positive Rate, and FNR = False Negative Rate.
@@
 In the empirical analysis, we study the conditional dependence relations on daily returns of S&P500500 companies in the period from $04/01/2016$ to $31/12/2022$ downloaded from Yahoo!
-Finance, using the proposed elliptical skew-(S)KEPTIC estimators and comparing it with the SKEPTIC.
-After cleaning the data, the dataset consists of 454 stocks, each counting 1762 observations.
-The log-returns of the stocks in the S&P500 index are empirically skewed and present heavier tails than the normal distribution (Cont, (2001)).
+Finance, using the proposed elliptical skew-(S)KEPTIC estimators and comparing them with the SKEPTIC estimator.
+After cleaning the data, the dataset consists of 454 stocks, each with 1762 observations.
+The log-returns of the stocks in the S&P500500 index are empirically skewed and exhibit heavier tails than the normal distribution (cont2001empirical).
 Figure 2: Illustration of leptokurtic distributions of daily log-returns, with a relevant skewness component. The stock data are from $04/01/2016$ to $31/12/2022$.
-As a motivating example, we benchmark the analysis of Han and Liu, (2014), reporting the asymmetry of log-returns in Figure 2, where we plot histograms of the returns of three stocks in our sample ”eBay Inc.”, ”Coca-Cola” and ”Wells Fargo” belonging to three different sectors, respectively, Consumer Discretionary, Consumer Staples, and Financials.
-We appreciate the departure from normality reflected in leptokurtic distribution with a relevant skewness component.
-Moreover, we investigate the Gaussian assumption of the asset log-returns by computing normality tests with different significance levels at 5% and 1%.
+As a motivating example, we benchmark the analysis of han2014scale, reporting the asymmetry of log-returns in Figure 2, where we plot histograms of the returns of three stocks in our sample ”eBay Inc.”, ”Coca-Cola” and ”Wells Fargo” belonging to three different sectors, respectively, Consumer Discretionary, Consumer Staples, and Financials.
+We appreciate the departure from normality reflected in a leptokurtic distribution with a relevant skewness component.
+Moreover, we investigate the Gaussian assumption of the asset log-returns by computing normality tests with different significance levels at 1% and 5%.
 Table 3 indicates a clear and complete deviation from normality in the stock data.
@@
 5% 454 454 454
-To estimate the graphs, given the large number of variables relative to observations we determine the tuning parameter $\lambda_{G}$ using the Extended Bayesian Information Criteria (EBIC) of Chen and Chen, (2008) to ensure a balanced trade-off between model fit and complexity.
-$\operatorname{EBIC}=-2L+E\log(n)+4\gamma E\log(p).$ (19)
-The hyperparameter $\gamma\in\left[0,1\right]$, with greater values of $\gamma$ implying a stronger penalization, $L$ denotes the log-likelihood, $n$ the sample size, $E$ the number of non-zero edges and $p$ the number of nodes.
-We set $\gamma=1$, ensuring a strong penalization, as desired.
-The tuning parameter $\lambda_{G}$ is set equal to $0.54$ the value that minimizes the scores of the first-order differences in EBIC.
-To align the log-returns to the unified skew-$t$ distribution for computing the elliptical skew-KEPTIC estimator, we estimate the degrees of freedom of the distribution.
-We follow Arellano-Valle and Genton, 2010a using the maximum likelihood estimation method.
-In likelihood methods, the underlying assumption is that the observations are independent, therefore, we apply the best ARIMA model to the series using the R function auto.arima to filter our returns into a white noise series, and we apply the likelihood method to the filtered series to estimate the degrees of freedom using the command mst.mle from the R package sn developed by Azzalini, (2023).
+To estimate the graphs, given the large number of variables relative to observations, we determine the tuning parameter $\lambda_{G}$ using the StARS approach of liu2010stability.
+The tuning parameter $\lambda_{G}$ is set equal to $0.58$.
+To compute the elliptical skew-(S)KEPTIC estimators, we need to estimate the skewness parameters for the closed skew-normal distribution and for the unified skew-$t$ distribution.
+The skewness parameters are estimated using the same procedure as described in the numerical simulations.
 Figure 3 shows the estimated graphs using SKEPTIC, elliptical skew-SKEPTIC, and elliptical skew-KEPITC estimators; all the correlation matrix estimators are based on Kendall’s tau statistic.
@@
 The stocks that belong to the same sector tend to cluster, and those at the periphery of the network are uncorrelated with the others during the considered period, thus, they are uncorrelated with the movement of the market.
-Peralta and Zareei, (2016) show that those stocks have a greater weight in constructing a portfolio using the Markowitz approach.
+peralta2016network show that those stocks have a greater weight in constructing a portfolio using the Markowitz approach.
 Therefore, in large-dimensions, a network approach can be valuable for identifying stocks with the least correlation to the market.
-Table 4 demonstrates that incorporating skewness and kurtosis components leads to a sparser network estimation compared to the normal case.
+That is, incorporating skewness and kurtosis components leads to a sparser network estimation compared to the normal case as reported in Table 4.
 Table 4: Summary statistics of the estimated networks using Kendall’s tau statistic for the correlation matrix estimation.
@@
 DIM SK STK SNK SK>>SNK SK>>STK SNK>>STK
-S&P 500 454 384 284 231 100 153 53
+S&P 500 454 344 329 341 3 15 12
 7 Concluding remarks
 We propose rank-based correlation matrix estimators, named the elliptical skew-(S)KEPTIC estimators, to provide semiparametric estimation of graphical models built on meta skew-elliptical distributions.
-The proposed framework includes two classes of estimators: the elliptical skew-KEPTIC estimator which employs Kendall’s tau statistic, and the elliptical skew-SKEPTIC estimator which also uses Spearman’s rho statistic to estimate the correlation matrix.
+The proposed framework includes two classes of estimators: the elliptical skew-KEPTIC estimator, which employs Kendall’s tau statistic, and the elliptical skew-SKEPTIC estimator, which also uses Spearman’s rho statistic to estimate the correlation matrix.
 We study the theoretical properties of the rank-based correlation matrix estimators and establish concentration bounds.
 These bounds yield statistical rates of convergence under the elementwise maximum norm for both the correlation matrix estimation and the resulting estimation of the precision matrix.
-We also study the sparsistency of the precision matrix, evaluating the performance of the elliptical skew-(S)KEPTIC estimators in correctly recovering the edges of the graphical model.
 The empirical analysis and numerical simulations demonstrate that the elliptical skew-(S)KEPTIC estimators yield the optimal performance when the data deviates from normality.
-Applied to real data, specifically the log-returns of S&P500 constituent stocks, the resulting graphs are sparser compared to those estimated under normality.
-Our findings show a reliable framework that is useful when dealing with data that exhibit high skewness which cannot be neglected, and its effectiveness is validated through the optimal graph recoveries of the numerical simulations.
+The application to the log-returns of the S&P500 constituent stocks shows resulting graphs that are sparser compared to those estimated under normality.
+Our findings demonstrate a reliable framework that is useful when dealing with data that exhibit high skewness which cannot be neglected, and its effectiveness is validated through the optimal graph recoveries of the numerical simulations.
 Appendix A Proofs of elliptical skew-(S)KEPTIC estimators
@@
 Proof.
-We consider $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa};f)$, where $f(\boldsymbol{X})\sim V_{0}^{-1/2}Z_{0}$, and
-$Z_{0}\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ where $D_{\kappa}=I_{p}$, and $V_{0}=1$.
-We compute the correlation matrix of $\boldsymbol{X}$ using Kendall’s tau statistic, $\Sigma^{0}_{\tau}=\left[\tau_{ij}\right]$, and we decompose $\Sigma^{0}_{\tau}=L^{\top}D_{\kappa}L$, aligning it $\Sigma^{0}_{\tau}$ to the assumption of $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa};f)$ exploiting (8).
-$\text{cov}(\boldsymbol{X})$ $=\left(L^{\top}(I-D^{2})L\right)^{-1},$
-$=(L^{\top}ML)^{-1}\quad\text{where:}\quad M=I-D^{2}.$
-The matrices $I$ and $D^{2}$ are both diagonal matrices, and the matrix $M$ is.
-Thus, we have a new decomposition of the correlation matrix in which the skewness parameter $\alpha$, belonging to the matrix $D_{\alpha}$, contributes to the determination of the linear relations.
-The matrix $(L^{\top}ML)$ is:
-$\left(\begin{array}[]{ccccc}m_{11}&\ell_{12}m_{11}&\ell_{13}m_{11}&\ldots&\ell_{1p}m_{11}\\ \ell_{21}m_{11}&\sum^{p-(p-2)}_{k=1}\ell_{k2}\ell_{2k}m_{kk}&\sum_{k=1}^{p-(p-2)}\ell_{k2}\ell_{2k}m_{kk}&\ldots&\sum_{k=1}^{p-(p-2)}\ell_{k2}\ell_{pk}m_{kk}\\ \ell_{31}m_{11}&\sum_{k=1}^{p-(p-2)}\ell_{k2}\ell_{3k}m_{kk}&\sum_{k=1}^{p-(p-3)}\ell_{k3}\ell_{3k}m_{kk}&\ldots&\sum_{k=1}^{p-(p-3)}\ell_{k3}\ell_{pk}m_{kk}\\ \vdots&\ldots&\ldots&\ddots&\vdots\\ \ell_{p1}m_{11}&\sum_{k=1}^{p-(p-2)}\ell_{k2}\ell_{pk}m_{kk}&\sum_{k=1}^{p-(p-3)}\ell_{k3}\ell_{pk}m_{kk}&\ldots&\sum_{k=1}^{p}\ell_{kp}\ell_{pk}m_{kk}\end{array}\right)$
-where the generic $m_{k,k}$ term is:
-$m_{kk}=I_{kk}-D^{2}_{kk}=1-\frac{2}{\pi}\frac{\alpha_{kk}^{2}}{1+\alpha_{kk}^{2}}.$
-Let $A_{{\tau}}\equiv L^{\top}ML$ denote the rescaled final correlation matrix of $\boldsymbol{X}$, estimated using Kendall’s tau statistic.
-We recall that the stochastic representation (7) of the CSN is a sum of two comonotic copula functions that have the invariance property, so the SKEPTIC estimator is the same for both the distributions, thus the elliptical skew-KEPTIC estimator is the sum of two SKEPTIC estimators and an estimated matrix, $A_{\tau}$, from $f(\boldsymbol{X})\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$.
-When $V_{0}\sim Gamma(\nu/2,\nu/2)$, we have that $f(\boldsymbol{X})\sim\operatorname{SUT}_{p,p}(0,\Sigma,D_{\alpha}L,\nu,0,D_{\kappa})$ or equivalently: $f(\boldsymbol{X})\sim\frac{Z_{0}}{\sqrt{V_{0}}}$.
-Since the positive support of Gamma random variable, and the fact that Kendall’s tau is invariant under monotone transformations, we state that the skew-KEPTIC estimator maintains the same ordinal measures of dependencies of the closed skew-normal distribution even when $f(\boldsymbol{X})\sim\operatorname{SUT}_{p,p}(0,\Omega^{-1},D_{\alpha}L,\nu,0,D_{\kappa})$.
-∎
+We consider $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}\!\big(\Omega^{-1},\,D_{\alpha}L,\,h^{(p)},\,0,\,D_{\kappa};\,f\big)$, where $f(\boldsymbol{X})\stackrel{{\scriptstyle d}}{{=}}V_{0}^{-1/2}Z_{0}$ and $Z_{0}\sim\operatorname{CSN}_{p,p}\!\big(0,\,\Omega^{-1},\,D_{\alpha}L,\,0,\,D_{\kappa}\big)$, with $D_{\kappa}=I_{p}$ and $V_{0}=1$.
+Under this framework, the stochastic representation of $\boldsymbol{X}$ coincides with (8); explicitly:
+$\boldsymbol{X}$ $\stackrel{{\scriptstyle d}}{{=}}L^{-1}D_{\kappa}^{-1/2}D_{\alpha}(I+D_{\alpha}^{2})^{-1/2}\,\boldsymbol{U}\;+\;L^{-1}D_{\kappa}^{-1/2}(I+D_{\alpha}^{2})^{-1/2}\,\boldsymbol{V}.$
+${\boldsymbol{X}}=D_{\alpha}(I+D_{\alpha}^{2})^{-1/2}\,\boldsymbol{U}\;+\;(I+D_{\alpha}^{2})^{-1/2}\,\boldsymbol{V}.$
+Here $\boldsymbol{V}\sim\operatorname{NPN}(0,\Sigma^{0},f)$, where $\Sigma^{0}=L^{-1}D_{\kappa}L$, meaning that its joint CDF is:
+$F_{V}(v_{1},\ldots,v_{p})=\Phi_{\Sigma^{0}}\!\big(\,\Phi^{-1}(F_{V_{1}}(v_{1})),\,\ldots,\,\Phi^{-1}(F_{V_{p}}(v_{p}))\,\big),$
+where each $F_{V_{j}}$ is estimated from normalized ranks as described in Section 2.2. Regarding the $U$ distribution:
+$U_{ij}=F_{\mathrm{HN}}^{-1}(\widehat{r}_{ij})=\Phi^{-1}\!\Big(\frac{\widehat{r}_{ij}+1}{2}\Big),$
+where $F_{\mathrm{HN}}^{-1}$ is the standard half-normal quantile function.
+By copula invariance under strictly increasing marginal transforms, $\operatorname{Copula}(U)=\operatorname{Copula}(V)$, (see e,g, embrechts2002correlation), this means that $U$ and $V$ share the same Gaussian copula dependence structure $\Sigma^{0}$.
+Therefore, the joint CDF of $U$ (Gaussian copula with half-normal marginals) is:
+$F_{U}(u_{1},\ldots,u_{p})=\Phi_{\Sigma^{0}}\!\big(\,\Phi^{-1}(F_{\mathrm{HN}}(u_{1})),\,\ldots,\,\Phi^{-1}(F_{\mathrm{HN}}(u_{p}))\,\big),\qquad u_{i}\geq 0,$
+and $\Sigma^{0}$ can be estimated with the SKEPTIC estimator, where the generic element of $\Sigma^{0}_{ij}=\sin\Big(\frac{\pi}{2}\tau_{ij}\Big)$.
+When $V_{0}\sim\mathrm{Gamma}(\nu/2,\nu/2)$, $\boldsymbol{X}$ reduces to a unified skew-$t$ (SUT) model and $V$ is a semiparametric $t$-copula.
+demarta2005t show that the identity $\Sigma^{0}_{ij}=\sin\!\big(\tfrac{\pi}{2}\tau_{ij}\big)$ still holds for $t$-copula, so the dependence structure is the same for the $t$ and half-$t$ distribution.
+Let $A_{0}=D_{\alpha}(I+D_{\alpha}^{2})^{-1/2}$ and $B_{0}=(I+D_{\alpha}^{2})^{-1/2}$, and denote by $S_{\tau}:=\sin\!\big(\tfrac{\pi}{2}\,T\big)$, where $T$ is the correlation matrix estimated with Kendall’s tau statistic, the elliptical skew-KEPTIC estimator is:
+$S^{\tau,\alpha}\;\equiv\;\operatorname{Cor}({\boldsymbol{X}})$ $=A_{0}\,S_{\tau}\,A_{0}\;+\;B_{0}\,S_{\tau}\,B_{0}$
+$=(I+D_{\alpha}^{2})^{-1/2}\,\Big(S_{\tau}+D_{\alpha}S_{\tau}D_{\alpha}\Big)\,(I+D_{\alpha}^{2})^{-1/2}.$
+In the element-wise form, for $i\neq j$, the estimator is:
+$S^{\tau,\alpha}_{ij}=\sin\!\Big(\tfrac{\pi}{2}\tau_{ij}\Big)\;\frac{1+\alpha_{i}\alpha_{j}}{\sqrt{(1+\alpha_{i}^{2})(1+\alpha_{j}^{2})}},\quad\text{and:}\quad S^{\tau,\alpha}_{ii}=1.$
 A.2 Proof of Lemma 3.5
@@
 We recall Lemma 2.1, where for the normal distribution, and for $i\neq j$, we have: ${S}_{ij}=2\sin\left(\frac{\pi}{6}{\rho}_{ij}\right)=\sin\left(\frac{\pi}{2}{\tau}_{ij}\right)$.
-Let $A_{\rho}$ denote the final correlation matrix of $\boldsymbol{X}$ with Spearman’s rho statistic, as in the normal distribution, we have: ${S}^{\rho,\alpha}_{ij}\equiv 4\sin\left(\frac{\pi}{6}A_{{\rho}_{t}{ij}}\right)=2\sin\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)$.
-∎
+Let $A_{0}=D_{\alpha}(I+D_{\alpha}^{2})^{-1/2}$ and $B_{0}=(I+D_{\alpha}^{2})^{-1/2}$, and denote by $S_{\rho}:=2\sin\!\big(\tfrac{\pi}{6}\,P\big)$, where $P$ is the correlation matrix estimated using Spearman’s rho statistic.
+Following the proof of Theorem 3.4, the elliptical skew-SKEPTIC estimator is:
+$S^{\rho,\alpha}\;\equiv\;\operatorname{Cor}({\boldsymbol{X}})$ $=A_{0}\,S_{\rho}\,A_{0}\;+\;B_{0}\,S_{\rho}\,B_{0}$
+$=(I+D_{\alpha}^{2})^{-1/2}\,\Big(S_{\rho}+D_{\alpha}S_{\rho}D_{\alpha}\Big)\,(I+D_{\alpha}^{2})^{-1/2}.$
+In the element-wise form, for $i\neq j$, the estimator is:
+$S^{\rho,\alpha}_{ij}=2\sin\!\Big(\tfrac{\pi}{6}\rho_{ij}\Big)\;\frac{1+\alpha_{i}\alpha_{j}}{\sqrt{(1+\alpha_{i}^{2})(1+\alpha_{j}^{2})}},\qquad S^{\rho,\alpha}_{ii}=1.$
 A.3 Proof of Theorem 5.1
-Proof.
-We have: $\hat{S}^{\rho,\alpha}_{ij}=$ $4\sin\left(\frac{\pi}{6}A_{\hat{\rho}_{ij}}\right)$ and we introduce $\Sigma^{\rho}$, the correlation matrix using Spearman’s rho statistic.
-We know that $A_{\rho}=L^{\top}ML$, where $M=I-D^{2}$.
-By Definition: $\Sigma^{\rho}=L^{\top}D_{\kappa}L$.
-This implies:
-$\left\lVert\widehat{D}_{\kappa}\right\rVert_{\max}\geq\left\lVert\widehat{M}\right\rVert_{\max},$
-$\left\lVert L^{\top}\widehat{D}_{\kappa}L\right\rVert_{\max}\geq\left\lVert L^{\top}\widehat{M}L\right\rVert_{\max},$
-$\left\lVert\widehat{\Sigma}^{\rho}\right\rVert_{\max}\geq\left\lVert A_{\widehat{\rho}}\right\rVert_{\max},$
-$\left\lVert 4\sin\left(\frac{\pi}{6}\widehat{\Sigma}^{\rho}\right)\right\rVert_{\max}\geq\left\lVert 4\sin\left(\frac{\pi}{6}A_{\widehat{\rho}}\right)\right\rVert_{\max}.$
-[…]
-$E(\widehat{\rho}_{ij})=\frac{6}{\pi(n+1)}\left[\arcsin\left(\Sigma_{ij}^{0}\right)+(n-2)\arcsin\left(\frac{\Sigma_{ij}^{0}}{2}\right)\right].$ (20)
-We rewrite (20) as:
-$\Sigma^{0}_{ij}=2\sin\left(\frac{\pi}{6}E(\hat{\rho}_{ij})+a_{ij}\right),\quad\text{where }a_{ij}=\frac{\pi E(\hat{\rho}_{ij})-2\arcsin(\Sigma^{0}_{ij})}{2(n-2)}.$
-The term $|a_{ij}|\leq\frac{\pi}{n-2}$. Therefore, for all $n>\frac{6\pi}{t}+2$, we have $|a_{ij}|<\frac{t}{6}$, and thus:
-$\text{pr}\left(\left\lVert 4\sin\left(\frac{\pi}{6}\hat{\Sigma}^{\rho}\right)-\Sigma^{0}\right\rVert_{\max}>t\right)$ $=\text{pr}\left(\max_{ij}\left|4\sin\left(\frac{\pi}{6}\hat{\rho}_{ij}\right)-2\sin\left(\frac{\pi}{6}E(\hat{\rho}_{ij})+a_{ij}\right)\right|>t\right)$
-$=\text{pr}\left(\max_{ij}\left|2\sin\left(\frac{\pi}{6}\hat{\rho}_{ij}\right)-\sin\left(\frac{\pi}{6}E(\hat{\rho}_{ij})+a_{ij}\right)\right|>\frac{t}{2}\right)$
-$\leq\text{pr}\left(\max_{ij}\frac{2\pi}{3}\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})-a_{ij}\right|>\frac{t}{2}\right)\quad\text{(Lipshitz function)}$
-$\leq\text{pr}\left(\max_{ij}\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})\right|>\frac{3t}{4\pi}-|a_{ij}|\right)$
-$\leq\text{pr}\left(\max_{ij}\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})\right|>\frac{3t}{4\pi}-\frac{t}{6}\right)$
-$=\text{pr}\left(\max_{ij}\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})\right|>\frac{(9-2\pi)t}{(2\pi-3)}\right)$
-$\leq\sum_{ij}\text{pr}\left(\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})\right|>\frac{(9-2\pi)t}{(2\pi-3)}\right)\quad\text{(Boole's inequality)}$
-$\leq p^{2}\exp\left(-\frac{nt^{2}(9-2\pi)^{2}}{(2\pi-3)^{2}}\right)\quad\text{(Hoeffding's inequality)}$
-Setting:
-$\beta=p^{2}\exp\left(\frac{(9-2\pi)^{2}nt^{2}}{(2\pi-3)^{2}}\right)$
-and solving for $t$, we get:
-$t=\sqrt{\frac{(2\pi-3)^{2}}{(9-2\pi)^{2}n}\log\left(\frac{p^{2}}{\beta}\right)}$
-When $\beta=\frac{1}{p}$, we have $\sqrt{\frac{(2\pi-3)^{2}}{(9-2\pi)^{2}n}\log\left({p^{3}}\right)}$. ∎
+Let $B_{ij}=\dfrac{1+\alpha_{i}\alpha_{j}}{\sqrt{(1+\alpha_{i}^{2})(1+\alpha_{j}^{2})}}$ and, for $i\neq j$, we write the target as $S^{\rho,\alpha}_{ij}=2\sin\!\big(\tfrac{\pi}{6}\rho_{ij}\big)\,B_{ij}$, and $S^{\rho,\alpha}_{ii}=1$.
+For $i\neq j$:
+$\left|\widehat{S}^{\rho,\alpha}_{ij}-{S}^{\rho,\alpha}_{ij}\right|$ $\leq\left|\widehat{B}_{ij}2\sin\!\big(\tfrac{\pi}{6}\widehat{\rho}_{ij}\big)-B_{ij}2\sin\!\big(\tfrac{\pi}{6}\rho_{ij}\big)\right|$
+$\leq|\widehat{B}_{ij}|\left|2\sin\!\big(\tfrac{\pi}{6}\widehat{\rho}_{ij}\big)-2\sin\!\big(\tfrac{\pi}{6}\rho_{ij}\big)\right|+|\widehat{B}_{ij}-B_{ij}|,$
+since, by construction, $|\widehat{B}_{ij}|\leq 1$. From liu2012high, we know that, with probability $1-1/p^{2}$, we have:
+$\max_{ij}\left|2\sin\!\big(\tfrac{\pi}{6}\widehat{\rho}_{ij}\big)-2\sin\!\big(\tfrac{\pi}{6}\rho_{ij}\big)\right|\leq 8\pi\sqrt{\frac{\log p}{n}}.$
+Regarding the term $|\widehat{B}_{ij}-B_{ij}|$, we define $f(a,b)=\dfrac{1+ab}{\sqrt{(1+a^{2})(1+b^{2})}}$ so that $\widehat{B}_{ij}=f(\widehat{\alpha}_{i},\widehat{\alpha}_{j})$ and $B_{ij}=f(\alpha_{i},\alpha_{j})$ .
+A direct calculation shows that:
+$\frac{\partial f}{\partial a}(a,b)=\frac{b-a}{(1+a^{2})^{3/2}(1+b^{2})^{1/2}},\qquad\frac{\partial f}{\partial b}(a,b)=\frac{a-b}{(1+a^{2})^{1/2}(1+b^{2})^{3/2}},$
+hence $|\partial f/\partial a|\leq 1$ and $|\partial f/\partial b|\leq 1$ on $\mathbb{R}^{2}$. By the mean value theorem in two variables,
+$|\widehat{B}_{ij}-B_{ij}|=|f(\widehat{\alpha}_{i},\widehat{\alpha}_{j})-f(\alpha_{i},\alpha_{j})|\leq|\widehat{\alpha}_{i}-\alpha_{i}|+|\widehat{\alpha}_{j}-\alpha_{j}|.$ (20)
+Therefore, we have that:
+$\max_{ij}\left|\widehat{S}^{\rho,\alpha}_{ij}-{S}^{\rho,\alpha}_{ij}\right|\leq 8\pi\sqrt{\frac{\log p}{n}}+2\max_{1\leq l\leq p}|\widehat{\alpha}_{l}-\alpha_{l}|$
+Define $A:=\max_{1\leq\ell\leq p}|\widehat{\alpha}_{\ell}-\alpha_{\ell}|$ and fix an arbitrary threshold $t_{\alpha}>0$. On the event $\{A\leq t_{\alpha}\}$ we have $2A\leq 2t_{\alpha}$. Hence:
+$\mathbb{P}\big(2A\leq 2t_{\alpha}\big)\;\geq\;1-\mathbb{P}(A>t_{\alpha}).$
+Equivalently, with probability at least $1-\mathbb{P}(A>t_{\alpha})$, the plug-in error term satisfies
+$2\,\max_{1\leq\ell\leq p}\big|\widehat{\alpha}_{\ell}-\alpha_{\ell}\big|\;\leq\;2t_{\alpha}.$
+Exploiting Boole’s inequality, we obtain:
+$\mathbb{P}\!\left(\|\widehat{S}^{\rho,\alpha}-\Sigma^{0}\|_{\max}\;\leq\;8\,\pi\,\sqrt{\frac{\log p}{n}}\;+\;2t_{\alpha}\right)\;\geq\;1-\frac{1}{p}-\mathbb{P}(A>t_{\alpha}).$
+In particular, if $t_{\alpha}$ is chosen so that $\mathbb{P}(A>t_{\alpha})\leq 1/p^{2}$, then with probability at least $1-2/p^{2}$,
+$\|\widehat{S}^{\rho,\alpha}-\Sigma^{0}\|_{\max}\;\leq\;8\,\pi\,\sqrt{\frac{\log p}{n}}\;+\;2t_{\alpha}.$
+Moreover, if we assume that the $\alpha$ parameters are correctly estimated using likelihood-based methods, we have that $A=O_{P}\!\left(\sqrt{\frac{\log p}{n}}\right)$, taking $t_{\alpha}=C_{\alpha}\sqrt{\frac{\log p}{n}}$, where $C_{\alpha}$ is a constant value that depends from $\alpha$, we obtain:
+$\|\widehat{S}^{\rho,\alpha}-\Sigma^{0}\|_{\max}\;\leq\;\big(8\,\pi+2C_{\alpha}\big)\sqrt{\frac{\log p}{n}}\quad,$
+with probability greater than $1-\tfrac{2}{p^{2}}$.
 A.4 Proof of Theorem 5.2
-Proof.
-Similarly to the case of Spearman’s rho, we have that $\hat{S}^{\tau,\alpha}_{ij}=$ $2\sin\left(\frac{\pi}{2}A_{\hat{\tau}_{ij}}\right)$.
-We introduce $\Sigma^{\tau}$, the correlation matrix using Kendall’s tau statistic.
-We know that $A_{\tau}=L^{\top}ML$, where $M=I-D^{2}$.
-By Definition: $\Sigma^{\tau}=L^{\top}D_{\kappa}L$.
-This implies that:
-$\left\lVert\widehat{D}_{\kappa}\right\rVert_{\max}\geq\left\lVert\widehat{M}\right\rVert_{\max},$
-$\left\lVert L^{\top}\widehat{D}_{\kappa}L\right\rVert_{\max}\geq\left\lVert L^{\top}\widehat{M}L\right\rVert_{\max},$
-$\left\lVert\widehat{\Sigma}^{\tau}\right\rVert_{\max}\geq\left\lVert A_{\widehat{\tau}}\right\rVert_{\max},$
-$\left\lVert 2\sin\left(\frac{\pi}{2}\widehat{\Sigma}^{\tau}\right)\right\rVert_{\max}\geq\left\lVert 2\sin\left(\frac{\pi}{2}A_{\widehat{\tau}}\right)\right\rVert_{\max}.$
-[…]
-$E(\hat{\tau}_{ij})=\tau_{ij},$
-where
-$\tau_{ij}:=\operatorname{Corr}(\operatorname{sign}(X_{i}-\widetilde{X}_{i}),\operatorname{sign}(X_{j}-\widetilde{X}_{j}))$
-is the population version.
-$\text{pr}\left(\left\lVert 2\sin\left(\frac{\pi}{2}\hat{\Sigma}^{\tau}\right)-\Sigma^{0}\right\rVert_{\max}>t\right)\quad$ $=\text{pr}\left(\max_{ij}\ \left|2\sin\left(\frac{\pi}{2}\hat{\tau}_{ij}\right)-\sin\left(\frac{\pi}{2}\tau_{ij}\right)\right|>t\right)$
-$\leq\text{pr}\left(\max_{ij}\pi\left|\hat{\tau}_{ij}-\tau_{ij}\right|>t\right)\quad\text{(Lipshitz function)}$
-$\leq\sum_{ij}\text{pr}\left(\left|\hat{\tau}_{ij}-\tau_{ij}\right|>\frac{t}{\pi}\right)\quad\text{(Boole's inequality)}$
-$\leq p^{2}\exp\left(-\frac{nt^{2}}{2\pi^{2}}\right)\quad\text{(Hoeffding's inequality)}$
-Setting
-$\beta=p^{2}\exp\left(-\frac{nt^{2}}{2\pi^{2}}\right).$
-And solving for $t$, we get:
-$t=\sqrt{\frac{2\pi^{2}\log\left(\frac{2p^{2}}{\beta}\right)}{n}}.$
-When $\beta=\frac{1}{p}$, we have $\sqrt{\frac{2\pi^{2}}{n}\log\left({p^{3}}\right)}.$
-A.5 Proof of Theorem 5.3
-We start by proving $\|\hat{L}-L\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$ and $\|\hat{M}-M\|_{\max}O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$.
-Kendall’s tau and Spearman’s rho statistics can be rewritten in terms of U-statistics, and their estimates of correlation matrix achieve $O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$ as rate of convergence.
-Under assumption (a), we know that $\Sigma^{0}$ has bounded eigenvalues, which implies:
-$\|\hat{L}-L\|_{\max}\leq\tau_{2}\|\hat{\Sigma}-\Sigma^{0}\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right).$
-Regarding the convergence of the estimated $\hat{M}$ matrix, we know that it has diagonal elements defined as: $\hat{m}_{kk}=1-\frac{2}{\pi}\frac{\hat{\alpha}_{kk}}{1+\hat{\alpha}_{kk}}$ where $\hat{\alpha}$ parameters are estimated via the EM algorithm introduced in Abe et al., (2021), for which we assume that the regulation conditions of Wu, (1983) hold.
-Then $\|\hat{M}-M\|_{\max}O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$.
-We find bounds for $\|\hat{\Omega}-\Omega\|_{\max}$.
-$\|\hat{\Omega}-\Omega\|_{\max}$ $=\|2\sin\left(\frac{\pi}{2}(\hat{L}^{\top}\hat{M}\hat{L}^{\top})\right)^{-1}-2\sin\left(\frac{\pi}{2}({L}^{\top}{M}{L}^{\top})\right)^{-1}\|_{\max}$
-$\leq\|(\hat{L}^{\top}\hat{M}\hat{L})^{-1}-({L}^{\top}{M}{L})^{-1}\|_{\max}$
-$=\|(\hat{L}\hat{M}^{-1}\hat{L}^{\top})-({L}{M}^{-1}{L}^{\top})\|_{\max}$
-$=\|(\hat{L}-L)\hat{M}^{-1}\hat{L}^{\top}+L(\hat{M}^{-1}-M^{-1})\hat{L}^{\top}+LM^{-1}(\hat{L}-L)\|_{\max}$
-We define $A_{1}=(\hat{L}-L)\hat{M}^{-1}\hat{L}^{\top}$, $A_{2}=L(\hat{M}^{-1}-M^{-1})\hat{L}^{\top}$, and $A_{3}=LM^{-1}(\hat{L}^{\top}-L^{\top}_{0})$.
-$\|A_{1}\|_{\max}=\|(\hat{L}-L)\hat{M}^{-1}\hat{L}^{\top}\|_{\max}\leq\|\hat{L}-L\|_{\max}\|\hat{M}^{-1}\|_{\max}\|\hat{L}^{\top}\|_{1}$
-$\|\hat{L}-L\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right),\quad\|\hat{M}^{-1}\|_{\max}=O(1),\quad\|\hat{L}^{\top}\|_{1}=O(s).$
-Where $s$ are the non-zero elements in the $L$ matrix.
-$A_{2}=\|L(\hat{M}^{-1}-M^{-1})\hat{L}^{\top}\|_{\max}\leq\|L\|_{\infty}\|\hat{M}^{-1}-M^{-1}\|_{\max}\|\hat{L}^{\top}\|_{1}$
-Where $\|\hat{M}^{-1}-M^{-1}\|_{\max}=\max_{j}\left|\frac{1}{\hat{M}_{jj}}-\frac{1}{M_{0,jj}}\right|=\left|\frac{\hat{M}_{jj}-M_{0,jj}}{\hat{M}_{jj}M_{0jj}}\right|$.
-This means that: $\|\hat{M}^{-1}-M^{-1}\||_{\max}=O_{P}(\|\hat{M}-M\|_{\max})=O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$.
-$\|A_{3}\|_{\max}$ $=\|LM^{-1}(\hat{L}^{\top}-L^{\top}_{0})\|_{\max}$
-$\leq\|L\|_{\infty}\|M^{-1}\|_{\infty}\|\hat{L}^{\top}-L^{\top}_{0}\|_{\max}=O_{P}\left(s\sqrt{\frac{\log p}{n}}\right)$
-We conclude that:
-$\|\hat{\Omega}-\Omega\|_{\max}=O_{P}\left(s\sqrt{\frac{\log p}{n}}\right).$
-A.6 Proof of Theorem 5.4
-As discussed in Lam and Fan, (2009), we only need to prove a bound in $\|\hat{\Omega}-\Omega\|_{\max}$ for the three shrinkage procedures proposed to recovery $\Omega$.
-For the glasso algorithm, the proof follows the proof of Theorem 2, in Lam and Fan, (2009).
-For the CLIME, the bounds are provided in Theorem 4 in Cai et al., (2011), under the Frobenius norm, which implies $\|\cdot\|_{\max}$.
-And for the graphical Dantzig selector, the bounds are provided in Section 3 of Yuan, (2010) work.
+From liu2012high, we know that with probability $1-1/p^{2}$ we have that:
+$\left|\widehat{S}_{ij}^{\tau}-\Sigma_{ij}^{0}\right|\leq 2.45\pi\sqrt{\frac{\log p}{n}}$
+The results for our estimator become:
+$\max_{ij}\left|\widehat{S}^{\tau,\alpha}_{ij}-{S}^{\tau,\alpha}_{ij}\right|\leq 2.45\pi\sqrt{\frac{\log p}{n}}+2\max_{1\leq l\leq p}|\widehat{\alpha}_{l}-\alpha_{l}|,$
+where choosing a probability $1-1/p^{2}$, and assuming the convergence of the likelihood-based estimator, we obtain:
+$\|\widehat{S}^{\tau,\alpha}-\Sigma^{0}\|_{\max}\;\leq\;\big(2.45\,\pi+2C_{\alpha}\big)\sqrt{\frac{\log p}{n}},$
+with probability greater than $1-\tfrac{2}{p^{2}}$.